A Complete Bibliography of Publications in Advances and Applications in Statistics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

01 July 2023 Version 1.06

Title word cross-reference

1/n [Vrb14a]. 2 [CT01]. 2 × 2 [SR14b, SRCT17, TTT08]. 3 [BMZ^{+}17, She20, TRA^{+}18]. 5 [AJP16]. A [Das09, SP17]. AR(1) [CaN16, ESESA16, ESESIA17]. AR(2) [ESESI14, Vrb10]. AR(k) [Vrb17b]. AR(p) [CL06]. ARCH(1) [ZZF08]. ARIMA(p, d, q)(P, D, Q)_{L} [AS16b]. ARMA(1, 1) [BU16]. χ^{2} [Vrb14b]. D [DP08, Das09]. D_{s} [MY15]. e [PK16]. E(s^{2}) [KM02]. ϵ [SB09]. G [Aze18, KE19, PCB13]. GARCH(1, 1) [AB11]. GARCH(p, q) [SS18]. I(d) [GA09a]. INMA(1) [SPA18]. K [Ami09, Moh14, Zha20, Abo10, AS16a, ASS10, BA14, KNG11, TPP07b]. L [AM10a, EOB07, AAAA18]. L_{p} [CKS01]. MA(1) [BU16]. n [KR13]. \overline{X} [Guo11, GS10]. p [FAB20, GA09b, KK09, SNaN14]. Q [SdA12]. r [DN17]. ρ [Wan13]. σ^{2}Σ_{0} [KNG11]. t [FKL06, HNG16, MH09, Pol19, TH14]. τ [Wan13]. T L [AM10a]. U [Kel06].

-Bayes [SB09]. -constraint [PK16]. -digit [KR13]. -dimensional [BMZ^{+}17]. -fold [AJP16]. -geometric-discrete [AAAA18]. -identification [DN17]. -Means [Zha20, Moh14]. -mixing [SP17]. -model [Vrb10]. -moment [EOB07]. -moments [AM10a]. -negative [PCB13]. -norm [CKS01]. -Optimal [MY15, DP08, Das09, KM02]. -parameter [CT01]. -record [Abo10]. -step [She20]. -test [Kel06]. -type [ASS10]. -unit [TRA^{+}18]. -value [FAB20, KK09, SNaN14]. -values [GA09b]. -variable [Ami09]. -variate [KNG11].

1 [ASS10]. 18th [TRCH12]. 19 [EMM21, Els21]. 19th [TRCH12].

2 [TTM08].

AANA [Ryu04]. abrupt [CG07]. Absolute [AMB11, AKA20]. academic [BSGOFGPF17, RIT18]. Accelerated [SV12, HAE19, LS07, Ram16b]. Acceptance [Bak03, DB20, Ham20b, KA20]. access [DNH20, dA10a]. accident [TW20]. accidents [Kam11, KAAA21]. accord [EH12]. according [KKBB18]. accounting [Con06, Tol12]. accuracy [KPS08, Vrb14a, Vrb14b]. Accurate [Vrb13a, Vrb15, RW18, Vrb13b]. acquisition [AA12]. action [SB11]. action-decision [SB11]. active [Maz04]. activity [Con06]. Adaptation [TRCH12, Dog14]. Adaptive [Gri11, LTvR09, BRR11, BMS^{+}19, CCS16, JD16, Kan01, QKH18]. addendum [Vrb13b]. Adding [Afi16]. additional [SB06]. additive [FEM^{+}09, GA09b, Wei15]. Adjusting [AARCCA16, SRCTAA13]. administrative [LVM11]. adolescent [MRS19]. adult [DBD^{+}07]. adulthood [AZ19]. advantages [Maz04]. aero [WZG11]. aero-engine [WZG11]. affect [FEM^{+}09]. affecting [AA15b, DNH20, Ela10, EAHH13, FMI20, Hoa20]. after [RORGTR20]. against [ES16, PK17]. age [AA17, DDJM09, LKP04]. age-specific [AA17]. agents [RZC08]. aggregate [MWB17]. Aging [BA14]. agricultural [SMHN19, SMP21]. agro [VPO^{+}12]. agro-climatic [VPO^{+}12]. agrometeorological [VPKG12]. AIC [DLM11]. aided [Hua06]. AIDS [DDJM09]. Ain [EAHH13]. air [AZRS11, Yİ10, YAA16, Zak20]. Akaike [MSIT05]. Albania [ST11]. algal [uaSYHNQ20]. algebra [Kel05]. algebraic [FM10, FG11, Zoi09]. Algorithm [Zur01, şTE^{+}11, CEG09, DGS08, Fuj13, IFVA06, Kel05, Kel06, LB16, Lin09, NM04, SM12b, SH09, TP02, Ven15]. algorithms [BRR11, DCKD13, DN10, DT19, Kub19, LTTD20, SOAD06]. allocation [GS10, GAP21]. allocations [BLB10, Das09]. Almon [ÇA12]. Almost [Han03, PDK06, SZZ08, WS15b, Wu16]. along [KDNH18]. alteration [Ram16c]. alternate [DZS15]. alternative [AE11, GOPLD09, GÖ01, PCSMGO11, TH14]. alternatives [FHS20]. aluminum [ARA^{+}18, ARA^{+}19]. American [AF08]. among [MHW^{+}19, NBL15, Nis09, RIT18, WS15a]. Amortization [Nko09]. amplified [FG11]. Analysed [Din19]. analyses [Maz04]. Analysis [AA15a, Akt15, GO15, MRS19, MQLA19, ST11, SPM20, AEEG19, AAN21, AM12a, AC19, AZA^{+}19, ARA^{+}18, ARA^{+}19, Al 20, AC20, Ald21, AGW19, AA20c, AAK10, AAKAK11, AM12b, ACM07, AIR18, AOAS17, Azi18, BU16, BuH04, BKK13, Bui20a, CG07, CL06, CJ20, DDJM09, DRM15, DBSH11, DT07, DB20, DA10b, EB18, EM10, ES12, EKK03, FWT12, GOG07, Gar17, GD14, GZS18, GLARM18, GOPLD09, HH02, IAS19, JM19, JCACTIGO11, JKS21b, Kam11, Kho08, Kip13, Kip17, KNMM13, KKMMK20, KW02, LB16, LK06, LTvR09, LS16, MHW^{+}19, MeMW19, MFPO20, MSIT05, Moa10, Mos15, NBL15, ORMPL17, PG10, PPT02, PC14, PC18, PLRaB18, SM12a, SMYH19, SS21, SMP21, SH09, TRA^{+}18, TK04, TP02, WYZ10, Wan14, WNB13, Zak20, Zha20]. analytics [BFG21]. analyze [JSCI18, PPT02]. Analyzing [ZHMK09, CTF11, Che09]. Angle [Hür15, ZRX10]. ANNBP [AMA13]. annual [TPP05]. annuity [WPJC21b]. anode [ARA^{+}18, ARA^{+}19]. anomalies [MeMW19]. ANOVA [GA09b, Nin08, Ras11]. anti [ORMPL17]. anti-ideal [ORMPL17]. apple [DZS15]. Application [BuH04, DRM15, EKK03, HCJ21, KG21, NMW20, NV20, SM12b, TKD19, Tor21, Afi16, AZA^{+}19, AMA20, Ald13, ACM07, Arb08, ABS08, AOM17, AJP16, AAH18, BB18, BKK13, ÇA12, CMT16, DMLN10, DNH20, EAHH13, EE14, EE20a, FFX01, Fay19, GAM03, GBHP^{+}12, Hür03, KA19, KM08, Mar14, MRSR10, MHW^{+}19, MRS19, MKO21, OMO19, OT07, PPT02, PHTY07, Par10, PKKHM^{+}16, RMdSC13, RCAAMC19, SYO04, SC19b, Slo13, WPJC21a, WPJC21b, Wan14, XDL16, YAA16]. Applications [TS10, AK12, AKEE18, CF05, CBHMRA17, EE20b, FWT12, HNG16, HE16, KNMM13, KKD09, Kvå11, Lop14, Maz04, Mos15, PCB13, RA21, Yİ10, ZAJ20, ZRX10]. Applied [PC18, AZRS11, BMS05, DHY^{+}15, DT19, FMI20]. Applying [Abd10, BL01, Doa20a, Doa20b]. approach [AME18, Afi17, AK06, Ali10, Arb08, Arr13, AOAS17, BAZ20, BA19, BS21, BRY13, CG07, CL06, DGI18, DSG^{+}09, EFAEH17, ESAAA18, ERH17, FEM^{+}09, FEW09, HNG16, HKPGN09, Hua06, IAED15, Kan16, LX12, Liu10, Mad08, MN15, MM02, Nko09, Nko10, OOA11, PHTY07, PK16, PM21, SGM20, Sin16b, SLCB11, TW13, Thi09, Tol12, Tor20, XDL16, YKK03, YM12, ZBF18]. approaches [AYM15, AS05, EB18]. Appropriate [KKD09]. Approximate [Sal19a, Vrb17a, PDK06, TTT08, TMO07]. Approximation [Mar14, AAA^{+}16, AAE11, DA09, KB11, Vrb11, Zim08]. approximations [AAE12, Gar08, Gar17, Mar13]. Arab [AA15b, Moh13, SA16]. Arabia [YAA16]. Arabic [AA08]. arbitrage [MRS09]. ARCH [BH08, SZZ08]. Archimax [Hür05]. archives [LVM11]. ARDL [BA19]. area [Con06, KA19, SLKP08, WNB13]. area-biased [KA19]. areas [TPP05]. ARFIMA [NK13]. ARIMA [AC20, APL19, AA17, Cha20, Moh13, MJ11]. ARIMA/GARCH [Moh13]. arithmetic [YMT10]. ARL [Are19, BAS19]. array [APD^{+}10]. arrival [AS13]. art [UABW21]. artificial [El 18, Moh13, SC19a]. ascent [KPP09]. Asia [MeMW19]. Asian [Lin09]. aspect [JKS21a, JKS21b]. aspects [DFBL^{+}03]. asphalt [MWB17]. assess [FU19, GOSLVA12]. Assessing [AASR17, ICBS14, Sad10]. Assessment [GOG07, EH12, HSL09, LK06, PBS^{+}04, TCD20]. assets [SSL07]. assistance [ACM07]. associate [GS14]. associated [BSGOFGPF17, EMM20, Gar08, KR01, KPS08]. association [HLBF02, ICBS14, TH02, Um08, Van20]. associations [Maz04]. assumptions [SBZ03]. assurance [WPJC21a]. asthma [AZ19]. asymmetric [SJ12, Sun11, Yel20]. asymmetry [TMO07]. Asymptotic [El 11, Hsi10a, SVSO02, Vrb14a, AJM11, AARCCA16, CBHMRA12, Dor05, Pol19, SRCT17, WJ13]. Asymptotics [RG15]. atrial [FEM^{+}09]. attempt [PPT02]. attenuation [CFFVGO10]. attributable [Wu01]. Attribute [AAAJ17, PSKF17, Zam19]. attributes [AC19]. augmentation [HR07]. authorship [BRY13]. autocorrelated [BU16]. Automatic [RB03, SPM20]. automobile [XDL16]. autoregression [BH08, Mar14]. autoregressions [WDGP10]. autoregressive [AS17, APL20, ESESI15, PSA18, QDG17, Tin18, WH20, Yeh16]. auxiliary [EJ04, GAP21, IKA^{+}20, Zam19]. availability [HM14]. average [AdC14, AS16b, AS17, AKA17, APL20, BAS17, Cha20, HK20, IAED15, IAS19, SSA16, SPA18, Tin18].

balanced [Ald20a, Ald21, BV21, MKM12, SB20, VVV06]. ballistic [IWC03]. Balochistan [RIT18]. bandit [Wan07]. bandwidth [Yan12]. Bangkok [WNT20]. Bangladesh [DS21, SHM19]. Bani [MA20]. bank [AF19, Din20, AA15b]. banking [LHC01]. bankruptcy [ASAF09]. banks [AAN21, AA12, Doa20b, EH12]. bar [SCG01]. Barlow [Sto06b]. based [AM15b, AEB13, AMAAEA19, AB11, AARCMC20, Ami14, AATD20, AAED13, AIR18, Bak03, BA15, BM21, BLLR01, BLV06, BS21, BMS^{+}19, Cho07, CCS16, DGI18, Das20, DBD^{+}07, DT07, Dor05, DNO18, EJ04, ES15, El 18, EE19, FKE16, FHS20, FMA19, FCG10, GZS18, IFVA06, JL15, Kel06, KB11, LB16, LYK03, Mak14, MeMW19, MKO21, MFR^{+}19, Oid11, PK17, PCSMGO11, PACW13, PK16, PKKHM^{+}16, Ram16c, RE18, SMYH19, Sel15a, Sha10, SSBA20, SBZ03, TH14, Vrb13a, Vrb13b, WM08, Wu16, YQCSG10, Yan12, ZZN13, Zha20, ZW10]. Basel [EH12]. baseline [AS05]. bases [EKK03, Zoi09]. basis [VPO^{+}12]. Batch [AS13]. bathtub [AA20a]. Baum [Sto06a]. Bayes [Ami09, Cho07, GNB01, SJ12, SB09]. Bayesian [Abd04, AM10b, AKGU^{+}16, AA20a, Ald21, BMZ^{+}17, CG07, CKC08, Cho08, CJ12, DRM15, DA10b, EM10, ESS06, Fuj13, Gar17, GOdlFSLF08, GOPLD09, HA08, HM10, IuZHZ20, IE05, IAED15, KB11, Liu10, Mad08, Moa10, Mos15, NBL15, NK13, PHTY07, PvR04, PACW13, Ram16b, RS14, Sal19a, Sal18, SAGKS17, Sha19, She20, SS18, SH09, SLCB11, WY14, WNB13, YKK03, YB13, ZHMK09]. beam [RSM^{+}11]. bearing [DHSW20, DZS15, ZRX10]. bearing-angle [ZRX10]. Behavior [AARCMC20, Lee01, LX12, MM02, WJ13]. behavioral [MRS19]. behaviour [Kip11]. being [BAZ20]. Benchmarking [CW10]. Benford [Hür03, KR13]. Bernoulli [AAA^{+}16, HBJ09]. Bernoulli-exponential [HBJ09]. Bertalanffy [DMLN10]. best [MSIT05, Tin18]. best-fitting [MSIT05]. beta [CF14, DA09, NBA04, RIH20, RE18, SK08, Wan08]. beta-family [Wan08]. better [AGW19, LCL02]. between [AAN21, AKA20, AKAA20, Al 20, BA19, CM13, EI07, Han11, HSL09, HLBF02, Kaw10, Kaw14, KHY10, KNMM13, Mah10, MFR^{+}19, Moh13, NW16, aN16, SMG12, Sha08, TNN18, Wan13, WS15a]. beyond [Res04]. Bias [AS19, ABS08, CFFVGO10, JOO18, Kan16]. Bias-Corrected [AS19]. biased [Akm05, AK06, ÇA12, ESESIA17, EE20a, KA19, Ras13]. BIC [DLM11]. big [AAR19, Alm21]. binary [AJP16, DGHP10, KSM16, PKKHM^{+}16]. binary-data [DGHP10]. binding [KDNH18]. binomial [Abd04, APL20, GOdlFSLF08, Kaw10, Kaw14, KY15, PCB13, RCDO11, Ros20, SJ10]. bioequivalence [ESS06]. biprediction [Mül02]. Birnbaum [AM10b]. birth [AE17, El 20, SOO08, Swi08]. birth-death [Swi08]. Bivariate [Fam08, AdC14, AR14, Aze18, şTE^{+}11, Cho07, CKC08, Cho08, CJ12, CBHMRA11, DA10b, HA08, KE19, RYS08, Sad10, SS19, Sha19, SLCB11, TG11, Wan13, YH09]. black [Ram16a]. black-box [Ram16a]. Blackwelder [AACARC13]. block [BV21, KNMM13, MKM12, SB20, Vic08]. blocking [EKK04]. bloom [uaSYHNQ20]. blow [Alt18]. BLUE [MS06]. bold [XW14]. Bologna [JCACTIGO11]. Boltzmann [OAA19]. bonus [LLDD17]. Bootstrap [Ald13, BR12, BK02, ÇA12, Che02, GPP12, Mak14, OOA11, RSW14, RP15, SVSO02, Sha10]. bootstrap-based [Sha10]. Bootstrapping [IM15, QDG17]. border [PC18]. bound [SNaN14]. boundary [Wu14]. bounded [DGW^{+}10, MESAF16]. Bounds [OA09, SA10, Sel15a]. box [Ram16a, TS10]. branching [UMV13]. brand [GOSLVA12]. Brazil [AZRS11]. Brazzaville [Nko09]. break [BMS^{+}19]. breakdowns [OA09]. breast [ASFS11, NBL15]. brief [LCP05]. British [Yİ10]. Brownian [AMO20, AA20b, GR13, IS10, Yos15]. bugs [Mos15]. built [CCM11]. Burr [GAAAA04, KAF12, NMN18, Sal19a]. Burr-type [KAF12]. business [AA20c, Azi16, Hoa20]. businesses [Ela10]. busy [OA09].

cable [TRA^{+}18]. Cairo [Azi16]. calendar [AKGU^{+}16]. calibration [Coz08]. call [Muz12]. cancer [ASFS11, Abd10, Akt15, MKM^{+}15, NBL15, STPS10]. cancer-specific [MKM^{+}15]. cannot [SdA12]. Canonical [AOAS17, KNMM13]. Capability [HH02, LCL02, SAM19, SV04]. capita [BA19, OO20]. capital [DNH20, EE14, IE05, Lev13]. capture [TW13]. capture-recapture [TW13]. car [LLDD17]. card [DB20]. cardio2000 [PPT02]. care [Con06, Par10]. Carlo [AAA19, AKAA20, BRR11, HBA04, RYAD12]. carrying [Ald20b]. cascades [BB06]. case [AA15b, AM12b, BRY13, DN17, HCJ21, LLDD17, PG10, Tza08, WRP16]. cases [EMM21, Els21]. Cash [Doa20a]. cassava [Cha20]. Catania [AM12b]. categorical [AGM^{+}07, Kvå11]. categories [IT18, TH02, TMS08, YMT10]. Cauchy [Zgh03]. causal [YB13]. CDS [ZZN13]. cell [LPKPS10]. censored [AM10b, AA19, BAAA02, Cho08, DA10b, ES15, FMA19, LL09, Li11, Li15, Ras15, SJ12, SJ10, SR14a]. censoring [AASI18, AMAAEA19, BA15, CT01, CBSBHM10, LS07, Ram16b, SD19, She20]. censorship [El 11]. central [KR01, NBA04, WPJC21a]. century [TRCH12]. certain [AS16a, Dôk06, EKK03]. Chad [AAY^{+}19]. chain [BRR11, DLM11, MM02]. chains [AA08, DN17, MQLA19]. Champernowne [WDM14]. chance [HEDA18]. change [AKGU^{+}16, BAAA02, BuH04, BA04, Cha12, Cha18, CG07, DS21, Dia07, GYH18, HBA04, PRVATS14, SWCT19, TRCH12, Yİ10]. change-point [PRVATS14]. change-points [Dia07]. changes [AKGU^{+}16, CBHMRA17]. changing [MRS09]. characteristic [Zgh03]. characteristics [MWB17]. Characterization [AS16a, KNG11, Wan10, BA14]. Characterizations [RIH20]. chart [AS16b, AS17, Are19, AKA17, BAS17, Che02, Dha18, FHS20, GS10, NL21, PSA18, SPA18, Sun11]. charts [Alb12, GP10, Guo11, Kam11, Ula15]. checking [Sho08]. checks [LB16]. chemotaxonomic [AB17]. Chi [AS18, HP04, Hua06, SRCT17]. Chi-square [AS18, HP04, Hua06, SRCT17]. Chiang [SWCT19, TS10]. chicken [DMLN10]. child [BAZ20, GD14, GOdlFSLF08, Han09, Han11]. childhood [AZ19]. children [Han11]. China [DLP12, YT13]. Chinese [LYZ15]. Choice [CC08, ABS08, DFBL^{+}03, Sha10, Yan12]. choosing [EI07, NV20]. churn [BFG21]. circular [GBHP^{+}12, JK03, LC01]. cirrhosed [Azi19]. city [AM12b]. claims [MED14]. Clapper [IFVA06]. class [AB01, BV21, CTM18, Dôk06, Fel14, GS14, SA10, TW13]. classes [FG11]. classic [PCSMGO11]. classical [EB18, PKKHM^{+}16, Yos15]. classical-new [Yos15]. Classification [Lee01, Moh14, PD07, ASFS11, AKÜ20, Maz04, RIT18, SM12b, Ven15, WNT20]. classifications [TH02]. Climate [AKGU^{+}16, DS21, TRCH12]. climatic [VPO^{+}12, VPG^{+}14]. clinical [Arb08]. Closed [OAA19, ESESIA17]. closing [FMI20]. cluster [BS04, CCS16, QKH18, Van20, YT10]. clustered [QKH18]. Clustering [MeMW19, SYO04, APD^{+}10, BS21, DCKD13, IFVA06, Moh14, Zha20]. co [Ald13, KKMMK20, Zoi09]. co-inertia [KKMMK20]. co-integration [Ald13]. coal [Yİ10]. cocoa [BTANKB19]. codes [Kel05, Kel06]. coding [EKK04]. Coefficient [ZLZ06, AK19, Ald20a, FMTW10, MS06, NKaN11, aN16, SNaN14, WaNN14]. coefficients [Kan01, You05]. Cohen [Um05]. coherence [GK08, Lop14]. coherent [SdA12, dA10a]. cointegration [BA19, GPP12]. college [RIT18]. Colombo [APL20, SC19a]. colorectal [STPS10]. Comarca [MRSR10]. combination [AAE12, PKKHM^{+}16, XDL16]. combinations [KLK04]. combine [GZS18]. combined [RZC08, TW20]. Combining [APL19, Hag14, LYK03]. commercial [Din20, Doa20b, EH12]. common [Gun12, LB16, TNN17a, TNN17b, TN18]. communication [SA16]. communities [RR03]. companies [DLP12, Doa20a, Moh14, WYZ10]. company [FMI20, YSK06]. Comparative [AA17, Kip13, Kip17, Al 20, ABD05, BV09, FDdPF10, HLBF02, LK06, Tan10]. compare [SM12a]. Comparing [BR12, CW08, SRCTAA13, WNT20].

Comparison [AC20, AAY^{+}19, Cha18, DCKD13, DBD^{+}07, Kaw10, LTTD20, PCSMGO11, Abu02, AKA20, AKAA20, BRIA17, DT19, FSH10, GOG07, HK20, KY15, LPLH08, Mak14, MFR^{+}19, Moh13, RCAAMC19, SKS08, SR14c, WK18]. Comparisons [Sae17, Cho07, HR07, Nis09, SS21, WM08]. competition [Kip17]. competitiveness [Hoa20]. complementarity [JWLZ14]. Complete [Ryu04, FMA19]. complex [Arb08, El 18, NBA04, PC14]. component [AEB13, Akm05, AIR18, Kam11, LTvR09]. components [Al 20, FR17, KPS08, Nis09]. composite [GYH18, KPS08]. compound [Hsi10a, IS10, NC03, SS16]. Comprehensive [LYZ15]. compressed [MNC17]. computation [ASSS20, KB11, WH17]. Computational [IM09, AYM15, PK16]. computations [Gar15]. computer [Hua06, Kel05]. computer-aided [Hua06]. Computing [Kel06, DS08, Kel05]. concerning [Bor12, CBHM14, Sto08, Tza08]. concomitants [SS19]. concordance [FMTW10]. concrete [RSM^{+}11]. condition [BRR11, LK06, WZG11]. conditional [El 11, EE14, FWT12, LW04, MN15, Mar14, SSRG16, SAGKS17]. conditions [OKM16, Sto06a, Wu14]. conference [KM02]. Confidence [AK12, Cha03, LL09, NKaN11, NW16, aN16, SV04, TNN17a, TN18, TNN18, WaNN14, AK19, CW08, FH14, HYZ13, Kaw10, Kaw14, Niw17, PCSMGO11, QTH04, RW18, RCAAMC19, RG15, TNN17b, Vrb11, Vrb13a, Vrb13b, Vrb15]. confidential [IM15]. configurations [LB16]. confirmatory [DBSH11]. confirmed [Els21]. conformance [Alt18]. congestion [JSCI18]. Congo [Nko09]. Congolese [Nko10]. conjoint [AC19, ACM07]. connected [LB16, SYO04]. connectivity [PCGOMF18]. consecutive [Won20]. consensus [LB16]. considering [RMS^{+}11]. consistency [SP17]. constrained [Ame16, HEDA18, PLRaB18]. Constraint [Ma05, PK16]. constraints [GG07]. constructed [KM02]. constructing [IM15, SD19]. Construction [GS14, KMT12, MKM12, MKN16]. Constructive [KL15]. consumer [AC19, Din18, Doa20a]. consumption [MJ11]. containment [BRR11]. contaminated [Lin02]. contemporaneous [GG07]. contents [Han09]. contingency [BR12, DT19, IT18, KFT13, Law12, LPKPS10, MSIT05, Mül02, TTT08, TMO07, TMS08, YMT10, YMT13, YKT14]. continuous [Kon13, MESAF16, RB03, RSM^{+}11, ZX05, ZP08]. contoured [MPZ09, SBZ03, ZKZ13, ZZK13]. contraceptive [AEH18]. contribution [TK04]. Control [GP10, AK08, AS16b, AS17, AKA17, BAS17, BAS19, Che02, Dha18, FHS20, GS10, Haj08, HSC09, Mek10, NL21, SPA18, SCG01, UABW21, Ula15, ZP08]. controlling [AAAJ17, Mun11]. Convergence [AACARC13, DN10, Gar08, FDGP13, Han03, HK20, Ryu04]. convex [XDL16]. convolution [LWZ^{+}17]. Cooperation [YM12]. coordinate [FHS20]. coordinate-wise [FHS20]. copula [AAEE17, AH21, Das20, DA10b, EE17, Hür05, SMYH19, SLCB11, ZBF18, ZZN13]. copula-based [Das20, SMYH19]. copulas [CC08]. corporate [JKS21a, JKS21b]. corporeal [DMLN10]. Corrected [AS19]. correction [BM21, CFFVGO10, HM10, HB17, Kub19]. Corrections [Ula15]. correlated [BV21, Has18, MDTD10, SB20, TK04, YT10]. Correlates [GD14]. correlation [AOAS17, Bor12, CBHMRA11, DMV13, FMTW10, Hür15, KNMM13, LT20, Vrb17b, Wei15]. correlations [Dog14, YT10]. correspondence [PPT02]. corruption [AM12a]. cost [Ald20b, AAED13, PG10, SRMS17]. Costa [DMV13]. cotton [AAY^{+}19]. Council [YM12]. count [AAK21, Els21, KKD09, MRS19]. countably [MQLA19]. counterexample [Sto08]. countries [JCGO15, Nwa09, SA16, SMHN19, Sim06]. counts [AM12b, DT19, PFL06, PSA18]. Covariance [FM10, CBHMRA12, KNG11, OT07, SSRG16, SS16]. covariates [AM10b, AGM^{+}07, DGM20, DA10b, Liu10, TW13]. Coverage [CaN16, HZFZ06, Zim08]. covering [CCM11]. COVID [EMM21, Els21]. COVID-19 [EMM21]. Cox [CJ20, Mos15, ZX05]. CPI [ST11]. creating [ST11]. creatinine [Wan14]. credit [Dry11, EH12, FU19, RZC08]. crisis [Bui20b]. criteria [DLM11, EK04, RCAAMC19, Sha10]. criterion [MSIT05, NV20, Yeh16]. crop [MA17, VPO^{+}12]. crop-yield [VPO^{+}12]. cropping [AKGU^{+}16]. crops [Ald13, SMP21]. cross [AEH18, AJP16, DBD^{+}07, TH02, WH17, YQCSG10]. cross-classifications [TH02]. cross-sectional [AEH18, YQCSG10]. cross-validation [DBD^{+}07, WH17]. crossings [KAA08]. crude [AOAS17, Din18]. CSR [DLP12]. Cuadras [AAEE17, EE17]. cubic [LT20]. cultivation [HCJ21]. cumulative [AS17, Dog14, IT18]. current [LC09]. curtailment [PD09]. curve [AATD20, AMB11, Fel11, WDGP10]. curves [GZS18, RAM15, WNB13]. customer [AA12, FG10, lÇl19]. customers [FL09, Maz04]. cusum [SSA16, BAS19]. cut [BF16]. CVaR [Hur12].

Dagum [SD21]. daily [AH21, EMM21, Els21, FMI20, Yel20]. dam [BMS^{+}19]. dam-break [BMS^{+}19]. data [AAR19, Abd04, AAA19, AA11, AAM^{+}13, AA15a, Abd10, APD^{+}10, AM10b, AZRS11, AASI18, AZA^{+}19, Akt15, AMAAEA19, AA12, AGW19, AAK21, Alm21, AAKAK11, ACG10, AKÜ20, ABS08, AOM17, AGM^{+}07, BMZ^{+}17, BDNN15, BAAA02, BA04, BKK13, CTF11, Che09, Cho08, DGM20, DCKD13, DBSH11, DGHP10, DMLN10, DA10b, Dry11, DT19, ES15, ESESI14, ESAAA18, El 11, ES12, FEW09, GOG07, GYH18, GZS18, GLARM18, GO15, GBHP^{+}12, HZFZ06, HNG16, HR07, IM20, JM19, JSCI18, JGAF10, KKBB18, Kho08, Laa14, LL09, Lop14, LS16, MRS19, Mek10, Mos15, NMW20, NL21, Nko10, OMO19, PPT02, Par10, PBS^{+}04, PLRaB18, RMdSC13, RA21, SGM20, SOAD06, SJ10, SLKP08, SS21, TS10, Tor20, Um05, WPJC21a, Wan14, WH20, Won20, Yan08, YT10, Yan12, Yİ10, Zak20]. data-based [Yan12]. database [PPT02]. dataset [SKS08]. de-amplified [FG11]. DEA [AM15b]. DEA-based [AM15b]. death [MKM^{+}15, Swi08]. deaths [TW20]. debit [DB20]. decaying [Kip20]. December [GBHP^{+}12].

Decision [ASFS11, Stu02, GS08, PSKF17, SB11]. decisions [Gri11]. decline [AEH18, WDGP10]. Decomposition [TMS08, IWC03]. decrement [DSG^{+}09]. defecting [Sun11]. Deferred [DFBL^{+}03]. defining [EKK04]. definition [FWT12]. deformations [RMS^{+}11]. degradation [DHSW20, HAE19]. degrees [IM09]. delay [GS03]. delayed [Mad08]. delete [MEL16]. delete-group [MEL16]. Delhi [Sal18]. Delta [Hoa20]. demand [Arr13, PDK06, Par10, Slo13]. dengue [APL20]. density [Abd08, AK06, DGI18, JOO18, Kan16, KB11, SSBA20, TTM08, YH09]. denumerable [DLM11]. departure [IT18, Law12, TSU01, TH02]. dependence [AF08, Dor05, Hür05, LM11, PK17, TW13]. dependency [SMYH19, ZBF18]. dependent [AKA17, CEG09, El 11, GS03, HEDA18, HZFZ06, Mar13, MA19, MM02]. depending [BL01, DDJM09]. depression [GOdlFSLF08, ICBS14]. depressive [MHW^{+}19]. deprivation [BV09]. depth [Bui20b, Mek10]. derived [DA10b]. deriving [DMV13]. describing [EE14]. design [Alt18, BS04, Coz08, KBS05, KMT12, Laa14, MKM12, Nis09, Ram16a, TO05, TKD19]. designed [Das09]. designs [BV21, DP08, EKK03, EKK04, EK04, FR17, GS14, KM02, MY15, MKM12, MKN16, PC14, SB20, VVV06, Vic08]. detect [GPP12]. Detecting [Nin08, SKS08, YMM11]. Detection [Cha12, AKGU^{+}16, DGI18, DHY^{+}15, MBSJ19, Oid11]. deterioration [GS03, GB07]. Determinants [AF19, Din20, YM12]. Determination [Ald20b, Ald13, BLB10]. determine [Ald20a, Kam11, KG21]. Determining [Ras11, PD07]. deterministic [Swi08]. Dev [CF05]. developing [MQAA21, SMHN19, Sim06]. Development [VPKG12, AM12a, AATD20, Bui20a, SA16, TKD19, Van20, ZHMK09]. developments [AAR19]. device [Chr09]. diabetes [MA20]. diacritized [AA08]. diagnosis [ASFS11, PKKHM^{+}16]. diagnostic [KPS08, TP02]. Diagnostics [GR15, Víš14]. diagonal [DP08, SB20]. diagonals [KFT13]. diagonals-parameter [KFT13]. Difference [Han11, SMG12, Hag19, HBJ09, JPW11, Kaw10, Kaw14, MKN16, NW16, aN16, PC13, TO05, TNN18, WZG11, WJ13, Wu16]. difference-based [Wu16]. Differences [RIT18, NBL15, Sel15a]. different [AK19, AMAHE12, AASI17, Ali10, BLLR01, Cha18, DRW19, FEW09, Lev13, LTTD20, MSIT05, RSM^{+}11, RMS^{+}11, SMP21, Yel20]. differentially [PM21]. diffusion [Abu02, Kon13, YY20, Zim08]. digit [KR13]. digital [Gri07]. dimension [Hag14]. dimensional [BMZ^{+}17, MKM12, PvR04, Ram16a]. dimensions [KMT12]. direct [KG21, YT13, YM12]. directed [Lop14]. directions [Kho08]. Dirichlet [DÇ18, TTM08]. disadvantages [Maz04]. disasters [Yİ10]. discarding [Lin09]. discharges [KKBB18]. disclosure [DLP12]. Discontinuous [Fel11]. discovery [SB06]. Discrete [FS08, AAAA18, JS03, MKO21, NR05, ZX05]. discretization [GD10, Ram16c]. discriminant [CL06, EB18, ES12, GOG07, JM19, KW02]. disease [APL20, FFX01]. disequilibrium [FFX01]. dispersed [Has20]. dispersion [Kip13, Kip17]. displacement [NV20]. disposable [OO20]. dissolution [HSL09]. distance [FHS20, Gar17, JGAF10]. distance-based [FHS20]. distances [CF14, IFVA06]. distress [FFX01]. distributed [ÇA12, CZ12, FDdPF10, KR13, MNC17, NL21]. Distribution [Al 16, FCG10, Sel15b, TPP05, AJM11, AM10a, AMAHE12, AM10b, Afi16, AR16, AZIE19, AMAAEA19, AMA20, AM20, AMEGM12, ASSS20, AS18, AAE12, Ald20a, AA20c, AKEE18, AS19, AKA17, AOM17, AR14, AAED13, BA20, Bak03, BA15, BLF11, BMS05, Bor12, BAS17, BAS19, CMT16, CT01, CF14, CF05, CBSBHM10, CBHMRA17, DS08, DC20, DÇ18, DGG05, EE20a, EE20b, Ess17, EE19, FMA19, Fay19, FKL06, GNB01, GAM03, GK08, Had19, Ham20a, Ham20b, HNG16, Has20, Hsi08, HP04, Hua06, HZ10, IuZHZ20, JSS19, JZ20, KA19, KKBB18, Kar20, KNG11, KAF12, KA20, Kub19, KM08, Kvå11, LL09, Li11, Li15, LWZ^{+}17, LVX19, Lug16, MED14, MS06, MME21, Mar14, MKO21, NBA04, NA05, NLI16, NMN18, NR05, NMW20, OMO19, OAA19, PHTY07, PAL11, Pol19, PLRV15]. distribution [RMdSC13, RIH20, RA21, Ras13, RAM15, RE18, RCDO11, RGlHnSC20, SD21, Sal19a, Sal19b, SJ12, SC19b, SAGKS17, SS19, Sha19, SKS08, SJ10, SSL07, SBZ03, SNaN14, SRCT17, Tan10, TG11, TH14, Vrb14a, Vrb17a, WaNN14, ZAJ20]. Distribution-free [FCG10]. distributions [AK12, AYM15, Afi17, AS16a, ASSSAZ18, Alk18, Ami09, ASM17, BL01, BLLR01, BRIA17, şTE^{+}11, CEG09, DA10b, ES15, EOB07, FH14, Had19, HE16, Hsi08, Hsi10a, JK03, KE19, Li12, MN15, MPZ09, Moa10, Niw17, RM18, Ram16c, SK08, SD19, TNN18, TPP07b, TPP07a, Wan08, Wan10, WDM14, Wu02, Yel20, ZKZ13, ZZK13]. disturbances [GA09a]. divergence [HK20, TSU01, TH02]. diversified [SMHN19]. diversity [Zha20]. dividend [YSK06]. Dma [PSA18]. DNA [KDNH18]. DoE [BMS05]. Does [FEM^{+}09, Bui20b]. dollar [ES16, RORGTR20]. domain [CL06]. domains [WNB13]. domestic [Cha20]. dominance [BAZ20]. Double [JAO06, BAS17, Sel15b]. doubly [Li11, Li15, MFR^{+}19]. Downweighting [RR03]. driven [Els21, KS18]. driver [SMHN19]. drought [Sim06]. DSCIA [KKMMK20]. dual [KKMMK20, SS19]. duality [Hsi10b]. due [SB06]. Duration [BK12]. durations [Has18]. Dynamic [ZP08, AAA19, BF16, ESAAA18, Gar15, RZC08, SMHN19]. dynamical [FDGP13]. dynamically [YSK06]. dynamics [GF19, GKKTS14, LP13].

e-service [Azi20]. earnings [Abd04]. earthquake [DSG^{+}09]. Easterbrook [BRY13]. echo [MNC17]. econometrical [Nko09, Nko10]. economic [AM15a, DH20, GG07, KG21, Nko10, Tor21, Van20, YT13]. economical [Haj08]. economics [AK12]. economy [Din19, MQAA21]. EDC [DLM11]. EDF [CBSBHM10, KNG11]. Edgeworth [Vrb11]. education [BA19, Han09, JCACTIGO11, JCGO15, KHY10, El 20, SSBA20]. educational [AAK21, LVM11]. EEG [Lop14]. Effect [AJM11, AK19, CFFVGO10, AM12a, AZ19, AZIE19, Akt15, AdC14, DRW19, DH20, GS10, Kip17]. effective [SCG01]. Effectiveness [JKS21a, Chi14]. effects [AAA19, AMO21, BU16, BMS05, JM19, Liu10, SV12, SR14a, SBZ03, STPS10, UH13]. efficiency [AA11, ÇA12, Cos09, EH12, ES12, GS10, SRMS17, Tza08, WNT20]. Efficient [AMO20, HYZ13, YB13, Zam19, HTD18, Kho08, Ram16a, SH09]. EGX30 [Azi18]. Egypt [AE17, AEH18, AA11, AMA13, AAM^{+}13, Ald13, AAH18, Ela10, ES12, EMM21, Els21, FMM19, IE05, Kam11, Kam19, El 20, Moh13, MA17]. Egyptian [AM15b, AM15a, AA15b, AA12, Azi18, AF19, EE14, ES16, Moh14]. eigenvectors [CBHMRA12]. eight [KY15]. election [Tin18]. electricity [JKS21a, JKS21b]. elements [FSH10]. Eliya [WH20]. elliptically [MPZ09, SBZ03, ZKZ13, ZZK13]. emphasis [SSBA20]. Empirical [MRS09, PC13, QTH04, Wei15, AKGU^{+}16, AMO21, CM13, DH20, DFBL^{+}03, ES16, GCN20, Hsi08, JM19, KLK04, Kub19, MDTD10, RE18, Zgh03, ZX05]. employment [AE17]. energy [MJ11]. engine [DHY^{+}15, WZG11]. engines [WZG11]. Enhanced [Oid11, SJH18]. enhancement [Gri07]. Enhancing [Alt18]. EnKF [FDGP13]. Ensemble [BFG21, WDM14, CTM18, DGW^{+}10]. enterprises [DNH20]. entire [HH02]. entropy [BRIA17, ÇA12, RIH20, SJH18, TSU01]. envelope [MNC17]. envelopment [ES12, GLARM18, Kho08, PLRaB18]. environment [Han11]. environmental [HNG16, MHW^{+}19, SM12a]. epidemiological [PPT02]. epileptic [DGI18]. epsilon [AMEGM12]. epsilon-skew [AMEGM12]. equality [AYM15, LKP04]. equation [GOPLD09, SS16, Wu14]. equations [Abu02, FKE16, JPW11, NV20, WJ13]. equipment [MFPO20]. equivalence [PBS^{+}04]. equivalent [IM09]. ergodicity [Ven15]. Erlang [ASS10, EE20b, NLI16]. Erlang-truncated [EE20b, NLI16]. Errata [Vrb13b]. Error [CZ12, CW10, GOG07, KS03, LVX19, Mun11, RZC08, SS16, SSL07, WS15b, Wu17, YMM11]. error-models [CW10]. errors [BU16, BV21, BH08, DP08, Kan01, LV18, Liu10, SS16, SB20, TW16]. errors-in-variable [TW16]. Estate [AA15b, Alm21, Bui20b, WYZ10]. estimable [Beg11]. estimate [HZFZ06, IKA^{+}20, MME21, RS14, SV12]. estimated [AEB13, IM15, Ula15]. estimates [BK02]. Estimating [Kip20, MED14, MN18, MKM^{+}15, MDTD10, STPS10, Zgh03, Abd04, AK19, Ald20a, Arr13, ABS08, CT01, DSG^{+}09, JAO06, MN15, PM21, UH13]. Estimation [AMAHE12, AAAA18, AGM^{+}07, Dia07, ES15, ESESA16, EE19, Fuj13, FDdPF10, Hoa20, IM20, KK09, Li11, Li12, LVM11, PRVATS14, RZ14, Ras15, SS16, Sin16a, TN14, AK12, Abd08, AM10a, AH21, AEB13, AZIE19, Akm05, AK06, AG07, AB11, AS05, AS19, Ami09, BK12, BA15, BLB15, BDNN15, BuH04, ÇA12, CMT16, CKS01, DGPF12, Das20, DC20, DÇ18, DGHP10, DGG05, FCG10, FS08, GNB01, GOdlFSLF08, Hag19, HA08, HM10, HZ10, IuZHZ20, JZ20, JPW11, Kan01, Kan16, Kou02, LM11, LPKPS10, LP11, LP13, Li15, LCP05, LC09, Ma05, MN20, MPZ09, NK13, Özt01, PG10, PHTY07, QKH18, RYAD12, RM18, RZC08, RG15, Sal19a, Sal19b, SVSO02, SJ12, SAGKS17, Sha19, SP17, SJ10, Sin16b, Sto06b, SR14c, SS18, TO05, Thi08, Thi09, Tor21, TNR10, WK18]. estimation [Wu17, YXY19, YY20, YH09]. estimations [KE19, SD19, SLKP08]. estimator [AARCMC20, AAH18, CCS16, EFAEH17, GK08, IKA^{+}20, JD16, Kha13, KT08, KSM16, LT20, MS06, Mar13, MEL16, Oma18, Oma19, QH18, QKH18, ŞS10, TTT08, TW16, WS15b, Wu16, Yan12]. Estimators [JL15, AMO20, EJ04, ESESI14, ESESI15, EOB07, FHS20, FR17, GOG07, KAK18, Mag19, Sae17, TCD20, TMO07, Tza08, Vrb10, YAA16, Zam19]. ETF [RZC08]. Ethiopia [GD14]. Europe [JCACTIGO11]. European [JCGO15]. EV [Wu17]. evaluate [AA11, ES12, LHC01]. Evaluating [Che02, EK04, Mun11, NV20, XB10]. Evaluation [PD09, Yan08, AAK21, HSC09, LYZ15, LCL02, RZC08, SCG01]. event [AE17, El 20, SGM20, SLKP08]. events [EMM20, GF19]. evidence [AEH18, Din19, Din20, JM19, JZTS17, JCGO15, MRS09, TRCH12, Bui20b]. evolution [DDJM09, Nko09]. EWMA [ARAA17, Are19, AKA17]. Exact [AB01, Hsi08, MH09, OO07, SSA16, Are19, Kaw10, SR14b]. examine [AZ19]. example [DCKD13, YQCSG10]. examples [Laa12]. Exceedance [FG11]. excess [MED14, WM08]. exchange [AAK10, BB18, ES16, Kam19, MRS09, Nwa09, RORGTR20]. exchangeable [Sto11]. existence [GÖ01, Kho08]. exogenous [AS17, Are19]. expansion [Tsa02, Vrb14a]. expectation [FWT12]. expected [Ald20b]. expenditure [BA19, GKKTS14, JCGO15]. experiment [Das09]. experimental [IFVA06]. experimentation [DFBL^{+}03]. experiments [ACM07, BS04, OO07, Ras11, Sal05]. explanatory [KKD09, YS17]. Explicit [AS17, BAS17, SPA18, AS16b]. Exploratory [LS16, DBSH11, JKS21b]. Exploring [DHSW20, FDGP13]. explosive [Nko09]. exponential [AMAHE12, ASSS20, AKEE18, Ami09, AKA17, AR14, APL19, BRIA17, Cho07, CKC08, Cho08, EE20a, EE20b, EE19, FMA19, GYH18, GZS18, Had19, HA08, HM10, HBJ09, IuZHZ20, Lug16, Mad08, MED14, MS06, MKO21, NLI16, OT07, PBS^{+}04, QKH18, Ras15, Sal19b, SJ12, Sel15b, SJ10, SSA16, TO05, TNN17a, TN18, TNN18]. exponential-lognormal [GYH18]. exponentially [AS16b, Kip20]. Exponentiated [AKEE18, ASM17, NMN18, Sal19b, Bar09, CBSBHM10, FMA19, MED14, RMdSC13, Ras15, WRP16, ZAJ20]. exponentiated-exponential [MED14]. export [KG21, SMHN19]. exposed [MHW^{+}19]. exposure [Zak20, Zim08]. expressed [PM21]. expression [APD^{+}10, AS16b, OAA19, SSA16]. expressions [BAS17]. Extended [Bor12, EE20a, Ham20b, PSKF17, TPP07b]. extension [AAEE17, AM07, LCP05]. Extensions [KPP09]. extinction [Kip11]. Extra [Abd04]. Extra-variation [Abd04]. extreme [AS18, LL09, TPP05, XDL16]. extremes [Bis11, Dor05]. Eye [HKPGN09].

factor [BSGOFGPF17, Cho07, JKS21b, KDNH18, MeMW19]. factorial [EKK03, EKK04, EK04, Sal05]. Factors [Ela10, EMM20, SHM19, AA15b, DLP12, DNH20, EAHH13, FMI20, GOSLVA12, Hoa20, KAAA21, Maz04, MA20, RIT18, Van20, uaSYHNQ20]. failure [AA15b, AA20a, AA20c, Li15, SV12]. False [SB06]. families [CC08, Hur12]. Family [CTF11, ASSSAZ18, Alk18, ASM17, Aze18, Han11, HE16, JK03, KE19, Mun11, OT07, PCB13, SK08, SD19, TO05, Wan08]. family-wise [Mun11]. farming [HCJ21]. Farrington [AASR17]. Fast [Kub19, WH17]. father [Han11]. father-children [Han11]. fatigue [SAM19]. fault [DHY^{+}15, HM10]. faults [HB17]. feasible [Kho08]. Feature [QF20]. Features [EW15, DGHP10, LYK03]. feedback [CY04]. Female [El 20, AE17, DMLN10]. Fertility [AEH18, AA17]. FGM [SLCB11]. fibrillation [FEM^{+}09]. FIFA2018 [Azi20]. file [Laa14]. filling [Tor20]. filter [AB11, DGW^{+}10, ZRX10]. filtering [DHY^{+}15, Tol12, XW14]. finance [LX12, SC19b]. Financial [Bis11, Bui20a, AA15b, Bui20b, Gri11, Maz04, Moh13]. find [LB16]. finding [Cha18, ZBF18]. Finetti [Hür15]. Finite [Mül02, You05, BLB10, Bor12, BMS^{+}19, DS08, EJ04]. first [BMS^{+}19, DP08, El 20, PSA18]. fish [RSR^{+}18]. Fisher [SR14b]. fisheries [AAM^{+}13]. fit [AR16, CBHMRA17, Els21, KNG11, RE18, XB10]. fitness [Hun05]. Fitting [DGM20, SGM20, JCACTIGO11, Mah10, MRS19, MSIT05]. five [GD14, MKN16]. Fixed [RSW14, AAA19]. Flexible [KDNH18, Afi16, AR16]. floods [Sim06]. flow [MRSR10, Yİ10]. flows [BF16]. fluctuations [EE14]. FMRI [BMZ^{+}17, PCGOMF18, XW14]. focused [KBS05]. Focusing [Dry11]. fold [AJP16]. food [KG21, MA17, SMYH19]. footwear [Van20]. force [ZBF18]. forced [DFBL^{+}03]. forecast [FMI20, ST11]. Forecasting [AA12, AA20b, AAK10, AATD20, ES16, IS21, MRSR10, OO20, RORGTR20, AMA13, AC20, BM21, FMM19, MJ11, Muz12, SC19a, TW20, Tin18]. Forecasts [Thi09, APL19, GG07, Sal18]. foreign [AA11, AATD20, KG21, YT13, YM12]. forging [BMS05]. form [ESESIA17, GÖ01, OAA19]. formation [Van20]. forms [CBHM14]. formula [SPA18]. formulas [AS17, Vrb17b]. formulation [KB11]. four [GS14, MKM12]. Fourth [EAM10]. Fourth-order [EAM10]. fraction [AAAJ17]. fractional [AMO20, AA20b, Cho07, EKK04, EK04, IS10]. frailty [DAL06, WPJC21a, WPJC21b]. framework [DT07, GS08, JCACTIGO11, OO20]. Fréchet [KA19]. free [FCG10]. freedoms [IM09]. fresh [uaSYHNQ20]. Freund [Cho07]. frontier [Das20]. fully [KHY10]. function [AKGU^{+}16, AG07, Al 16, Ald20a, Dia07, ES15, GKKTS14, JZ20, Kip20, LL09, NA05, Niw17, OAA19, RA21, RAM15, Sae17, SJ12, Vrb17b, ZAJ20, ZBF18, Zgh03]. Functional [ZLZ06, El 11, GZS18, SS16]. Functional-Coefficient [ZLZ06]. functions [AASI17, Ald21, AF08, Beg11, CBHM14, DA10b, HM14, Hsi08, Kub19, Ram16a, SA10, Wan10]. fund [LYZ15, ZP08]. future [Abo10]. Fuzzy [BV09, Thi08, Dha18, RJ20, Thi09]. FWER [Mun11].

G [YXY19]. G-M [YXY19]. GA [EW15]. GA-KNN [EW15]. GA-SVM [EW15]. game [Ros20]. gamma [CZ12, CBSBHM10, DAL06, GP10, Ham20a, MESAF16, OMO19, Ras13, SAGKS17, WPJC21a]. gamma-distributed [CZ12]. gap [MA17]. gaps [Tor20]. GARCH [BB18, Moh13, RORGTR20, SSL07, Yel20]. GARCH-jump [SSL07]. Gauss [ASSSAZ18]. Gaussian [AK06, Gar08, KAA08, Ven15, ZZN13]. GDP [BA19]. GEE [MRS19]. gender [Mah10]. gene [APD^{+}10]. General [Vrb17b, AE11, EE14, MEL16, PCB13]. generalization [AAE12, EE20b, Um05]. Generalized [AASI17, AASI18, AEB13, Alk18, Gun12, NLI16, Niw17, PFL06, QH18, YMT13, Akt15, AMAAEA19, Al 16, AR14, BA20, BA15, ÇA12, CTF11, DC20, FEM^{+}09, FS03, Fam08, FKH16, FKL06, GAM03, GAAAA04, GP10, GA09b, Ham20a, Has20, Hsi08, HSL09, Hür03, JSS19, KAF12, LB16, Lug16, Mos15, NMN18, Par10, RIH20, Ras13, RGlHnSC20, SdA12, SK08, SS19, Sha19, SSL07, SNaN14, SH09, TTM08, TPP07b, TPP07a, TC05, WM08, WDM14, WRP16, ZAJ20]. generated [Han03, HE16, KR01, SD21]. generated-Dagum [SD21]. generating [SA10]. generation [JKS21a, JKS21b, SOAD06]. Generators [MKO21, KLK04]. genes [PM21]. genetic [NM04, PD07, SOAD06, SM12b]. Gensini [KT02]. genuine [Coz08]. genus [AB17]. geochemistry [AOAS17]. Geometric [Ven15, AMO20, AA20b, AA20c, AAAA18, GAM03, IS10, YMT10]. geometry [BKK13]. germination [KM08, OKM16]. Ghana [BTANKB19]. given [Fel11]. glitter [KRS09]. global [Bui20b, Ram16a]. GLS [RYAD12]. Gmm [Das20, Doa20a, Doa20b, RYAD12]. goal [ERH17, HAE19]. gold [KRS09]. goodness [AR16, KNG11, RE18]. goodness-of-fit [KNG11]. goods [Doa20a]. Google [FMI20]. governance [AM12a]. government [BA19]. Governorate [MA20]. grades [Mah10]. Granny [DZS15]. graphic [ZW10]. graphic-based [ZW10]. graphical [Gar15, Mek10]. gravitational [BKK13]. gravity [YM12]. Great [Azi16]. grid [BMS^{+}19]. grid-based [BMS^{+}19]. Gröbner [EKK03]. group [EB18, KA20, MEL16]. grouped [Das09]. groups [JM19, Lee01]. growth [AM15b, AM15a, Ami14, DMLN10, DH20, KG21, SMHN19, YT13]. Gulf [YM12]. Gunung [AKGU^{+}16].

Haar [FKE16]. half [AKEE18, Fay19, NMN18]. Harmonic [YMT10, Azi18]. Harnessing [Zak20]. harvesting [UMV13]. Haryana [VPKG12, VPO^{+}12, VPG^{+}14]. having [AS13, ES15, Ham20b, TH02]. hazard [AE17, Afi17, AK06, DSG^{+}09, ES15, Wan10]. hazards [Mos15]. health [AK12, AZA^{+}19, ACM07, FAB20, PD07]. Healthcare [lÇl19, GKKTS14]. heavy [FH14]. hedging [Chi14, ZP08]. Hellinger [Gar17]. helpful [Kha13]. hemodialysis [CJ20]. Heterogeneity [YQCSG10]. heterogeneous [LYK03, UH13]. Heteroscedastic [DMLN10]. heteroscedasticity [GA09b, LW04]. hidden [DGG05, GPP12, IM15]. hierarchical [BS21, DT07, Moh14, ZHMK09]. hierarchy [Maz04]. high [FH14, PvR04]. high-dimensional [PvR04]. higher [DLM11, JCACTIGO11, JCGO15]. HIV [UABW21, DDJM09]. HIV/AIDS [DDJM09]. HNBUE [SA10]. Hodgkin [PKKHM^{+}16]. holdings [Doa20a]. Holt [Azi20]. home [Han11]. homes [ZHMK09]. hometown [SSBA20]. homogeneous [Bar12, DDJM09, DN10, Kon13, Wan13]. Hong [Hoa20]. hospital [Con06]. hospitals [ES12, EAHH13]. hot [BMS05]. Hotat [MA20]. Hourly [SC19a]. HRFS [BMZ^{+}17]. human [AM12a, MRSR10, UABW21]. Hybrid [AMA13, AAED13]. Hybridization [Afi17, Had19]. hydraulic [DRW19]. hydrology [ZLZ06]. hyperbolastic [GKKTS14]. hyperbolic [WDGP10]. hypergeometric [ASSSAZ18, NA05]. hypotheses [DN17, PCSMGO11]. hypothesis [dA10a].

Iceland [TRCH12]. ICT [SA16]. ideal [ORMPL17, PDK06]. identifiable [Nin08]. Identification [IWC03, IAED15, uaSYHNQ20, DN17, IE05, Víš14]. identify [AAN21]. identifying [AA15b, DNH20, YB13]. II [AM20, BA15, CBSBHM10, ES15, EH12, FMA19, LS07, SJ12, SJ10, She20]. III [Sal19a, GO15]. image [Gri07, LYK03]. images [SM12b]. imbalanced [CTM18, JM19]. immigrant [AM12b]. immigration [Dôk06]. Immune [UABW21]. immunohistochemical [PKKHM^{+}16]. Impact [DKS19, Nko10, VPG^{+}14, AE17, Chi14, Din18, KG21, YT13]. imperfect [Gas12]. implication [dA10a]. implicit [KL15]. import [IS21]. importance [MN20, WY14, YS17]. imports [Ald13]. improve [Azi20, Dry11]. Improved [Kar20, TMO07, BSGOFGPF17, MME21, TTT08]. Improving [Vrb14b, AKGU^{+}16]. Imputation [Won20]. Inadmissibility [ŞS10]. incapability [Lin02]. incidence [Akt15, BV09]. income [Abd04, HLBF02, OO20, SK08]. Incomplete [KFT13, AGM^{+}07, BV21, MKM12, SB20, Tol12, Vic08]. inconsistent [YSK06]. increasing [Li15]. Independence [WS15a, CKC08, CM13, TMS08, Vrb14b]. independent [AK12, CJ12, DS08, Kaw10, Kaw14, KY15, LPKPS10, Um08, YKK03]. Index [AA20b, AEEG19, AM12a, AM15b, AF08, CMT16, CBHMRA17, Din18, EE14, HLZ06, KT02, Lin02, SA16, SJH18, Zah14]. indexed [Yos15]. indexes [Dor05]. India [VPKG12, VPO^{+}12, VPG^{+}14, KG21]. indicators [AM12a]. indices [Hun05, KPS08, SAM19, SV04, Stu02]. individual [CZ12, EAHH13, Lee01, NL21]. individuals [LB16]. Indonesia [AKGU^{+}16, SPM20, Tin18]. industrial [Azi16, Ela10]. industry [ARA^{+}18, ARA^{+}19, KG21, Van20]. inequalities [Hür15]. inequality [AEH18, AMB11]. inertia [KKMMK20]. infarction [EAHH13]. Inference [ZX05, AM20, BA20, BLV06, DRM15, FMA19, Gas12, GOdlFSLF08, HSL09, RYS08, She20, Tan10, WY14]. Inferences [HLZ06, AM10b, Gun12, YT10]. Inferiority [KY15, AACARC13, AARCCA16, Mun11, SRCTAA13]. infinite [MQLA19]. inflated [AZA^{+}19, FS03, MFR^{+}19, PC13]. inflation [AC20, Din19]. inflows [YM12]. Influence [MH12, Bui20b, GOSLVA12, QKH18, uaSYHNQ20]. Influencing [Van20, DLP12, SHM19]. influential [Sha08]. information [EJ04, GD10, IKA^{+}20, Kel05, Lee01, MSIT05, PD07, SA16, Tol12, Zur01]. infrequent [Arr13]. initial [Wu14]. injection [Alt18]. injury [ACG10]. input [Ven15]. instantaneous [YXY19]. institute [RIT18]. institutions [JCACTIGO11]. insurance [AMO21, LLDD17, PHTY07, Sim06, WPJC21b, XDL16, YSK06, ZBF18, Zim08]. integer [PSA18]. integer-valued [PSA18]. integral [FKE16]. integrals [Gar08]. integrated [AS17, APL20, Tin18, YXY19, Zoi09]. integrating [Alt18]. integration [Ald13]. intelligence [El 18]. Intention [DB20]. interest [CY04]. Internet [Han09]. interplay [AME18]. interpolation [GBHP^{+}12, Ram16a]. interruptions [AS13]. Interval [Li15, AK12, AK19, AA19, BK02, CaN16, FU19, Kaw14, LL09, QTH04, Sae17, SD19, Um05]. intervals [AA20a, Cha03, HYZ13, IM15, Kaw10, Laa12, NKaN11, NW16, aN16, Niw17, PCSMGO11, RW18, RCAAMC19, SV04, TNN17a, TNN17b, TN18, TNN18, WaNN14]. Intervention [Kam19, YB13]. intra [BV21, YT10]. intra-class [BV21]. intra-cluster [YT10]. intrinsically [Sin16b]. introducing [AZIE19]. Introduction [Laa14]. Inverse [SRMS17, AR16, AK06, AM20, AOM17, BF16, KA20, NMW20, NW16, OMO19, SS19, TNN17b, WaNN14]. inverted [BA15, DA09]. investigate [AAM^{+}13, MA20]. Investigating [DRW19]. Investigation [AGW19]. investment [AA11, KG21, YT13, YM12]. investments [ZP08]. investor [Chi14]. Iranian [SMYH19]. irregular [PMS15]. Issues [Mun11]. Italy [AM12b]. item [FCG10]. Iterative [Akm05, MN20]. Iteratively [Ame16]. ivar [PMS15].

jackknife [GCN20, MEL16]. Jazan [JZTS17]. Jeffery [IuZHZ20]. Jenkins [TS10]. job [RIT18]. Johnson [SRCTAA13, Yan08]. Joint [AH21, HEDA18, Has18, QKH18, SGM20, WPJC21b]. joint-life [WPJC21b]. Jordan [BAZ20]. Judge [BRY13]. judgment [GO15, Kou02]. judicial [BRY13]. jump [Abu02, SSL07, YY20]. jump-diffusion [Abu02]. juvenile [DBD^{+}07].

Kalman [AB11, DGW^{+}10, DHY^{+}15]. kappa [Um05]. Katz [Sto06a]. Kendall [Wan13]. Kernel [YY20, EAM10, HZFZ06, JD16, KA06, KKD09, SP17, Yan12]. key [Van20]. Kidul [AKGU^{+}16]. kind [Bak03]. knit [RR03]. KNN [EW15]. knowing [YT10]. known [MS06, NKaN11, aN16, SNaN14, WaNN14]. Kotler [SRCTAA13]. Kubbard [DMLN10]. Kullback [JGAF10]. Kumaraswamy [AMA20, ASM17, AOM17, EE19, FMA19, HE16, NMW20]. kurtosis [Wu02]. Kuwait [AAKAK11].

L [Sto08, Ham20b]. labor [LP11, Slo13]. labour [LVM11, Nko10]. LAD [Oma18]. lag [Ald13, ÇA12, FDdPF10, NM04]. lagged [Mar13]. Lagunera [MRSR10]. Lai [Sto08]. Lake [AAM^{+}13]. lambda [Kvå11]. LAN [Dog14, Kon13]. Land [SWCT19]. landings [RSR^{+}18]. language [AA08]. Lanka [APL20, DKS19, HCJ21, JKS21a, JKS21b, SC19a, WH20]. Laplace [AMEGM12, AS18, AAE12, JK03, RE18]. large [ABD05, DGHP10, SZZ08, Sto11, TNN17b, TP02]. large-sample [DGHP10]. larger [LCL02]. larger-the-better [LCL02]. Laster [SRCTAA13]. latent [KHY10, MHW^{+}19, TW13]. law [CBHMRA17, Hür03, SZZ08, Sto06a, Sto11]. laws [JS03]. layout [Abd04, KBS05]. learning [Alt18, DT19, GF19, LTTD20, XW14]. Least [Kou02, AKA20, AKAA20, FH14, JZTS17, KT08, ŞS10, Víš15]. Least-squares [Kou02]. leather [Van20]. left [ORMPL17]. Leibler [JGAF10]. lemma [BL01, PKKHM^{+}16]. Length [AK06, Akm05, AS16b, AS17, BAS17, EE20a, FDdPF10, Ras13, SSA16, SPA18]. Length-biased [AK06, Akm05, EE20a, Ras13]. Leone [AM20, SD21]. level [BDNN15, EKK03, GS03, MKN16, Sal05, Tor21, VPO^{+}12]. levels [AARCCA16, SR14b, WDM14]. Leveraging [AAR19]. Lévy [DNO18]. lexical [Zha20]. Libby [RIH20]. life [BK12, Bak03, Dia07, ES15, Ham20b, LS07, Ram16b, SAM19, WPJC21a, WPJC21b, YSK06, ZAJ20]. lifetime [AA20a, AAED13, BA14, FG10, KA20, Li15, PAL11, RMdSC13, SLCB11]. lifetimes [WM08]. light [CCM11, DB20]. light-models [CCM11]. lightning [KKBB18]. Likelihood [BLV06, YT10, AS19, BS21, DGHP10, FCG10, GCN20, HM10, KSM16, Ma05, MN20, MSIT05, PC13, QTH04, RM18, Vrb17b, Wei15, ZX05]. Likelihood-based [BLV06, BS21]. likelihoods [TH14]. limit [Dôk06, KR01, SZZ08]. Limiting [BLLR01]. limits [SR14c]. Lindley [Alk18, AA20c, AS19, DC20, JSS19, Kar20, Sha19]. line [AAH18, EAM10]. Linear [LV18, LVX19, AMAHE12, AASI17, AASI18, Akt15, AE11, AAE12, Ald20a, AAE11, BU16, Beg11, Cha03, FLMD11, FDGP13, GR15, HCJ21, Han03, HTD18, JM19, KR01, KS03, LCP05, LPLH08, LTvR09, MS06, MBSJ19, Özt01, PG10, PDK06, SRMS17, Sae17, Sha10, SBZ03, TCD20, Tol12, Tor21, TN14, TC05, WDM14, WS15b, Wu16, Wu17, ZLZ06]. linearity [Mak14]. Linex [GNB01]. link [AASI17]. listed [DLP12, Moh14]. liu [EFAEH17, Oma19, ŞS10, WS15b, Wu16]. liu-type [EFAEH17, Oma19]. liver [Azi19]. load [BMS05]. loading [Zim08]. loans [LX12]. Local [AF08, LM11, TW13, Zur01]. localization [BMZ^{+}17]. location [MS06, MPZ09, Thi08, Tza08, Zgh03]. location-scale [Thi08]. loci [FFX01]. Log [ASM17, AAED13, GP10, GR15, KB11, MNC17, TTT08]. log-compressed [MNC17]. log-density [KB11]. Log-Exponentiated [ASM17]. log-gamma [GP10]. log-linear [GR15]. log-logistic [AAED13]. log-odds [TTT08]. logistic [Abd10, AKEE18, AATD20, AJP16, AAED13, BA20, BR12, DNH20, DT19, FEM^{+}09, Fay19, GD14, GÖ01, Mah10, MA20, NMN18, TP02, XB10]. logistics [uaSYHNQ20]. logit [Das09]. lognormal [AK12, GYH18, Niw17, SNaN14]. Logone [AAY^{+}19]. Lomax [Ald21, Ess17, Fay19, RMdSC13, Sal19a, SAGKS17]. long [AMO21, CTF11, El 18, GAAAA04]. long-term [CTF11, El 18, GAAAA04]. longevity [Lev13]. Longitudinal [Wan14, AEEG19, BU16, BV09, DGM20, DBSH11, SGM20]. Lorenz [Fel11, RAM15]. loss [Ald20a, Ald21, FL09, MED14, SJ12, Sun11]. lot [AAAJ17, GS03]. lot-size [GS03]. lower [Dry11]. LTRC [BA04]. Lugannani [Tsa02]. lymphoma [PKKHM^{+}16].

M [ASS10, ASS10, YXY19]. M/M/1 [ASS10]. Machine [GF19, Alt18, DT19, KA06, LTTD20]. magnesium [FEM^{+}09]. magnitude [GK08]. magnitude-squared [GK08]. Mai [SWCT19, TS10]. main [Ald13, Kam11, SR14a]. maintenance [Ald20b, Gas12, GB07]. makers [PSKF17]. making [GS08, PSKF17]. Malaysia [RSR^{+}18]. Mali [LLDD17]. Malmquist [AM15b]. malus [LLDD17]. Management [WYZ10, ACG10, HCJ21, Zim08]. Mandel [SdA12]. Manipulating [CCM11]. Mann [Kel06]. Manning [AASR17]. manufacturing [PG10]. many [Al 20, Arr13, PC14]. Maple [DGS08]. mapping [Zak20]. Marginal [YKT14, FCG10, YMT13]. MARIMA [AMA13]. marine [RSR^{+}18]. marked [Has18]. markers [PKKHM^{+}16]. market [Azi18, Bui20b, CBHMRA17, EE14, LP11, LVM11, Moh13, Yel20, ZBF18]. Markov [BRR11, CEG09, DDJM09, DN17, DGG05, DLM11, GS08, IM15, MRSR10, MQLA19, Vrb15]. Markov-model [Vrb15]. Markovian [AA08, GB07, RM14]. marks [Has18]. Marshall [Afi16, şTE^{+}11, CKC08, Cho08, EE20a]. martingales [WS15a]. mass [Sel15a]. Mathematical [FKH16, EFAEH17, EB18, KT02]. matrices [CBHMRA12, FLMD11, KM02]. matrix [FKE16, NBA04, SSRG16, TTM08, ZKZ13, ZZK13]. matrix-variate [NBA04, TTM08]. MAVE [AY13]. maximization [SJH18]. Maximum [DGHP10, HM10, RM18, Ros20, AH21, AS19, ÇA12, FCG10, KSM16, Ma05]. Maxwell [OAA19]. MCEM [DN10]. MCMC [AM10b]. mean [AMB11, BA14, Dha18, Dia07, DNO18, EJ04, FSH10, IKA^{+}20, JAO06, KK09, KA20, Kip11, LCP05, MDTD10, Nis09, QTH04, QKH18, SJ12, SNaN14, Swi08, TO05, TNN17a, TNN17b, Tsa02, WaNN14, Zam19]. Means [Zha20, AK12, BAS17, BAS19, CT01, DDJM09, HBJ09, Moh14, NKaN11, NW16, aN16, PC13, RW18, YMT10, YKK03]. Measure [BV21, IT18, Law12, Bor12, FG11, Hur12, LM11, Nis09, TSU01, TH02, Um08, Wu01, Wu02, YKT14]. Measurement [JKS21b, AM15b, CFFVGO10, Liu10, SS16, WS15b, Wu17]. measurements [AS05, CZ12, NL21]. measures [Arb08, BRIA17, BV09, DMV13, Hür05, ICBS14, Kvå11, RIH20, TMO07, YMT10, YMT13]. Measuring [AdC14, Dor05, YS17]. Median [RJ20, AS18, PCSMGO11, SMG12]. Median-Variance [RJ20]. medical [GS08, Mos15, Par10, RA21, SM12b]. medium [DNH20]. Melia [OKM16]. memory [AMO21]. merger [AA12]. meta [DT07, GZS18, PC18, SS21]. meta-analysis [DT07, GZS18, PC18, SS21]. metabolomics [AGW19]. method [AH21, AK08, AE11, AGW19, Azi20, BK02, DS08, Dog14, EAM10, FH14, FR17, GZS18, Hag19, IM15, JWLZ14, KAF12, NV20, OOA11, PK16, PSKF17, Ram16a, Ram16b, RS14, Sad10, SYO04, SD19, Sho08, Sin16b, SR14c, SH09, SCG01, TKD19, Tor21, TNR10, Tza08, WCW14, YB13, YS17, YH09, ZZN13]. methodologies [Cha12, SVSO02]. methodology [KPP09, OKM16]. methods [AK19, AA15a, AMAHE12, AM10b, AKGU^{+}16, AEB13, ÇA12, DBD^{+}07, EW15, ESS06, Fuj13, GOG07, Gri11, Hag14, HYZ13, JM19, KY15, KA06, Moh14, Mun11, SSBA20, Tan10, TH14, WNT20, WK18]. metric [AM07]. metrology [RA21]. MEWMA [FHS20]. Mexican [CBHMRA17, RORGTR20]. microarray [PM21]. migration [GOSLVA12, MRSR10]. minimize [Ald20b, FR17]. minimizing [XDL16]. Minimum [SJ10, AH21, BF16, MS06]. mining [AA11, Yİ10]. MIRNAS [AME18]. mis [LVX19]. mis-specified [LVX19]. misconceptions [FAB20]. missing [AK08, DGM20, ESESA16, IM20, TN14, Won20]. missingness [Liu10]. misspecified [KT08]. Mixed [ARAA17, SSL07, SBZ03, AA19, BU16, Beg11, Cha03, GOG07, JGAF10, Liu10, Ras15, SS21, STPS10, VPO^{+}12, Wan14, ZX05]. Mixed-effects [SBZ03, Liu10]. mixing [HZFZ06, SP17]. MIXMAX} [Alb12]. Mixture [BH08, Akm05, DT07, GAAAA04, Ham20a, Has18, HBJ09]. mixtures [şTE^{+}11, MWB17, MKO21, RM18]. MLE [Mar14, Vrb14a]. MLEs [Vrb13b, Vrb13a]. mobile [Han09]. model [AJM11, AMA13, AM15a, Abd10, AA15b, Afi16, AA19, Akt15, ASAF09, AAA^{+}16, Al 20, AA20a, Ald13, Ald21, AA20c, AAAA18, AS17, Are19, ABS08, AJP16, Azi16, AF19, Azi19, BK12, Bar09, Bar10, Bar12, BA14, Beg11, BTANKB19, BAS19, ÇA12, Cha20, CY04, CG07, CT01, CKC08, Cho08, CJ12, DDJM09, DGI18, Das20, DC20, DAL06, DT07, DMLN10, DNH20, Dry11, DNO18, ESESI14, ESESIA17, ESAAA18, EAHH13, EE14, Fam08, FKH16, FCG10, FLMD11, Gas12, GYH18, GAAAA04, GS03, GOSLVA12, GA09b, GAP21, HM10, HB17, HSC09, IM15, IS10, JSCI18, JCGO15, Kam19, KT02, KS03, KG21, KFT13, KSM16, Law12, Lee01, LP11, LP13, LNC10, Mah10, MME21, MH09, Maz04, MRS19, MQAA21, MSIT05, Mos15, MA17, NK13, OOA11, Oma19, PG10, PRVATS14]. model [PFL06, Par10, PACW13, PMS15, PLRV15, QDG17, RJ20, RS14, RYS08, SMHN19, SAM19, SC19a, Sha08, Sha10, SJH18, Sin16a, STPS10, Sto06b, SS18, SLCB11, TW16, TCD20, TW20, Tin18, TSU01, TMS08, Víš14, Vrb10, Vrb15, WPJC21b, Wan13, Wan14, WH20, WS15b, Wu16, WRP16, Wu17, XW19, XB10, YY20, You05, YM12, ZZN13, ZX05, lÇl19]. Modeling [AK16, AA08, AAK21, EMM21, FMM19, GAAAA04, GKKTS14, Yel20, Alt18, BMZ^{+}17, Has18, Moh13, NMW20, SGM20, Tor21, TH14, Wu14, Zak20, ZHMK09]. Modelling [Azi18, Con06, KM08, KRS09, LW04, AZ19, APL20, BMS05, KDNH18, MNC17, RSR^{+}18, Sim06, VPO^{+}12]. Models [ACG10, AEEG19, Abd04, AAA19, AZRS11, AASI17, AASI18, Akm05, AE11, AMO21, AAY^{+}19, AB11, AS16b, APL19, AAED13, AAS12, AA17, AGM^{+}07, BU16, BLB15, BB18, BAS17, CCM11, Cha03, CW10, DGHP10, DGG05, ESESI15, ESESA16, EM10, Els21, FS03, FMI20, FG10, FU19, FDdPF10, Gar15, GA09a, GOPLD09, GÖ01, GR15, Hag19, HCJ21, HLZ06, HA08, Has18, IAED15, IAS19, IM20, JK03, KHY10, Kip13, Lev13, LPLH08, Liu10, LK06, MY15, MBSJ19, MFR^{+}19, Moh13, MH12, Oma18, Özt01, PACW13, QF20, RZ14, Ras11, RG15, RSM^{+}11, Ros20, RORGTR20, SV12, SS16, SSL07, Sho08, ST11, SBZ03, Sin16b, SB11, TW13, Thi08, Thi09, TS10, Tol12, TN14, TC05, UH13, VPKG12, Vrb17b, WH17, Wei15, Yeh16, Yel20, ZLZ06]. modern [SMYH19]. Modification [Oma19, EE17, JD16, LC01]. Modified [AR16, AZA^{+}19, EJ04, GCN20, VVV06, CTF11, Vic08, WDM14, WRP16]. molding [Alt18]. moment [EOB07, SA10, Sto06a]. Moments [Vrb10, ZKZ13, AM10a, ASSS20, JSS19, MDTD10, Sel15b, Sin16b, Tza08]. monetary [DH20]. money [Din19]. monitoring [Dha18, Dog14, MRS19, PSA18, WZG11]. Monte [AAA19, AKAA20, BRR11, HBA04, RYAD12]. monthly [Cha20]. mortality [AK16, GD14, HLBF02, KHY10, Lev13, LTTD20, WNB13, ZBF18, ZP08]. mortgage [AMO21, LX12]. mother [AZ19]. motion [AMO20, AA20b, GR13, IS10, Yos15]. Motor [MFPO20]. movement [Nwa09]. movements [RMS^{+}11]. moving [AS16b, AS17, AKA17, APL20, BAS17, HK20, IAED15, IAS19, SPA18, Tin18]. MPI [DA09]. MR3060565 [Vrb13b]. MRNAS [AME18]. MS [AGW19]. MSE [MS06]. Muktiharjo [SPM20]. Multi [VPO^{+}12, CTM18, EJ04, EB18, KNMM13, PK16, YKT14]. multi-auxiliary [EJ04]. multi-block [KNMM13]. multi-group [EB18]. Multi-level [VPO^{+}12]. multi-objective [PK16]. multi-way [YKT14]. multicollinearity [KT08, QF20]. Multidimensional [BAZ20]. Multilevel [AZ19, AAS12]. multimodal [LWZ^{+}17]. multinomial [BB06, uaSYHNQ20]. multiobjective [HTD18]. Multiple [Cho07, ARA^{+}18, ARA^{+}19, Al 20, Ami09, AKA17, DN17, KW02, KG21, LCP05, LNC10, Nis09, PPT02, PACW13, PSKF17, Tor21, TC05, WY14]. multiplicative [JOO18]. multipoint [Das09]. multisample [NC03, OO07]. Multistage [JAO07]. Multivariate [ASAF09, GD10, TPP07b, TPP07a, AZRS11, AMEGM12, Arb08, Beg11, CMT16, DGS08, FSH10, Haj08, LCP05, PMS15, TK04, Um05, Um08, WH17, ZW10]. mutual [GD10]. myocardial [EAHH13].

Nadaraya [JD16]. Nakagami [MNC17]. names [AA08]. Nasser [AAM^{+}13]. nationwide [YQCSG10]. natural [LT20]. navigation [Fuj13, YAA16]. NBU [LKP04]. NBU-ness [LKP04]. Negative [APL20, PCB13, RCDO11]. neighborhood [SSBA20]. neighborhoods [SYO04]. neoclassical [Slo13]. neonatal [AK16]. ness [LKP04]. nested [Akt15, FR17]. network [BF16, LYK03, Moh14, SRMS17, SC19a]. network-based [LYK03]. networks [CTM18, IM09, LK06, Moh13, PvR04]. Neural [LYK03, LK06, CTM18, IM09, Moh13, Moh14, SC19a]. neutrosophic [PSKF17]. Neyman [BL01, PKKHM^{+}16]. NHANES [GO15]. NHPP [Ami14, Bar10]. Nigeria [OO20, SOO08]. Nigerian [AOAS17]. noise [CL06, Cos09, Gar08, LHC01, TK04]. noisy [AS05]. nominal [AARCCA16, Um08, YMT10, YMT13, YT10]. Non [BDNN15, KY15, NBA04, Sha19, Sto06b, WPJC21a, AAM^{+}13, AA15a, Ald20a, AACARC13, AARCCA16, Are19, Bar10, Bar12, DDJM09, DN10, EW15, FLMD11, GOG07, GA09a, Hag19, Kon13, LV18, LCL02, LK06, LTvR09, Mun11, Sae17, SRCTAA13, SR14c, Ven15, Won20, Yan08]. Non-central [NBA04, WPJC21a]. non-homogeneous [Bar12, DDJM09, DN10, Kon13]. Non-Inferiority [KY15, AACARC13, AARCCA16, Mun11, SRCTAA13]. non-linear [Ald20a, LTvR09]. non-normal [LV18, LCL02, Yan08]. non-normality [Sae17]. non-orthogonality [FLMD11]. Non-Parametric [BDNN15, Sto06b, AAM^{+}13, AA15a, Bar10, Bar12, EW15, GOG07, GA09a, Hag19, LK06]. non-separable [Ven15]. non-stationary [Won20]. noncentral [HP04, Hua06]. nonclassicality [SdA12]. nonconforming [AAAJ17]. noninformative [Moa10]. nonlinear [AB01, Ald21, CKS01, FKE16, JWLZ14, LYK03, Liu10, MN18, SSRG16, Sin16a, Sin16b, STPS10, WJ13, XW14]. nonlinearities [EI07]. nonmonotone [JWLZ14]. nonnegative [DS08]. Nonparametric [Aly16, ACM07, BAAA02, BA04, LM11, CW08, El 11, FEW09, HZFZ06, RZ14, SP17]. Norm [HR07, CKS01]. normal [AJM11, AZIE19, AAE12, FSH10, Han11, Hsi10a, KK09, KNG11, LV18, LCL02, Lin02, LCP05, Mag19, NR05, NKaN11, NW16, aN16, RM18, SC19b, TNN17b, WaNN14, Yan08, YKK03]. normal-Laplace [AAE12]. normal-normal [AJM11]. normality [CBHMRA11, El 11, Sad10, Sae17, Sho08]. normalized [Sto11, Ven15]. North [AF08]. Northern [VPO^{+}12, Van20]. Note [UH13, Che09, Cos09, FMTW10, GR13, KSM16, LP13, LPLH08, LC09, MSIT05, NR05, Pol19, RP15, Ram16b, Sha08, SLKP08, Wan13, WJ13, Wu16]. novel [CT01, KB11]. Novick [RIH20]. null [Bor12]. number [FM10, HB17, KLK04, KPS08, MED14, TW20]. numbers [PSKF17, RJ20, SZZ08, Sto11]. Numerical [FKE16, CEG09, DS08, KAK18, Sho08]. nursing [ZHMK09]. Nuwara [WH20].

O [MS06]. O-BLUE [MS06]. Obesity [HSC09, YQCSG10]. objective [PK16]. observable [GS08]. observation [Els21]. observations [AK08, Che09, IM20, Mad08, Sha08, SR14a]. observers [Um05]. obtain [AM10b]. occasions [LC09]. occidental [AAY^{+}19]. occurrence [AA15b, EAHH13, WNT20]. occurrences [KDNH18]. occurring [HB17]. oceanology [RA21]. Odd [Lug16, AOM17, Sal19a, SMG12]. odds [TTT08, XB10, YQCSG10]. offense [DBD^{+}07]. official [AAR19]. offset [AZIE19, Fuj13]. oil [AOAS17, Din18, SV12]. Olkin [Afi16, şTE^{+}11, CKC08, Cho08, EE20a]. one [Ram16a, TNN17a, TN18, TNN18]. one-dimensional [Ram16a]. one-parameter [TNN17a, TN18, TNN18]. openness [Bui20a]. operation [SV12]. operational [FKE16, LYZ15]. opinions [SHM19]. Optimal [Beg11, CKS01, GLARM18, KBS05, LS07, MY15, TP02, YSK06, Zim08, Ald20b, BLB10, DP08, Das09, HCJ21, KM02, PvR04, PKKHM^{+}16, Sha10, TNR10, UMV13]. Optimization [OKM16, Gri07, KS18, PK16, Ram16a]. Optimizing [Gar15, AE11]. Optimum [GAP21, HAE19, She20, KMT12]. option [FKL06, Muz12]. options [Lin09]. oracle [AY13]. order [ASSS20, Al 16, Ald13, BV21, BMS^{+}19, CBHMRA12, CBHM14, DP08, DLM11, Dry11, EAM10, GS03, JSS19, KMT12, MKO21, MA19, MKM12, MKN16, PK17, SS19, SB20, TPP07b, VVV06, Vic08, Wan07, Yeh16, Zoi09]. ordered [Aly16, IT18, Kou02, KFT13, TH02, TMS08]. ordering [MA19, RAM15, YKK03]. orderings [OT07, WM08]. ordinary [ŞS10, Zim08]. organizations [GLARM18, XW19]. Ornstein [AMO20, DNO18]. Orthogonal [EOB07, AKGU^{+}16]. orthogonality [FLMD11]. oscillators [AB01]. other [DMV13]. outcomes [FEW09, PD07, SB06]. Outlier [MBSJ19, YMM11]. outliers [AA20a, SKS08]. over-dispersed [Has20]. overall [ORMPL17]. overlap [ERH17]. overweight [YQCSG10]. overweight-obesity [YQCSG10].

p [Che02, Ham20b]. P-A-L [Ham20b]. package [XB10]. pair [SB20]. pairing [GS14]. Pairwise [KW02, VVV06]. Pakistan [BA19, RIT18]. pandemic [RORGTR20]. panel [AAA19, AA12, ESESI14, ESAAA18, IM20, YM12]. paradigm [AAK21, OO07]. Parameter [AB11, CMT16, JPW11, LP11, AEEG19, Afi16, AZIE19, Alt18, AS19, BL01, BKK13, CT01, DC20, EOB07, FKH16, Gun12, IuZHZ20, KAF12, KA20, Kip20, KFT13, Mad08, MS06, MME21, MPZ09, PAL11, RM18, SdA12, Sin16b, Tan10, TNN17a, TN18, TNN18, Tza08, Vrb10, XW14]. Parameters [BA15, KE19, AJM11, AYM15, AKAA20, Ald20a, AS19, AGM^{+}07, DRW19, DÇ18, ES15, HEDA18, HA08, IM15, JZ20, Kou02, Li11, Li12, LP13, Moa10, NV20, Özt01, RS14, Ras15, SAGKS17, SD19, Sin16a, Thi08, Tor21, TC05, UH13, Vrb15, WZG11, Zgh03]. Parametric [AAM^{+}13, BDNN15, TC05, AA15a, Abd08, Bar10, Bar12, Beg11, BK02, CC08, EW15, GOG07, GA09a, Hag19, JOO18, LK06, MN15, Sto06b, SR14c, TH14]. parametrized [Kon13]. Pareto [AYM15, Bak03, Gun12, HZ10, Kam11, Kou02, PHTY07, Sal19b, WDM14]. parity [Muz12]. Parsimonious [BTANKB19]. Partial [Lop14, AMB11, JZTS17, KSM16, Wu16]. partially [GS08, KHY10, TCD20, ZLZ06]. Particle [XW14, ZRX10]. partitions [KPS08]. past [BA14]. path [IFVA06, lÇl19]. path-distances [IFVA06]. Patient [DDJM09]. patients [CJ20, FAB20, NBL15, STPS10]. Pattern [Nwa09, Nin08]. patterns [AM12b, AKÜ20, Kip11]. pavement [BK12, LK06]. payments [GS03]. PBIB [GS14]. peakedness [Wu02]. Pearson [BL01, PKKHM^{+}16, SRCT17]. penalties [AY13]. Penicillium [AB17]. pension [LYZ15, MQAA21, ZP08]. percentile [JAO06, JAO07]. Percentiles [JZ20]. perception [GOPLD09]. perceptron [Ven15].

Performance [BK02, AAK21, AA20b, Azi20, BSGOFGPF17, Dha18, FMI20, Fuj13, HSC09, LHC01, NL21, SMP21, SPM20, SR14c]. performances [FU19]. period [AAN21, AH21, Ald20b, OA09]. Periodic [WH20, BH08]. periodicals [Arr13]. permissible [GS03]. Permutation [Sal05, Arb08, OO07]. persistence [BB18]. person [Lee01]. perspective [BV09]. perturbed [AMO20]. peso [RORGTR20]. pharmacokinetics [BLB15]. phase [KM08]. Philippine [AC20]. Physical [Hun05]. Pinto [DMV13]. pipes [RMS^{+}11]. plan [ARAA17, AAAJ17, Ham20b, KA20, MQAA21, SB09]. planning [HCJ21, Ram16b]. plans [LS07]. plant [ARA^{+}18, ARA^{+}19, KBS05, TRA^{+}18]. playing [Han11]. plethora [AME18]. PLS [lÇl19]. plus [CL06]. point [AKGU^{+}16, AAA^{+}16, BAAA02, BuH04, BA04, CG07, Has18, PRVATS14, YKT14, Yİ10]. point-symmetry [YKT14]. points [BL01, Cha12, Cha18, Dia07, Sel15a]. Poisson [AZA^{+}19, AAA^{+}16, AAK21, Bar12, BLF11, BAS17, BAS19, FS03, Fam08, Has18, Has20, Hsi10a, Hsi10b, IS10, JSCI18, Kar20, Kip20, MME21, MFR^{+}19, MM02, OMO19, PFL06, Par10, PSA18, RMdSC13, RYS08, SAGKS17, SS21, SPA18, Tan10, Yİ10]. policies [Fel11, Fel14]. policy [DH20, GB07, YSK06]. policyholder [YSK06]. political [DFBL^{+}03]. pollution [AZRS11, Zak20]. polykays [DGS08]. polynomial [FM10, Kel06, MY15]. polynomials [FG11]. pooling [ACM07]. population [AK19, BLB10, BLB15, EJ04, HBJ09, IKA^{+}20, JAO06, Kip17, LC09, QH18, QKH18, RZ14, SMG12, Swi08, UMV13, YQCSG10, Zam19]. population-based [YQCSG10]. populations [Cho07, GOdlFSLF08, Gun12, LCL02, PC13, TNN17a, TNN17b, TN18]. Portfolio [KS18, SSRG16, RJ20, ZP08]. positive [PK17]. possibilities [Nko09]. possible [Tor20, dA10a]. post [GO15, Wan14]. post-stratified [GO15]. post-transplant [Wan14]. Potato [Sal18]. potential [Zak20]. pound [ES16]. poverty [AAH18, BV09, Nko09]. Power [AA20c, TK04, TSU01, TH02, Al 16, Alk18, ABD05, ASM17, BU16, CBHMRA17, DGI18, ES15, Ess17, Fay19, JSS19, JZ20, JKS21a, JKS21b, Muz12, RA21, RAM15, YMM11, ZAJ20]. Power-divergence [TSU01, TH02]. power-law [CBHMRA17]. power/sample [BU16]. practical [Arr13, GZS18, KMT12, WCW14]. practice [Con06]. pre [AGW19, SJ12]. pre-test [SJ12]. pre-treatment [AGW19]. precipitation [TPP05, WNT20]. predict [MA17, SA16, SSBA20]. predicted [BM21]. predicting [AAY^{+}19, BTANKB19, TW20]. Prediction [AK08, AR14, PACW13, AM15a, Abo10, ASAF09, AA20a, Cha20, CaN16, DBD^{+}07, El 18, EE19, HYZ13, JCGO15, LTTD20, SR14c, TP02, VPG^{+}14]. Prediction-based [PACW13]. Predictive [RSR^{+}18, AK06, WH17]. predictor [WY14]. predictors [IE05, JZTS17]. preference [ACM07]. preferences [AC19, YSK06]. pregnancy [AZ19]. pregnant [MHW^{+}19]. Preliminary [MPZ09, BuH04]. prepayment [LX12]. presence [AM10b, AA20a, DA10b, GAAAA04, PC14]. preserving [GD10]. pressure [ZHMK09]. prevalence [GOdlFSLF08, HSC09]. preventive [Ald20b]. price [Cha20, CMT16, Din18, FMI20, Sal18, SMYH19]. prices [DRM15, Din18, FKL06, MQLA19]. pricing [ZZN13]. primary [Abd10, Con06]. principal [AIR18, LTvR09]. principle [Kam11]. prior [AZ19, IuZHZ20, Moa10]. priori [SB11]. priors [SAGKS17]. private [ES12, Hoa20]. probabilistic [CTM18]. probabilities [Abd04, CZ12, Kip11, LPKPS10, Ros20]. probability [CaN16, CF05, DS08, Gar15, IT18, KAK18, MN15, ND17, PC14, RP15, Tsa02]. probit [Azi16]. problem [BF16, CTM18, GÖ01, JGAF10, UMV13, UH13]. problems [CEG09, CY04, DN17, JWLZ14, PK16, ST11, ZRX10]. procedure [GPP12, Mek10, SSA16]. procedures [PCGOMF18, VPKG12]. Process [Lin02, SAM19, AMO20, Bar12, CaN16, Dha18, DNO18, Han03, HB17, Has18, IS10, JCACTIGO11, KBS05, KR01, Kip20, MM02, NL21, SPA18, Swi08, XW19, Yİ10]. process-focused [KBS05]. processes [AA12, CL06, Gar15, GP10, GS08, KAA08, KR01, LCL02, Lin02, MRSR10, El 20, PSA18, Pol19, SP17, Wan07, WM08, Zoi09]. processing [KG21]. Procrustes [LB16]. product [HH02, KAA08, LC09, MSP^{+}13, QKH18]. Production [PG10, BTANKB19, DKS19, FMM19, SMYH19, Tin18]. productivity [AM15b]. products [AAED13, Sim06, YKK03]. professionals [FAB20]. profiles [AB17, HSL09]. Profitability [Doa20b, HK20, AF19, Din20]. profits [AAN21]. programming [Ald20a, Ald21, EFAEH17, ERH17, EB18, FKH16, HEDA18, HCJ21, HAE19, HTD18, SRMS17, SSRG16]. progressive [AA19, BA15, LS07, SJ12, SD19]. progressively [ES15, SJ10]. projection [EOB07]. projects [PLRaB18]. proof [Bor12]. propensity [LX12]. Properties [Hür05, NA05, APD^{+}10, AMA20, AKEE18, BA20, BA14, ESAAA18, EE20a, EE20b, Fay19, Fel14, Gar17, KA19, MA19, ND17, PCB13, SD21, Sal19b, TCD20, TP02, You05]. Propolis [AGW19]. proportion [AK16, AARCMC20, RCAAMC19]. proportional [Afi17, Mos15, XB10, Zim08]. proportions [Kaw10, Kaw14, KY15, SRCTAA13]. Proposal [KKMMK20]. Proposed [AA15b, EAHH13]. prostate [MKM^{+}15]. proteomic [DCKD13]. prove [Kho08]. provide [Maz04]. Province [SWCT19, TS10]. provinces [EMM20, PC18]. pseudo [Aze18, ND17]. pseudo-bivariate [Aze18]. pseudo-probability [ND17]. psychological [CFFVGO10]. public [AZA^{+}19, JCACTIGO11, LYZ15]. Put [Muz12]. Put-call [Muz12].

quadrant [PK17]. quadrat [AM12b]. quadratic [CBHM14]. qualitative [KNMM13]. quality [AC19, Azi20, Haj08, KA20, Mek10, SCG01, lÇl19]. quantify [Ali10]. quantile [AAA19, Ame16, DS21, ESAAA18, El 11, FH14, FS08, OAA19, WK18, Yan12]. quantiles [CT01, Vrb17a]. quantitative [MQAA21]. question [MME21]. questionnaires [ACG10]. questions [DFBL^{+}03]. queue [AS13, SPM20]. queueing [ASS10, RM14]. queues [FL09, OA09]. race [WNB13]. race-sex [WNB13]. racial [NBL15]. Radial [ORMPL17]. radiation [SC19a]. rainfall [DKS19, TS10]. Random [FKL06, AAA^{+}16, CY04, CT01, Cho08, CM13, DS08, ERH17, FM10, HB17, JL15, KR13, KLK04, MA19, MDTD10, QH18, Ryu04, SV12, UH13, WS15a, You05, Zam19]. randomization [Chr09, SR14b]. Randomized [Chr09, Arb08]. randomly [BAAA02, SR14a]. range [Sel15b]. rank [Ald13, AAE11, AIR18, Bor12, DMV13, TH14]. rank-based [TH14]. ranked [CCS16, GO15, JAO06, JAO07, Kou02]. ranked-set [Kou02]. rankings [RR03]. rare [MME21, SLKP08]. rate [AK06, AA20a, AC20, AAK10, DSG^{+}09, ES16, HZFZ06, Hsi10b, Kam19, KAA08, Kip20, Li15, Mun11, Nwa09, RORGTR20, SB06]. rates [AA17, ASS10, Kel05, LS16, YMM11]. ratings [ACM07]. Ratio [CCS16, KAK18, AK12, BAS17, BAS19, Cha03, Cos09, EJ04, IKA^{+}20, LHC01, MSIT05, NKaN11, PC13, QTH04, QKH18, SH09, TTT08, Wei15]. ratio-of-uniforms [SH09]. ratios [KB11, YQCSG10]. Rayleigh [AMAAEA19, AMA20, AM20, AAAA18, BA15, Li12, Lug16]. RBC [DRW19]. Re [AAS12]. Re-sampling [AAS12]. reaction [ABS08]. Real [AA15b, AAN21, Alm21, Bui20b, FM10, WYZ10, ZAJ20, Zak20]. real-time [Zak20]. reasons [Kam11]. recapture [TW13]. recognition [HKPGN09]. Recommendation [Han09]. record [Abo10, AS16a, Ald21, Ess17, EE19]. records [BA14]. recurrence [AS16a, EAHH13]. recurrent [STPS10]. Recursive [ASSS20, IAS19]. Recycled [BLB15]. reduce [KAAA21]. reduced [Ram16a]. reduction [BKK13, CFFVGO10, Hag14, JOO18, Kan16]. reductions [Stu02]. reforms [JKS21a, JKS21b]. regarding [SHM19]. Regency [AKGU^{+}16]. regime [Sal18]. region [AAY^{+}19, JWLZ14, KKBB18, Van20]. regions [CW08, FH14, KL15, MeMW19, Vrb11, Vrb13a, Vrb13b, Vrb15]. Regression [PC14, AE17, Abd04, AAN21, AAA19, AM12a, Abd10, AZA^{+}19, AKA20, AKAA20, AAK21, AAY^{+}19, AJP16, APL20, AAS12, AGM^{+}07, Azi18, AF19, BR12, CJ20, DHSW20, DGHP10, DT19, ESAAA18, FEM^{+}09, FS03, Fam08, FKH16, GD14, GÖ01, Hag14, Hag19, JD16, Kan01, Kha13, KKD09, KL15, KS03, KG21, LV18, LVX19, LCP05, LPLH08, MY15, MBSJ19, MA20, MESAF16, NV20, Oma18, Özt01, PG10, PRVATS14, PFL06, Par10, PACW13, RZ14, RSW14, RS14, RYS08, RG15, SP17, Sha10, Sin16a, Sin16b, Tor21, TC05, TH14, TP02, WY14, WDM14, WK18, WS15b, YS17, You05, ZX05, uaSYHNQ20]. regressions [Ame16, MSP^{+}13]. regressors [RSW14]. Regular [EKK04]. regularly [Sto06a]. regulation [Han11]. regulatory [RM14]. rehabilitation [BK12]. reinforced [RSM^{+}11]. reinforcing [SCG01]. reinsurance [Zim08]. related [AM07, CF14, MSP^{+}13, RAM15, TPP07a]. relating [Vrb17b]. Relations [JSS19, AS16a, EKK04]. Relationship [BA19, AAN21, JKS21b, KHY10, Mah10, Sha08]. relationships [MN18, VPO^{+}12, Wan13, YB13]. relative [AAH18, BV09, CBHMRA17, Ma05, MKM^{+}15, WY14, YS17]. relevant [BMS05]. Reliability [ARA^{+}18, ARA^{+}19, DC20, KAF12, TRA^{+}18, AEB13, AB17, AMAAEA19, Ald20a, Ami09, Ami14, BLV06, Bar09, Bar10, Bar12, HM10, HB17, MFPO20, UABW21]. removals [SJ10]. Renewal [HM14, WM08]. repairable [Ami14, AAED13]. repairman [Al 20]. repairmen [ARA^{+}18, ARA^{+}19, Al 20]. repeated [AS05, Arb08, ICBS14, Nis09]. repetitive [ARAA17]. replicate [Coz08]. replicates [IFVA06]. Representation [Víš15, Zoi09]. representations [ZZK13]. Republic [Moh13]. requirements [Lev13]. resampling [JM19]. Research [WZG11, PC18]. Reserve [DGPF12]. reservoir [DGW^{+}10]. reservoirs [uaSYHNQ20]. reset [Lin09]. Residual [OOA11, Dia07]. resolution [EKK04]. resource [HCJ21]. resources [GLARM18, GS10, PLRaB18]. respiratory [FFX01]. response [AE11, Ali10, ABS08, BV21, Chr09, DGM20, FCG10, KPP09, Mül02, OKM16, VVV06]. responses [PC14, Wan07]. Restricted [GG07, KT08, TW16]. restrictions [KT08]. resubmitted [AAAJ17]. result [CBHMRA12, CBHM14]. results [Abu02, BM21, Bor12, RAM15, SB11, Tza08, Yos15]. retail [Sal18]. retailers [Arr13]. retention [AA12, DRW19, SSBA20]. retrial [ASS10, FL09]. retrieval [LYK03]. return [AH21, WDM14]. returns [AF08, LVM11, Moh13, ORMPL17, SSL07, Yel20]. reveal [SdA12]. reversed [Wan10]. reversion [DNO18]. review [Bis11, LCP05, PCGOMF18]. revisited [HP04]. reweighted [Ame16]. rho [DMV13]. rice [AC19, Tin18, Tsa02]. Ridge [Wu17, AKA20, AKAA20, FKH16, Kha13, Oma18]. Ridge-Type [Wu17]. right [CT01, ORMPL17, Ram16b, SR14a]. risk [ACG10, BFG21, EH12, FEM^{+}09, FU19, Lev13, Ma05, MA20, Stu02, WYZ10, Wu01, ZP08, uaSYHNQ20]. risks [LYZ15]. RMDSE [AM07]. road [KAAA21, TW20]. Robust [AAH18, FHS20, IAS19, Özt01, TH14, AG07, DP08, PCSMGO11, SP17, TC05, Víš14, WCW14]. robustness [EI07]. ROC [ASFS11]. rocket [Lin09]. rodding [ARA^{+}18, ARA^{+}19]. role [Han11]. root [DA09, WCW14]. roots [KB11]. rotatability [BV21]. rotatable [KMT12, MKM12, MKN16, SB20, VVV06, Vic08]. rotation [TNR10]. RRT [MME21]. rspca [Zah14]. rspca-type [Zah14]. RSS [AS18]. Ruin [CY04]. rules [PD09]. run [AS16b, AS17, BAS17, SSA16, SPA18, TPP07a]. run-related [TPP07a]. Russia [AATD20].

Saddle [AAA^{+}16]. Saddlepoint [AAE12, AAE11, Mar13]. safety [GOPLD09, KAAA21, Zim08]. Salary [BSGOFGPF17]. sales [AMA13]. sample [AJM11, AAE11, ABD05, BU16, BLB10, BK02, DGHP10, Guo11, JGAF10, PK17, SMG12, SJ12, Sel15b, Tan10, TNN17b, Tsa02, TP02, UH13, Vrb17a, You05, ZW10]. samples [BLLR01, ERH17, FMA19, GO15, JAO06, JAO07, Kou02, Li11, Li15, Ras15]. Sampling [Ram16c, AK19, ARAA17, AAAJ17, AKA17, AAS12, Bak03, BLB10, BR12, BS04, CCS16, EAM10, ERH17, GS10, Ham20b, JL15, KAK18, KA20, Laa14, LC01, LC09, MN20, MH09, MEL16, PC14, QH18, QKH18, RP15, SMP21, TNR10, Vrb14a, Zam19]. São [AZRS11]. SARIMA [OO20]. SAS [QDG17, XB10]. satisfaction [JZTS17, RIT18, lÇl19]. Saudi [YAA16]. Saunders [AM10b]. scalable [YB13]. scale [AYM15, AZIE19, BMS05, LNC10, ORMPL17, Özt01, Thi08, Zgh03]. scaled [CMT16]. scales [FEW09]. scaling [AM07]. scheme [Ami14, BMS^{+}19]. schemes [BAS19]. scholarships [BSGOFGPF17]. school [RIT18]. scores [WH17]. scoring [Dry11]. screening [Sal05]. SDE [Gar08]. seasonal [Are19, CaN16, IAS19, LW04]. seasonality [DKS19, SOO08]. second [Abd10, Bak03, BV21, KMT12, MKN16, SB20, VVV06, Vic08, Zoi09]. sectional [AEH18, YQCSG10]. sections [RSM^{+}11]. sector [Alm21, ES12, JKS21a, JKS21b, LLDD17]. securities [Moh13]. seed [KM08]. seemingly [AF19, Oma18]. segmented [RG15]. seizure [DGI18]. selecting [FKH16]. Selection [EB18, BMZ^{+}17, Dry11, EW15, MSIT05, NM04, PG10, PACW13, PLRaB18, QF20, RJ20, Res04, SSRG16, Sha08, Sha10, Yeh16]. self [Han11, RM14, Sto11]. self-normalized [Sto11]. self-regulation [Han11]. Semarang [SPM20]. semi [DDJM09, JOO18, SR14c, WS15a]. semi-Markov [DDJM09]. semi-martingales [WS15a]. semi-parametric [JOO18]. semiparametric [Wu17]. sensation [SVSO02]. sensitive [MME21]. Sensitivity [AEEG19, KT08, ORMPL17]. separability [TMS08]. separable [Ven15]. sequence [Han03, KDNH18, Mun11, SZZ08]. sequences [Sto11, ZZF08]. sequential [SB09, XW14]. serial [Vrb17b, Wei15]. series [Alk18, AAK10, AAKAK11, ASM17, BM21, BLV06, BMZ^{+}17, Bui20a, Cha20, CL06, GG07, KSM16, LM11, MBSJ19, NM04, OO20, PPT02, PMS15, QDG17, RORGTR20, SM12a, Slo13, Sto08, TK04, Won20, ZLZ06]. serum [Wan14]. server [AS13]. servers [RM14]. service [ACM07, ASS10, AS13, Azi20, BK12, Maz04, PLRV15, RZC08, SPM20, lÇl19]. services [ASS10, DB20]. set [BDNN15, CCS16, GO15, JM19, JAO06, JAO07, Kou02, Yİ10, Yos15]. set-indexed [Yos15]. sets [MKN16]. settlement [AM12b]. seven [Nwa09]. several [Gun12, LB16, LPLH08, PKKHM^{+}16, TNN17a, TNN17b, TN18, TH14, Um05]. severe [EMM20]. sex [WNB13]. Shams [EAHH13]. shape [AA20a, Gun12]. Share [AA20b, CMT16, MQLA19]. Shared [WPJC21b, AEEG19]. shares [FMI20]. sharp [RMS^{+}11]. shock [Wan13]. Short [El 18, Bis11]. shqc [Yeh16]. shrinkage [AARCMC20, AS05, BuH04, LCP05]. Sibuya [LM11]. sided [SSA16]. sieve [GPP12]. sigma [Mag19]. sigma-estimators [Mag19]. signal [AS05, Cos09, LHC01]. signal-to-noise [Cos09, LHC01]. signals [Kon13, XW14]. signed [KB11]. Significance [DZS15, AARCCA16, SR14b]. similarity [HSL09]. simple [FEW09, Ram16b, Sad10, Sho08]. Simulated [MN20]. simulation [AAA19, AKA20, BMS^{+}19, GR13, GZS18, KPS08, MFR^{+}19, PCGOMF18, RYAD12, SH09, TCD20, YAA16, ZW10]. simulations [KA06]. Simultaneous [CW08, AE17, BLB10, KKMMK20, El 20]. simultaneously [Mun11]. Single [DSG^{+}09, FL09, Al 20, GZS18, Ham20b, HLZ06, PSKF17]. singular [AC20]. sites [KDNH18]. six [KAF12]. size [ABD05, BU16, BLB10, GS03, Guo11, Ras11]. size-power [ABD05]. sizes [AACARC13, AARCCA16, AASR17, SMG12, Tan10, Vic08]. skew [AJM11, AMEGM12, HNG16, Hur12, TH14, ZKZ13, ZZK13]. skew-symmetric [Hur12]. slope [BV21, SB20, Vic08]. slopes [MSP^{+}13]. small [Azi16, BK02, Con06, DGHP10, DNH20, Ela10, SLKP08, Tan10, WNB13]. small-sample [DGHP10]. Smith [DZS15]. smoke [AZ19, MHW^{+}19]. smoking [MM02]. smooth [CG07]. smooth-abrupt [CG07]. smoothing [APL19, GZS18, JWLZ14, RB03]. SNPS [EW15]. Soares [DMV13]. social [JKS21a, JKS21b]. society [Lee01]. soft [YS17]. Software [HB17, Ami14, Bar09, Bar10, Bar12, HM10]. solar [SC19a]. solution [Are19, BF16, FKE16, HTD18, Maz04]. solutions [AB01, Abu02, TKD19]. Solvency [Lev13, Ela10]. solving [JWLZ14]. SOM [Moh14]. Some [APD^{+}10, Abu02, Gar17, RAM15, Tza08, Wu02, Yos15, AAR19, AB17, BB06, Bor12, ÇA12, CF05, CBHMRA17, HA08, ND17, YMM11, ZW10]. Song [Hoa20]. sonorous [SVSO02]. soonest [CEG09]. source [Tor20]. sources [TK04]. South [AF08, AAY^{+}19, MeMW19]. southern [PC18]. southernmost [EMM20]. southwestern [SOO08]. space [BKK13, DLM11, ND17, Ram16a]. Sparse [AY13]. Spatial [GBHP^{+}12, KKBB18, PDK06, AM12a, BMZ^{+}17, BDNN15, HLBF02, JCACTIGO11, Kip11, SH09, WH17]. Spearman [DMV13, Wan13]. special [DN17]. specific [AA17, BMZ^{+}17, MKM^{+}15]. specification [KS03]. specified [LVX19, LKP04]. spectral [Azi18, DGI18, VPKG12]. spectrum [AC20]. speed [DNO18]. spending [JCACTIGO11]. spline [AG07, LT20]. splintering [XW19]. square [AS18, FH14, HP04, Hua06, IT18, KFT13, Law12, SRCT17, TMO07, YMT13]. squared [GK08]. squares [AKAA20, JZTS17, Kou02, KT08, SZZ08, ŞS10, Víš15, ZZF08]. squeezed [SdA12]. Sri [APL20, SC19a, WH20, DKS19, HCJ21, JKS21a, JKS21b]. stable [JS03, TKD19]. stage [AAA19, ESAAA18, FEW09, MME21, TO05]. stages [AS13]. stair [FR17]. starch [Cha20]. start [JOO18]. State [VPKG12, AKA17, DLM11]. states [MQLA19, SdA12, Slo13]. Stationarity [ZZF08]. Stationary [DGG05, AB01, KR01, Won20]. statistic [SRCT17, TMS08]. Statistical [AM15a, AM20, BA20, BU16, EK04, FMA19, FG10, Gas12, Gri07, Mah10, MNC17, RYS08, SMP21, AME18, AAM^{+}13, AA15b, APD^{+}10, AMA20, AKEE18, Azi16, Azi19, BB18, BRY13, CG07, Coz08, DLP12, EAHH13, EE14, FMI20, Fay19, GA09a, HKPGN09, JCACTIGO11, KDNH18, Kel05, KAAA21, LK06, MWB17, MA17, SA16, Sho08, Sim06, TKD19, VPKG12, YB13, Zak20]. statistics [AAR19, ASSS20, Al 16, Ald21, BLLR01, CBHMRA12, CBHM14, El 18, Hsi10a, HLBF02, JSS19, JL15, KY15, Kel06, MKO21, MA19, PK17, SS19, SKS08]. steel [SCG01]. steepest [KPP09]. Stein [MPZ09]. Stein-type [MPZ09]. StEM [DN10]. Step [AMAAEA19, HAE19, LS07, Ram16b, She20]. step-stress [LS07, Ram16b, She20]. Stepwise [LPKPS10]. stimulation [Han11]. Stochastic [Wan07, WM08, Wu14, ZZK13, AZRS11, AMO21, Das20, EM10, FKE16, Gar08, IS10, JPW11, Lev13, MH12, PD09, PLRaB18, SGM20, WJ13, XW19, YY20]. stochastically [UMV13]. stock [AAKAK11, AF08, Azi18, BB18, CBHMRA17, DRM15, HK20, Moh14, MRS09, Stu02, Yel20]. stockouts [Arr13]. stopping [PD09]. strategies [Wan14]. strategy [KPP09, MME21, ZP08]. stratification [GAP21, Wu01]. stratified [BLB10, ERH17, GO15, JL15, MEL16, Zam19]. stratum [JL15]. stream [Oid11]. strength [DC20]. stress [AMAAEA19, DC20, HAE19, LS07, Ram16b, She20]. stress-strength [DC20]. stroke [EAHH13]. Strong [SP17, Kho08, SZZ08]. structural [AKÜ20, GYH18, GOPLD09, GOSLVA12, MN18]. structure [BV21, IE05, KNG11, KS03, SB20]. structured [SS16]. student [JZTS17, Mah10, SSBA20, FKL06, MH09, Pol19]. Student- [FKL06]. studies [AEEG19, BU16, ESS06, IM09, PCGOMF18, YQCSG10, YB13]. Study [BB06, RSM^{+}11, AEH18, AAA19, AM12a, AA15b, AB17, AMO21, ABD05, BRY13, DLP12, DH20, EAHH13, ES16, FMI20, FDdPF10, GZS18, HBA04, HCJ21, HLBF02, JCACTIGO11, KHY10, KAK18, KAAA21, KPS08, LLDD17, MWB17, MFR^{+}19, PG10, RYAD12, SA16, SB20, Tan10, Wu02, WRP16, YQCSG10, YAA16, ZW10]. studying [AA15b]. sub [PK17]. sub-sample [PK17]. subset [Res04]. subsystem [TRA^{+}18]. successful [GOSLVA12]. successive [CW08]. Sudanese [AAN21]. sufficiency [FLMD11]. sugarcane [VPG^{+}14]. Suggested [EE14, FMI20, Azi16, Azi19]. sum [AAA^{+}16, AS17, DS08, FR17, MDTD10]. summary [Kvå11]. sums [Ryu04]. superiority [Mun11]. superpopulation [GAP21]. superprocesses [Dôk06]. supersaturated [KM02]. supervised [LTTD20, QF20]. supplementary [MKN16]. supply [Din19, Nko10]. Support [KA06, ASFS11, DHSW20, Ram16c, SOAD06]. supports [RMS^{+}11]. sure [Han03, SZZ08, YAA16]. surface [BV21, KPP09, LTvR09, OKM16, VVV06]. surveillance [Ami14]. survey [Cha12, Laa12, MEL16]. surveys [Ali10, DFBL^{+}03, Zah14]. Survival [CJ20, RA21, AA15a, AOM17, CTF11, Ela10, FEW09, FWT12, GZS18, Hsi08, LL09, LNC10, MKM^{+}15, NMW20, OMO19, Wan07, XW19]. survivors [GAAAA04]. suspended [Oid11]. sustainability [JKS21a, JKS21b, Kip17]. sustainable [HCJ21, MQAA21]. SVM [EW15]. switching [Sal18]. symmetric [ESESI15, HH02, Hur12, Kan01, SS16, WZG11, Yel20]. symmetry [ABD05, IT18, KFT13, Law12, NC03, YMT13, YKT14]. symptoms [MHW^{+}19]. syndrome [FFX01]. synthesis [TKD19]. system [ASFS11, AEB13, ASS10, BLV06, BMS05, FDGP13, GS14, JPW11, LLDD17, LHC01, Nko10, PDK06, RM14, WJ13, Yan08, Zim08]. systematic [LC01, PCGOMF18, QH18, SOAD06]. systems [Ami14, Kel05, KBS05, RMS^{+}11, SPM20]. t [CMT16]. T. [Sto08]. table [BR12, SRCT17]. tables [DT19, IT18, KNMM13, KFT13, Law12, LPKPS10, MSIT05, Mül02, SR14b, TTT08, TMO07, TMS08, YMT10, YMT13, YKT14]. Tadawul [AA20b]. tag [EW15]. Taguchi [Alt18, Cos09]. tail [Dor05, Tsa02]. tailed [FH14]. Taiwan [HLBF02]. Tamim [MA20]. target [Ram16c, ZRX10]. TASI [AA20b]. tax [AMA13, Fel14]. tea [DKS19, PBS^{+}04]. teacher [PC18]. teachers [RIT18]. Technical [EH12, ES12, SB11]. technique [AKA20, CTM18, KS18, Mos15]. techniques [AAM^{+}13, AA12, AGW19, AAS12, BMS05, Cha18, DBSH11, El 18, GF19, SMP21]. technology [DB20, SA16]. telephone [Zah14]. teller [LHC01]. temperature [BMS05, DS21, DNO18, MeMW19, WH20]. temperatures [AH21]. temporal [KKBB18]. term [CTF11, El 18, GAAAA04, WPJC21a]. terminal [BL01]. terms [AA12]. terrorist [GF19, XW19]. Test [Yİ10, AS18, AACARC13, AARCCA16, AASR17, AARCMC20, Aly16, ABD05, BB18, BuH04, CM13, CBHMRA11, DA09, FFX01, FLMD11, HAE19, Ham20b, KY15, Kel06, LS07, MPZ09, MSIT05, PC13, PCSMGO11, Ram16b, RE18, SJ12, SOAD06, SKS08, SR14a, SRCTAA13, SR14b, TMS08, Vrb14b, Wei15, YMM11]. Testing [AYM15, BMS^{+}19, GA09a, HBJ09, LKP04, MSP^{+}13, NC03, Slo13, CKC08, Cho08, CJ12, DZS15, DN17, HBA04, Mun11, PK17]. Tests [CBHMRA17, AR16, AAE11, Bak03, BR12, BAAA02, BA04, CBSBHM10, Dor05, HBA04, KNG11, KLK04, Mak14, Mül02, Sal05, She20, TC05, TP02, WCW14, YMM11, ZW10]. Texas [HLBF02]. texture [IWC03]. th [AS16a, BA14]. Thailand [Cha20, EMM20, IS21, PC18, SWCT19, TW20, TS10]. their [Has18, KPS08, Mah10, Maz04, RCAAMC19]. theorem [Dôk06, KR01, SZZ08, Zoi09]. theoretical [DFBL^{+}03, GR13]. theories [Slo13]. theory [DB20, EKK04, Ros20, XDL16, ZAJ20]. Therapeutic [PBS^{+}04]. third [AE17, MKM12]. three [Abd04, AS13, DC20, EOB07, EKK03, KMT12, PAL11, RSM^{+}11, Yel20]. three-parameter [EOB07, PAL11]. three-way [Abd04]. threshold [ABS08, GYH18]. thresholds [XDL16]. tightly [RR03]. Time [Cha20, CL06, NM04, WDGP10, AAK10, AAKAK11, ABS08, BM21, BMZ^{+}17, Bui20a, CEG09, DRW19, DNO18, GG07, GS03, HM10, HB17, Has18, KSM16, LT20, LM11, LNC10, MBSJ19, MM02, OO20, PMS15, QDG17, RORGTR20, SGM20, SM12a, SV12, SMYH19, SWCT19, Slo13, STPS10, TK04, Wan07, Won20, YSK06, YQCSG10, Zak20, ZLZ06]. time-dependent [MM02]. time-to-event [SGM20]. time-varying [LT20]. tobacco [MHW^{+}19]. tolerances [HH02]. toll [SPM20]. Tomsk [KKBB18]. Tool [SOAD06]. tooling [BMS05]. Topp [AM20, SD21]. TOPSIS [PSKF17]. total [Ald20b, RZ14]. tracking [SJH18, ZRX10]. trade [AATD20, Bui20a]. traditional [AKGU^{+}16]. traffic [TW20, Yİ10]. train [Kam11]. training [AAN21]. trajectories [DBD^{+}07]. transcription [KDNH18]. transect [EAM10]. transfer [Fel11]. transform [RE18]. transformation [KAF12]. Transformations [RW18]. transforming [Yan08]. transition [KSM16, MHW^{+}19]. Transitions [MHW^{+}19]. Transmission [FFX01]. transmutation [Afi17]. Transmuted [SC19b, ASSS20]. transparency [Tor21]. transplant [Wan14]. trapezoidal [RJ20]. treatment [AGW19, AdC14, Mül02, NBL15, SS21]. treatment-response [Mül02]. treatments [CW08]. tree [SSBA20]. tree-based [SSBA20]. trees [PD07]. Trend [BM21, DS21, YQCSG10]. tri [DP08, SB20]. tri-diagonal [DP08, SB20]. trial [Arb08]. trials [CEG09]. triplets [GS14]. tripped [Azi16]. trivariate [CF14]. truncated [Bak03, BRIA17, EE20b, Ham20b, HZ10, NLI16]. trust [JWLZ14]. Tsunami [GBHP^{+}12]. ttr [Zha20]. tuberculosis [LTTD20]. Tunisia [MJ11]. tv [PMS15]. tv-ivar [PMS15]. Two [AAA19, AZRS11, TO05, AK12, AYM15, AAY^{+}19, AAE11, AS19, BAS17, BAS19, ESAAA18, EI07, EKK03, FFX01, FEW09, GOPLD09, GA09b, JM19, JGAF10, KAA08, Kaw10, Kaw14, KY15, KNMM13, LC09, MME21, Maz04, Nin08, Oma19, RSM^{+}11, Sal05, SRCTAA13, SSA16, TMS08, UH13, WNT20, WDM14, XDL16, YMT10, ZW10]. two-level [Sal05]. two-parameter [AS19]. two-sample [JGAF10]. two-sided [SSA16]. Two-stage [AAA19, ESAAA18, FEW09, MME21]. two-way [GA09b, Nin08, TMS08, YMT10]. Type [Wu17, ASS10, CBSBHM10, EFAEH17, ES15, Hür15, IKA^{+}20, KAF12, KAK18, KM08, LS07, LCL02, MPZ09, Mun11, NA05, Oma19, Ras15, SJ12, SJ10, Sin16a, TSU01, TH02, YMT10, YXY19, Zah14, ZBF18, AA19, AASI18, AM20, BA20, BA15, FMA19, Fay19, Ram16b, She20, TTM08, YMM11]. Type-2 [TTM08]. type-I [Mun11, AA19, AASI18, Ram16b]. type-II [ES15, LS07, She20].

U.S. [ES16, GKKTS14]. UAV [TKD19]. Uhlenbeck [AMO20, DNO18]. ulcer [ZHMK09]. ultimate [Kip11]. ultrasound [MNC17]. Unbalanced [GA09b]. unbiased [SJ10, TTT08, TMO07, WS15b, Wu16]. uncertainties [Wu14]. Uncertainty [FU19, Laa12, YMT10]. unconditional [AGM^{+}07, SAGKS17]. unconditionally [Nin08]. under-dispersed [Has20]. underdeveloped [Nwa09]. underlying [BAS17, BAS19]. underreported [PFL06]. Underreporting [Par10, SS21]. understand [GF19]. understanding [AME18, LX12]. undetected [Che09]. unemployment [LP13, LS16]. unequal [KAK18, RP15, Vic08]. unified [DB20, FWT12, Zoi09]. Uniform [HZFZ06, Tsa02, RGlHnSC20, TH02]. uniformly [NL21]. uniforms [SH09]. unimodal [LWZ^{+}17]. unit [DA09, Kho08, ORMPL17, TRA^{+}18, WCW14]. United [Slo13]. units [ARA^{+}19]. univariate [ABD05, FSH10, KK09]. universe [Mag19]. universities [AM15b]. university [LVM11, RIT18, EAHH13, JZTS17]. unknown [Ald13]. unlinked [FFX01]. unrelated [AF19, MME21, Oma18]. unsupervised [GD10]. until [AATD20]. Up-crossings [KAA08]. upon [BL01, WM08]. upper [AS16a, EE19, SNaN14]. urban [SOO08, TPP05]. USA [LS16]. Use [AAN21, AM10b, DGW^{+}10, Had19, Kam11, AEH18, BS04, DBSH11, DB20, JGAF10, SWCT19]. Using [AA11, Ald20a, Alm21, Azi20, ES12, Els21, Kho08, LHC01, MA20, SSBA20, AE17, Abd04, ASFS11, AM12a, AA15a, AM15b, AMAHE12, AH21, AKGU^{+}16, Afi16, Afi17, ARAA17, Akt15, AKA20, AS18, AA12, AAK21, AA20b, AARCCA16, AM12b, AKÜ20, AKA17, Azi16, Azi18, AF19, BR12, BV21, CTM18, CW10, CJ20, DBD^{+}07, DC20, EAM10, ESAAA18, EB18, EW15, FEM^{+}09, FHS20, GS14, Gar15, GAAAA04, GLARM18, GO15, GKKTS14, GA09b, HAE19, Has18, HKPGN09, IKA^{+}20, IT18, IWC03, JCACTIGO11, JL15, KK09, KS18, KHY10, KA20, KSM16, LT20, Lee01, LYK03, LVM11, MED14, MN20, MME21, Mek10, Moa10, Moh14, Mos15, Mun11, MDTD10, MKN16, NM04, ORMPL17, OO20, PPT02, PD07, PLRaB18, PSKF17, QDG17, RZ14, Ram16a, SGM20, SJ12, SC19a, SAGKS17, SOAD06]. using [SJ10, STPS10, SMP21, SB20, TH14, TP02, VPKG12, VVV06, Vic08, WDM14, Yel20, Zam19, Zgh03, ZHMK09, Zim08, uaSYHNQ20]. UTAUT [DB20]. utilization [GLARM18].

vacations [AS13]. validates [SR14b]. validating [Gar17]. validation [AJP16, DBD^{+}07, OOA11, WH17]. validity [Tsa02]. valuation [Lin09]. value [AKA20, FAB20, FG10, IS21, KK09, LL09, PC14, SNaN14, XDL16]. valued [PSA18, PSKF17]. values [Abo10, AS16a, DGM20, ESESA16, Ess17, EE19, GA09b, TN14, Won20]. VAR [Ald13, GG07, Yan12]. variability [Dha18, MWB17]. Variable [Guo11, Ami09, AS17, AAAJ17, ASS10, BMZ^{+}17, GAP21, KKD09, Mar13, TW16, Um08]. variables [Are19, DS08, EB18, KR13, KNMM13, MA19, MESAF16, MDTD10, Ryu04, VPG^{+}14, WS15a, XDL16, YS17]. Variance [Maz04, RJ20, AJM11, Ali10, Cha03, FR17, HBA04, Hag19, MEL16, MDTD10, Oid11, QH18, Sel15a, SJ10, TN18, TNN18, Vrb17b, XDL16]. variance-based [Oid11]. variances [Aly16, FR17, Niw17, Sae17, TNN18]. variants [SA10]. variate [KNG11, NBA04, TTM08, ZKZ13, ZZK13]. variation [Abd04, AK19, Con06, EE14, MS06, NKaN11, aN16, SNaN14, WaNN14]. variational [Fuj13]. various [VPKG12]. varying [DNO18, LT20, SMYH19, Sto06a, WDGP10]. vector [DHSW20, KA06, Nis09]. vectors [CM13, Ven15]. version [Ras13]. versions [DN10]. vertical [KNMM13]. via [EKK04, SH09, TH14, ZX05, ZP08]. Vibrations [RMS^{+}11]. Vietnam [AATD20, Din19, Din20, DNH20, DH20, Hoa20, Van20]. Vietnamese [HK20]. view [Chi14]. VIKOR [PSKF17]. village [HCJ21]. violence [SHM19]. violent [EMM20]. virtual [ORMPL17]. Virus [UABW21]. visualization [Laa12]. vitality [OT07]. volatility [AZRS11, AMO21, BB18, EM10, FKL06, IS10, LT20, MH12, Thi09, YXY19, Yel20, YY20]. volkensii [OKM16]. Volterra [FKE16]. volume [BMS^{+}19]. voter [Ros20]. voxel [BMZ^{+}17]. voxel-specific [BMZ^{+}17]. waiting [CEG09]. Wald [AARCMC20]. warranty [AAED13]. water [uaSYHNQ20]. Watson [JD16]. wave [BKK13, Wu14]. Wavelet [AS05, Cha18, IWC03, SM12a]. wavelets [FKE16]. way [Abd04, GA09b, Nin08, TMS08, YMT10, YKT14]. weak [Sto11]. wealth [SSBA20, Zah14]. weather [VPO^{+}12]. Web [RR03]. Weibull [Abd08, AM10a, Afi16, AR16, AASI17, AASI18, AMA20, Ald20a, AOM17, Aze18, Bar09, BLF11, şTE^{+}11, CTF11, CT01, CJ12, DA10b, GNB01, Ham20a, Ham20b, HE16, Has18, KE19, KA20, Li11, Moa10, NMW20, OMO19, PRVATS14, Ras15, RCDO11, SS19, Sin16a, TG11, WRP16]. Weibull- [Aze18]. Weibull-generalized [Ham20a]. Weibull-generated [HE16]. Weibull-type [Sin16a]. Weight [Oma18, DMLN10, HSC09]. Weighted [BS21, ESESI15, KA19, Ald20a, AS16b, BRIA17, DMV13, Kub19, OT07, RA21, Ryu04, Víš15, YY20, YH09]. weighting [PvR04]. weights [JL15]. well [AOAS17, BAZ20]. well-being [BAZ20]. wells [SV12]. wheat [FMM19, MA17, VPKG12]. whether [Azi16, Azi19]. which [AGW19]. white [Gar08]. Whitney [Kel06]. Whittle [NK13]. Wide [RR03]. Wilcoxon [AIR18]. wild [Mak14]. wind [KS18]. Winsorized [QTH04]. Winter [Azi20]. wise [FHS20, Mun11]. within [PLRaB18]. without [Chr09, YT10]. women [MHW^{+}19, SHM19]. working [CW10]. World [Din18, RR03]. worldwide [AM12a]. wrapped [JK03]. WWLLN [KKBB18].

X [NMN18]. XBAR [Ula15]. XGBoost [RORGTR20]. XII [GAAAA04, KAF12].

yield [VPKG12, VPO^{+}12, VPG^{+}14]. yields [AAY^{+}19].

zero [AZA^{+}19, DT19, MFR^{+}19, PC13, PC14]. zero-inflated [PC13]. zerovalue [PC14]. zeros [FM10]. ZIB [BAS17, BAS19]. zone [VPO^{+}12].

References

Al-Anzi:2008:MAL

[AA08] Fawaz S. Al-Anzi. Modeling Arabic language diacritized names Markovian chains. Advances and Applications in Statistics, 8 (2):177–192, April 2008. CODEN ???? ISSN 0972-3617. URL

               http://www.pphmj.com/abstract/3083.htm.

Abdelaal:2011:UDM

[AA11] Medhat M. A. Abdelaal and Essam Fawzy Aziz. Using data mining to evaluate the foreign investment efficiency in Egypt. Advances and Applications in Statistics, 23(2):135–152, August 2011. CODEN ???? ISSN 0972-3617. URL http://www.

               pphmj.com/abstract/6085.htm.

Alber:2012:FCR

[AA12] Nader Alber and Essam Fawzy Aziz. Forecasting of customers’ retention in terms of merger and acquisition processes of Egyptian banks using panel data techniques. Advances and Applications in Statistics, 29(1):63–79, July 2012. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
7065.htm.

Abdelaal:2015:ASD

[AA15a] Medhat Mohamed Ahmed Abdelaal and Sally Hossam Eldin Ahmed. Analysis of survival data by using non-parametric methods. Advances and Applications in Statistics, 46(2):107– 117, August 2015. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/9360.htm.

AbdulRahman:2015:PSM

[AA15b] Amr Ibrahim Abdul Rahman and Mohamed Alsayid Abdel Salam Abdel Rasoul. Proposed statistical model for identifying and studying the factors affecting the occurrence of financial failure case study: Egyptian Arab Real Estate Bank. Advances and Applications in Statistics, 47(3):189–210, December 2015. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/9558.htm.

Auda:2017:CAM

[AA17] Hend Auda and Hend Attia. Comparative ARIMA models for age-specific fertility rates. Advances and Applications in Statistics, 51(5):367–388, November 2017. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/11277.
htm.

Afify:2019:PTI

[AA19] W. M. Afify and Dina H. Abdelhady. On progressive Type-I interval censored under mixed model. Advances and Applications in Statistics, 57(1):61–87, July 2019. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12714.
htm.

Alanzi:2020:BPI

[AA20a] Ayed R. A. Alanzi and S. F. Niazi Ali. Bayesian prediction intervals for a new bathtub shape failure rate lifetime model in the presence of outliers. Advances and Applications in Statistics, 63(1):1–22, July 2020. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/13231.htm.

Alhagyan:2020:FPT

[AA20b] Mohammed Alhagyan and Fuad Alduais. Forecasting the performance of Tadawul All Share Index (TASI) using geometric Brownian motion and geometric fractional Brownian motion. Advances and Applications in Statistics, 62(1):55–65, May 2020. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/13161.htm.

Alkarni:2020:PLG

[AA20c] Said H. Alkarni and Wafaa A. Alshehri. Power Lindley geometric distribution: a new model for failure analysis in business. Advances and Applications in Statistics, 61(1):1–18, March 2020. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/13093.htm.

AlMutairi:2016:SPA

[AAA^{+}16] Alya O. Al Mutairi, Ameenah Alofi, Mai Alsoubhi, Sara Alradadi, and Huda Alblwi. Saddle point approximation for the random sum Poisson–Bernoulli model. Advances and Applications in Statistics, 49(3):231–243, September 2016. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/10081.htm.

Abdel-Aziz:2019:TSQ

[AAA19] Alaa A. Abdel-Aziz and Hossameldin A. Afify. Two-stage quantile regression for dynamic panel data models with fixed effects: Monte Carlo simulation study. Advances and Applications in Statistics, 54(1):21–30, January 2019. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
12341.htm.

Alshangiti:2018:EGD

[AAAA18] Arwa M. Alshangiti, Samah A. Alghamdi, A. M. Abouammoh, and Asma H. Althagafi. Estimation of l-geometric-discrete Rayleigh model. Advances and Applications in Statistics, 52 (5):281–306, May 2018. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/11686.htm.

Aslam:2017:AVR

[AAAJ17] Muhammad Aslam, Mansour Sattam Aldosari, Muhammad Azam, and Chi-Hyuck Jun. Attribute–variable resubmitted sampling plan for controlling lot fraction nonconforming. Advances and Applications in Statistics, 50(3):243–259, March 2017. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/10630.htm.

Almendra-Arao:2013:CTS

[AACARC13] Félix Almendra-Arao, José Juan Castro-Alva, and Hortensia Reyes-Cervantes. Convergence of test sizes for the Blackwelder’s non-inferiority test. Advances and Applications in Statistics, 32(1):15–31, January 2013. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/7388.
htm.

AlyaO:2011:SAT

[AAE11] Almutairi Alya O. and Ehab F. Abd-Elfattah. Saddlepoint approximation for two sample linear rank tests. Advances and Applications in Statistics, 23(1):43–52, July 2011. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/6079.htm.

Alahmadi:2012:SAL

[AAE12] Shatha A. Alahmadi and Ehab F. Abd-Elfattah. Saddlepoint approximations for linear combination and generalization of normal-Laplace distribution. Advances and Applications in Statistics, 27(1):55–64, March 2012. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/6742.htm.

Attia:2013:HWC

[AAED13] Ahmed F. Attia, Essam A. Ahmed, and Eman A. El-Dessouky. Hybrid warranty cost models for repairable products based on log-logistic lifetime distribution. Advances and Applications in Statistics, 33(1):37–53, March 2013. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/7456.htm.

Abdelghaly:2017:ECC

[AAEE17] Abdalla Abdelghaly, Essam Ahmed, Rasha Ebaid, and Walid Elbadawy. An extension of the Cuadras copula. Advances and Applications in Statistics, 50(2):97–106, February 2017. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/10580.htm.

Auda:2018:RER

[AAH18] Hend Auda, Nagwa Albehery, and Esraa Hassan. Robust estimator for relative poverty line with application to Egypt.

Advances and Applications in Statistics, 53(4):379–400, October 2018. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/12111.htm.

Alsaleh:2010:FER

[AAK10] Mohammed Ahmed Ali Alsaleh and Ahmad Mohamad Al-Kandary. Forecasting exchange rate: a time series analysis. Advances and Applications in Statistics, 16(2):191–199, June 2010. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/5162.htm.

Alghamdi:2021:MCD

[AAK21] Nawal G. Alghamdi, Muhammad Aslam, and Khushnoor Khan. Modeling count data using Poisson regression in evaluation of educational performance: a new paradigm. Advances and Applications in Statistics, 66(1):77–95, January 2021. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13618.htm.

Alsaleh:2011:TSA

[AAKAK11] Mohammad Ahmad Alsaleh, Ahmad Mohamad Al-Kandary, and Wafaa Abdullah Al-Kharief. A time series analysis of Kuwait stock data. Advances and Applications in Statistics, 22(1):1–11, May 2011. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/5911.htm.

Abdelaal:2013:PNP

[AAM^{+}13] Medhat Mohamed Ahmed Abdelaal, Muhamed Wael Farouq Abdelazim, Hisham Abdel-Tawab Mahran Morsy, Mona Mahmoud Mohamed Ebada, Hisham Mohamed Abdelaziz Saad, and Rashad Raouf Thabet Wahba. Parametric and nonparametric statistical techniques to investigate fisheries data in Lake Nasser Egypt. Advances and Applications in Statistics, 36(1):47–73, September 2013. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/7905.htm.

Abdallah:2021:URA

[AAN21] Ahmed Saied Rahama Abdallah and Hashmia Salih El-Jack Al-Nour. Use of regression analysis to identify the relationship between training and real profits for Sudanese banks for the period (2001–2014). Advances and Applications in Statistics, 66(2):115–144, February 2021. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/13666.htm.

Abbas:2019:LBD

[AAR19] Syed Wasim Abbas, Munir Ahmad, and Sajid Rasul. Leveraging big data for official statistics: some recent developments. Advances and Applications in Statistics, 54(1):99–136, January 2019. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/12346.htm.

Almendra-Arao:2016:ANS

[AARCCA16] Félix Almendra-Arao, Hortensia Reyes-Cervantes, and José Juan Castro-Alva. Adjusting nominal significance levels and test sizes when using an asymptotic non-inferiority test. Advances and Applications in Statistics, 49(1):67–86, July 2016. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/10027.htm.

Almendra-Arao:2020:BWT

[AARCMC20] Félix Almendra-Arao, Hortensia J. Reyes-Cervantes, and Marcos Morales-Cortés. Behavior of the Wald’s test for a proportion based on a shrinkage estimator. Advances and Applications in Statistics, 63(2):207–218, August 2020. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
13295.htm.

Auda:2012:RST

[AAS12] Hend A. Auda, Essam A. Ahmed, and Mahmoud M. Sadek. Re-sampling techniques in multilevel regression models. Advances and Applications in Statistics, 31(1):55–76, November 2012. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/7187.htm.

Afsana-Al-Sharmin:2017:GWL

[AASI17] Afsana-Al-Sharmin and M. Ataharul Islam. Generalized Weibull linear models with different link functions. Advances and Applications in Statistics, 50(5):367–384, May 2017. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/10824.htm.

Afsana-Al-Sharmin:2018:GWL

[AASI18] Afsana-Al-Sharmin and M. Ataharul Islam. Generalized Weibull linear models for Type-I censoring data. Advances and Applications in Statistics, 53(2):87–102, August 2018. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/11842.htm.

Almendra-Arao:2017:ATS

[AASR17] Félix Almendra-Arao and David Sotres-Ramos. Assessing test sizes for the Farrington–Manning test. Advances and Applications in Statistics, 51(5):351–366, November 2017. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/11276.htm.

Anikin:2020:FDF

[AATD20] B. A. Anikin, O. B. Anikin, N. Q. Thuong, and P. T. Dong. Forecasting the development of foreign trade of Russia and Vietnam until 2025 based on logistic curve. Advances and Applications in Statistics, 60(1):63–78, January 2020. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12997.htm.

Ali:2019:CTR

[AAY^{+}19] Ouchar Cherif Ali, Stéphane Metchebon Takougang Aime, Rasmane Yaro, Youssouf Pare, and Blaise Some. Comparison of two regression models for predicting cotton yields in the Logone occidental region of South Chad. Advances and Applications in Statistics, 59(1):75–87, November 2019. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12921.htm.

Abbaoui:2001:ESS

[AB01] L. Abbaoui and A. Bendjeddou. Exact stationary solutions for a class of nonlinear oscillators. Advances and Applications in Statistics, 1(1):99–106, April 2001. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/37.htm.

Allal:2011:PEM

[AB11] Jelloul Allal and Mohammed Benmoumen. Parameter estimation for GARCH(1, 1) models based on Kalman filter. Advances and Applications in Statistics, 25(2):115–130, December 2011. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/6477.htm.

Al-Bossly:2017:RSS

[AB17] Afrah Al-Bossly. A reliability study on some chemotaxonomic profiles of the genus Penicillium. Advances and Applications in Statistics, 50(1):1–20, January 2017. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/10502.
htm.

Abdalla:2004:EVB

[Abd04] Ibrahim M. Abdalla. Extra-variation and Bayesian binomial regression models for estimating earnings probabilities using a three-way layout income data. Advances and Applications in Statistics, 4(3):253–264, December 2004. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
435.htm.

Annaert:2005:NLS

[ABD05] J. Annaert, G. Brys, and M. De Ceuster. A new large sample test of univariate symmetry: a comparative size-power study. Advances and Applications in Statistics, 5(3):353–370, December 2005. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/667.htm.

Abdelfattah:2008:PEW

[Abd08] Ezz H. Abdelfattah. A parametric estimation of Weibull density. Advances and Applications in Statistics, 10(1):131–147, October 2008. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/3504.htm.

Abdelrahman:2010:ALR

[Abd10] Amr I. Abdelrahman. Applying logistic regression model to the second primary cancer data. Advances and Applications in Statistics, 17(2):191–210, August 2010. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/5237.
htm.

Aboubecrine:2010:PFR

[Abo10] Mohamed M. Aboubecrine. On prediction of future k-record values. Advances and Applications in Statistics, 17(1):79–93, July 2010. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/5167.htm.

Asimit:2008:TME

[ABS08] Jennifer L. Asimit, W. John Braun, and William A. Simpson. On a threshold model for estimating response bias and its application to choice reaction time data. Advances and Applications in Statistics, 10(2):207–218, December 2008. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/3722.htm.

Abundo:2002:SCR

[Abu02] Mario Abundo. Some comparison results of solutions of diffusion and jump-diffusion equations. Advances and Applications in Statistics, 2(1):51–78, April 2002. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/72.htm;
 http://www.pphmj.com/abstract/80.htm.

Agbas:2019:CAC

[AC19] Nikka S. Agbas and Roel F. Ceballos. On the conjoint analysis of consumer’s preferences on quality attributes of rice. Advances and Applications in Statistics, 58(1):45–55, September 2019. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/12820.htm.

Alderite:2020:CAS

[AC20] Arniel A. Alderite and Anthony F. Capili. Comparison of ARIMA and singular spectrum analysis in forecasting the Philippine inflation rate. Advances and Applications in Statistics, 65(1):19–31, November 2020. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/13493.htm.

Argentero:2010:MQD

[ACG10] Piergiorgio Argentero, Paola Cerchiello, and Paolo Giudici. Models for questionnaires data in injury risk management. Advances and Applications in Statistics, 14(1):77–100, January 2010. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/4522.htm.

Arboretti:2007:NPP

[ACM07] Rosa Arboretti, Livio Corain, and Marco Marozzi. Nonparametric pooling of preference ratings for conjoint analysis experiments with application on health assistance service. Advances and Applications in Statistics, 7(1):65–80, April 2007. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/2427.htm.

Alves:2014:MBA

[AdC14] Patrick Franco Alves and Gustavo T. L. da Costa. Measuring bivariate average treatment effect. Advances and Applications in Statistics, 43(2):91–105, December 2014. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/8859.htm.

Al-Eid:2011:AMO

[AE11] Eid M. Al-Eid. An alternative method for optimizing response in general linear models. Advances and Applications in Statistics, 20(2):101–108, February 2011. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/5787.htm.

Abd:2017:IFE

[AE17] Mostafa Sayed Mostafa Abd and El Hameed El Misery. The impact of female employment and the third birth event in Egypt using simultaneous hazard regression. Advances and Applications in Statistics, 51(4):283–293, October 2017. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/11236.htm.

Ahmed:2013:GMS

[AEB13] Essam A. Ahmed and Walid M. El-Badawy. Generalized methods of system reliability estimation based on estimated component reliability. Advances and Applications in Statistics, 33 (2):115–136, April 2013. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/7560.htm.

Abd-ElWahab:2019:SAI

[AEEG19] Heba A. Abd-ElWahab, Rasha B. Elkholy, and Ahmed M. Gad. Sensitivity analysis index for shared parameter models in longitudinal studies. Advances and Applications in Statistics, 57(1):1–20, July 2019. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/12711.htm.

Abd:2018:FDC

[AEH18] Mostafa Sayed Mostafa Abd, El Hameed El Misery, and Mohamed Ali Hussein. Fertility decline and contraceptive use inequality: evidence from a cross-sectional study in Egypt. Advances and Applications in Statistics, 53(2):137–152, August 2018. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/11845.htm.

Ane:2008:LDF

[AF08] Thierry Ane and Guillaume Fabre. Local dependence functions for North and South American stock index returns. Advances and Applications in Statistics, 9(2):153–176, August 2008. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/3445.htm.

Aziz:2019:DEB

[AF19] Essam Fawzy Aziz and Nader Alber Fanous. Determinants of Egyptian bank’s profitability using seemingly unrelated regression model. Advances and Applications in Statistics, 57 (2):145–153, August 2019. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/12769.htm.

Afify:2016:ANP

[Afi16] W. M. Afify. Adding a new parameter in flexible Weibull distribution using Marshall–Olkin model and its application. Advances and Applications in Statistics, 48(2):157–167, February 2016. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/9783.htm.

Afify:2017:HDU

[Afi17] W. M. Afify. Hybridization of distributions using transmutation and proportional hazard approach. Advances and Applications in Statistics, 51(6):445–468, December 2017. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/11388.htm.

Al-Ghamedi:2007:RES

[AG07] Ateq A. Al-Ghamedi. A robust estimation of spline function. Advances and Applications in Statistics, 7(3):435–440, December 2007. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/2722.htm.

Azam:2007:EPU

[AGM^{+}07] K. Azam, A. Grami, K. Mohammad, Gh. Jandaghi, M. Karimlou, and A. Kazemnejad. Estimation of parameters in unconditional categorical regression models with incomplete data in covariates. Advances and Applications in Statistics, 7(2):281– 290, August 2007. CODEN ???? ISSN 0972-3617.

Alghamdi:2019:IMT

[AGW19] Abdulaziz Alghamdi, Alison Gray, and David Watson. Investigation of metabolomics techniques by analysis of MS Propolis data: which pre-treatment method is better? Advances and Applications in Statistics, 58(1):13–34, September 2019. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12817.htm.

Abraj:2021:JRP

[AH21] M. A. M. Abraj and A. P. Hewaarachchi. Joint return period estimation of daily maximum and minimum temperatures using copula method. Advances and Applications in Statistics, 66(2):175–190, February 2021. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/13669.htm.

Auda:2018:WRB

[AIR18] Hend A. Auda, Mohamed A. Ismail, and Ali A. Rashed. Wilcoxon rank based principal component analysis. Advances and Applications in Statistics, 52(6):431–447, June 2018. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/11723.htm.

Abbasi:2011:EPS

[AJM11] Narges Abbasi, Amrolah Jafari, and Iman Makhodom. Effect of parameters of skew normal-normal model on asymptotic distribution of sample variance. Advances and Applications in Statistics, 20(2):153–157, February 2011. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
5791.htm.

Attanayake:2016:AFC

[AJP16] A. M. C. H. Attanayake, D. D. M. Jayasundara, and T. S. G. Peiris. An application of 5-fold cross validation on a binary logistic regression model. Advances and Applications in Statistics, 49(6):443–451, December 2016. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/10381.htm.

Akman:2006:LBI

[AK06] Olcay Akman and Sevtap Kestel. Length-biased inverse Gaussian hazard rate estimation: a predictive density approach. Advances and Applications in Statistics, 6(2):217–233, August 2006. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/1850.htm.

Abramov:2008:PMO

[AK08] Vyacheslav M. Abramov and Fima C. Klebaner. Prediction of missing observations by a control method. Advances and Applications in Statistics, 8(1):109–129, February 2008. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/2842.htm.

Abdel-Karim:2012:CIE

[AK12] Amany Hassan Abdel-Karim. Confidence interval estimation for the ratio of means of two independent lognormal distributions with applications to health economics. Advances and Applications in Statistics, 26(1):41–62, January 2012. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/6619.htm.

Abdel-Karim:2016:MPN

[AK16] Amany Hassan Abdel-Karim. Modeling the proportion of neonatal mortality. Advances and Applications in Statistics, 48(2):123–139, February 2016. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/9781.htm.

Abdel-Karim:2019:EDS

[AK19] Amany Hassan Abdel-Karim. Effect of different sampling methods on estimating the confidence interval of population coefficient of variation. Advances and Applications in Statistics, 55(2):221–236, April 2019. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/12500.htm.

Aslam:2017:EMA

[AKA17] Muhammad Aslam, Nasrullah Khan, and Osama H. Arif. The EWMA moving average control chart for exponential distribution using multiple dependent state sampling. Advances and Applications in Statistics, 50(1):51–71, January 2017. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/10505.htm.

Al-Kassab:2020:CBR

[AKA20] M. M. T. Al-Kassab and M. Almousa. A comparison between ridge regression and least absolute value using simulation technique. Advances and Applications in Statistics, 60(2):137–146, February 2020. CODEN ???? ISSN 0972-3617. URL http:/

/www.pphmj.com/abstract/13042.htm.

Al-Kassab:2020:MCC

[AKAA20] Mowafaq Muhammed Al-Kassab and Mohammed Qasim Al-Awjar. A Monte Carlo comparison between least squares and the new ridge regression parameters. Advances and Applications in Statistics, 62(1):97–105, May 2020. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
13164.htm.

Almarashi:2018:EHL

[AKEE18] Abdullah M. Almarashi, Mohamed G. Khalil, M. Elgarhy, and Mamhoud M. ElSehetry. Exponentiated half logistic exponential distribution with statistical properties and applications. Advances and Applications in Statistics, 53(4):423–440, October 2018. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/12114.htm.

Adi-Kusumo:2016:CCD

[AKGU^{+}16] Fajar Adi-Kusumo, Gunardi, Herni Utami, Aluicius Irwan Endrayanto, Mulyono Nitisapto, Tony Agus Wijaya, Ardhasena Sopaheluwakan, and Titus Christiawan. Climate changes detection using Bayesian change point methods and empirical orthogonal function for improving the traditional cropping calendar in Gunung Kidul Regency of Indonesia. Advances and Applications in Statistics, 49(5):387–398, November 2016. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/10275.htm.

Akman:2005:IME

[Akm05] Olcay Akman. Iterative mixture estimation for models with length-biased component. Advances and Applications in Statistics, 5(2):171–182, August 2005. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
451.htm.

Aktas:2015:ACI

[Akt15] Serpil Aktaş. Analysis of cancer incidence data using generalized linear model with nested effect. Advances and Applications in Statistics, 44(1):43–56, January 2015. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
9016.htm.

Arslan:2020:USP

[AKÜ20] G. Arslan, B. Karabulut, and H. M. Ünver. On using structural patterns in data for classification. Advances and Applications in Statistics, 65(1):33–56, November 2020. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13494.htm.

AlMutairi:2016:DOS

[Al 16] Alya O. Al Mutairi. Distribution of the order statistics of the generalized power function. Advances and Applications in Statistics, 49(2):87–104, August 2016. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/10074.
htm.

AlRahbi:2020:CAB

[Al 20] Yaqoob Al Rahbi. A comparative analysis between a model of multiple components with a single repairman and a model of multiple components with many repairmen. Advances and Applications in Statistics, 65(1):143–150, November 2020. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13500.htm.

Albers:2012:MC

[Alb12] Willem Albers. MIXMAX charts. Advances and Applications in Statistics, 31(2):77–101, December 2012. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
7328.htm.

Alderiny:2013:BDC

[Ald13] Mahmoud M. Alderiny. Bootstrap determination of the cointegration rank with unknown lag order in VAR model: application on Egypt’s imports from the main crops. Advances and Applications in Statistics, 37(1):73–94, November 2013. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/7936.htm.

Alduais:2020:UNL

[Ald20a] Fuad Alduais. Using non-linear programming to determine weighted coefficient of balanced loss function for estimating parameters and reliability function of Weibull distribution. Advances and Applications in Statistics, 62(1):31–53, May 2020. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/13160.htm.

Alduais:2020:DOP

[Ald20b] Fuad S. Alduais. Determination of the optimal period for carrying out preventive maintenance to minimize expected total cost. Advances and Applications in Statistics, 65(2):227–246, December 2020. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/13575.htm.

Alduais:2021:BAR

[Ald21] Fuad S. Alduais. A Bayesian analysis of record statistics from the Lomax model under balanced loss functions with nonlinear programming. Advances and Applications in Statistics, 67(1): 33–45, March 2021. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/13711.htm.

Alimohammadi:2010:DAQ

[Ali10] Roshanak Alimohammadi. A different approach to quantify response variance in surveys. Advances and Applications in Statistics, 16(2):141–145, June 2010. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/5158.
htm.

Alkarni:2018:GLP

[Alk18] Said H. Alkarni. Generalized Lindley power series family of distributions. Advances and Applications in Statistics, 53(5): 537–567, November 2018. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/12163.htm.

Almansour:2021:UBD

[Alm21] Aseel Almansour. Using big data in the real estate sector. Advances and Applications in Statistics, 67(2):207–224, April 2021. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/13759.htm.

Altarazi:2018:ECI

[Alt18] Safwan Altarazi. Enhancing conformance of injection blow molding by integrating machine learning modeling and Taguchi parameter design. Advances and Applications in Statistics, 53(5):519–535, November 2018. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12162.
htm.

Aly:2016:NTO

[Aly16] Emad-Eldin A. A. Aly. Nonparametric test for ordered variances. Advances and Applications in Statistics, 49(5):399–408, November 2016. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/10276.htm.

Arenas:2007:RER

[AM07] C. Arenas and F. Mestres. RMDSE: an extension to related metric scaling. Advances and Applications in Statistics, 7(1): 97–113, April 2007. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/2429.htm.

Abdul-Moniem:2010:MME

[AM10a] Ibrahim B. Abdul-Moniem. T L-moments and L-moments estimation for the Weibull distribution. Advances and Applications in Statistics, 15(1):83–99, March 2010. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
4862.htm.

Achcar:2010:UMM

[AM10b] Jorge Alberto Achcar and Fernando Antônio Moala. Use of MCMC methods to obtain Bayesian inferences for the Birnbaum–Saunders distribution in the presence of censored data and covariates. Advances and Applications in Statistics, 17(1):1–27, July 2010. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/5163.htm.

Abdelaal:2012:SEW

[AM12a] Medhat Mohamed Ahmed Abdelaal and Hisham Abdel-Tawab Mahran Morsy. A study of the effect of worldwide governance and corruption indicators on the human development index using spatial regression analysis. Advances and Applications in Statistics, 26(2):97–112, February 2012. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/6683.htm.

Altavilla:2012:AIS

[AM12b] Anna Maria Altavilla and Angelo Mazza. On the analysis of immigrant settlement patterns using quadrat counts. The case of the city of Catania (Italy). Advances and Applications in Statistics, 29(2):111–123, August 2012. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
7129.htm.

Abdelaal:2015:SME

[AM15a] Medhat Mohamed Ahmed Abdelaal and Saeed Farouk Saber Mohamed. Statistical model of Egyptian economic growth prediction. Advances and Applications in Statistics, 47(3):225– 246, December 2015. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/9560.htm.

Abdelaal:2015:MPG

[AM15b] Medhat Mohamed Ahmed Abdelaal and Hisham Abdel-Tawab Mahran Morsy. The measurement of productivity growth in Egyptian universities using DEA-based Malmquist productivity index. Advances and Applications in Statistics, 46(3):159–181, September 2015. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/9425.htm.

Al-Marzouki:2020:SIT

[AM20] Sanaa Al-Marzouki. Statistical inference of Type II Topp Leone inverse Rayleigh distribution. Advances and Applications in Statistics, 63(2):141–150, August 2020. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
13291.htm.

Abdelaal:2013:HMM

[AMA13] Medhat M. A. Abdelaal, Moustafa Galal Moustafa, and Dalia Mohamed Samy Aglan. Hybrid model of MARIMA and ANNBP for forecasting sales tax in Egypt. Advances and Applications in Statistics, 37(1):37–71, November 2013. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/7935.htm.

Al-Marzouki:2020:KWR

[AMA20] Sanaa Al-Marzouki and Sharifah Alrajhi. Kumaraswamy Weibull Rayleigh distribution: statistical properties and application. Advances and Applications in Statistics, 64(2):251– 266, October 2020. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/13441.htm.

Al-Marshadi:2019:SSR

[AMAAEA19] A. H. Al-Marshadi, A. H. Alharbey, A. M. Abd-Elfattah, and Muhammad Aslam. Step–stress reliability for generalized Rayleigh distribution based on censoring data. Advances and Applications in Statistics, 58(2):159–178, October 2019. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/12878.htm.

Abdul-Moniem:2012:ELE

[AMAHE12] I. B. Abdul-Moniem, H. F. Abdel-Hameed, and Y. M. Abo Essa. Estimation in linear exponential distribution using different methods. Advances and Applications in Statistics, 26(1): 63–77, January 2012. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/6620.htm.

Arora:2011:APM

[AMB11] Sangeeta Arora, Kalpana K. Mahajan, and Pooja Bansal. Absolute partial mean curve: a new curve of inequality. Advances and Applications in Statistics, 24(2):139–156, October 2011. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/6281.htm.

Amerise:2016:IRC

[Ame16] Ilaria L. Amerise. Iteratively reweighted constrained quantile regressions. Advances and Applications in Statistics, 49(6): 417–441, December 2016. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/10380.htm.

Aal:2018:SAU

[AME18] Medhat M. Abdel Aal, Mamdouh A. Alim Saad Mowafy, and Nehal Etman. A statistical approach for understanding the plethora of MIRNAS and MRNAS interplay. Advances and Applications in Statistics, 53(6):693–713, December 2018. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12232.htm.

Al-Mousawi:2012:MES

[AMEGM12] Howraa Al-Mousawi, Hassan Elsalloukh, Jose H. Guardiola, and Thomas McMillan. The multivariate epsilon-skew Laplace distribution. Advances and Applications in Statistics, 27(1): 9–26, March 2012. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/6739.htm.

Aminzadeh:2009:BEV

[Ami09] M. S. Aminzadeh. Bayes estimation of K-variable reliability for multiple exponential distributions. Advances and Applications in Statistics, 12(1):127–136, June 2009. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
4004.htm.

Aminzadeh:2014:SSR

[Ami14] M. S. Aminzadeh. A surveillance scheme for reliability growth of repairable systems and software based on NHPP. Advances and Applications in Statistics, 39(2):149–160, April 2014. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/8321.htm.

Alhagyan:2020:EEG

[AMO20] Mohammed Alhagyan, Masnita Misiran, and Zurni Omar. Efficient estimators for geometric fractional Brownian motion perturbed by fractional Ornstein–Uhlenbeck process. Advances and Applications in Statistics, 62(2):203–226, June 2020. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/13215.htm.

Alhagyan:2021:ESV

[AMO21] Mohammed Alhagyan, Masnita Misiran, and Zurni Omar. On effects of stochastic volatility and long memory towards mortgage insurance models: an empirical study. Advances and Applications in Statistics, 66(2):165–174, February 2021. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13668.htm.

Niwitpong:2016:CIDb

[aN16] Sa aat Niwitpong. Confidence intervals for the difference between normal means with a known coefficient of variation. Advances and Applications in Statistics, 49(1):49–65, July 2016. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/10026.htm.

Ayodeji:2017:CCA

[AOAS17] I. O. Ayodeji, T. O. Obilade, O. A. Adelodun, and K. O. Sofowora. Canonical correlation analysis approach for the geochemistry of a Nigerian crude oil well. Advances and Applications in Statistics, 50(1):37–49, January 2017. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
10504.htm.

Atem:2017:OKI

[AOM17] Bol A. M. Atem, George O. Orwa, and Levi N. Mbugua. The odd Kumaraswamy inverse Weibull distribution with application to survival data. Advances and Applications in Statistics, 51(5):309–335, November 2017. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/11274.htm.

Abreu:2010:SSP

[APD^{+}10] G. C. G. Abreu, A. Pinheiro, R. D. Drummond, S. R. Camargo, and M. Menossi. Some statistical properties of gene expression clustering for array data. Advances and Applications in Statistics, 14(2):191–204, February 2010. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
4607.htm.

Attanayake:2019:CFA

[APL19] A. M. C. H. Attanayake, S. S. N. Perera, and U. P. Liyanage. Combining forecasts of ARIMA and exponential smoothing models. Advances and Applications in Statistics, 59(2):199– 208, December 2019. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/12957.htm.

Attanayake:2020:NBR

[APL20] A. M. C. H. Attanayake, S. S. N. Perera, and U. P. Liyanage. Negative binomial regression + autoregressive integrated moving average modelling of dengue disease in Colombo, Sri Lanka. Advances and Applications in Statistics, 64(1):103– 112, September 2020. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/13337.htm.

Ateya:2014:PUB

[AR14] Saieed F. Ateya and Mohamed M. Rizk. Prediction under bivariate generalized exponential distribution. Advances and Applications in Statistics, 38(2):129–147, February 2014. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/8172.htm.

Afify:2016:MGF

[AR16] W. M. Afify and Ahmed Ramzy. Modified goodness of fit tests for the inverse flexible Weibull distribution. Advances and Applications in Statistics, 48(4):257–272, April 2016. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/9852.htm.

AlRahbi:2018:RAR

[ARA^{+}18] Yaqoob Al Rahbi, S. M. Rizwan, B. M. Alkali, Andrew Cowell, and G. Taneja. Reliability analysis of rodding anode plant in an aluminum industry with multiple repairmen. Advances and Applications in Statistics, 53(5):569–597, November 2018. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/12164.htm.

AlRahbi:2019:RAM

[ARA^{+}19] Yaqoob Al Rahbi, S. M. Rizwan, B. M. Alkali, Andrew Cowell, and G. Taneja. Reliability analysis of multiple units with multiple repairmen of rodding anode plant in aluminum industry. Advances and Applications in Statistics, 54(1):151–178, January 2019. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/12348.htm.

Ahmad:2017:MRS

[ARAA17] Liaquat Ahmad, Iqra Rafiq, Mansour Sattam Aldosari, and Muhammad Aslam. Mixed repetitive sampling plan using EWMA. Advances and Applications in Statistics, 51(3):167– 186, September 2017. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/11133.htm.

Arboretti:2008:PAM

[Arb08] Rosa Arboretti. A permutation approach for multivariate repeated measures with application to a complex randomized clinical trial. Advances and Applications in Statistics, 9(2): 261–282, August 2008. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/3450.htm.

Areepong:2019:ESA

[Are19] Yupaporn Areepong. The exact solution of ARL on EWMA chart for non seasonal AR model with exogenous variables. Advances and Applications in Statistics, 57(1):41–59, July 2019. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/12713.htm.

Arrizabalaga:2013:PAE

[Arr13] Jose Mari Eguzkitza Arrizabalaga. A practical approach for estimating demand in periodicals with many retailers and infrequent stockouts. Advances and Applications in Statistics, 32(1):1–13, January 2013. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/7387.htm.

Amato:2005:WSA

[AS05] Umberto Amato and Theofanis Sapatinas. Wavelet shrinkage approaches to baseline signal estimation from repeated noisy measurements. Advances and Applications in Statistics, 5(1): 21–50, April 2005. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/443.htm.

Ayyappan:2013:BAQ

[AS13] G. Ayyappan and K. Sathiya. Batch arrival queue with three stages of service having server vacations and service interruptions. Advances and Applications in Statistics, 35(2):111–126, August 2013. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/7845.htm.

Al-Shomrani:2016:CRR

[AS16a] Ali A. Al-Shomrani. Characterization and recurrence relations of certain distributions of k th upper record values. Advances and Applications in Statistics, 48(5):349–361, May 2016. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/9913.htm.

Areepong:2016:EEA

[AS16b] Yupaporn Areepong and Saowanit Sukparungsee. An explicit expression of average run length of exponentially weighted moving average control chart with ARIMA(p, d, q)(P, D, Q)_{L} models. Advances and Applications in Statistics, 49(5):369– 385, November 2016. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/10274.htm.

Areepong:2017:EFA

[AS17] Yupaporn Areepong and Saowanit Sukparungsee. Explicit formulas of average run length of cumulative sum control chart for autoregressive integrated moving average with exogenous variable model. Advances and Applications in Statistics, 50(1): 21–36, January 2017. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/10503.htm.

Al-Subh:2018:CST

[AS18] Sameer Ahmad Hasan Al-Subh. Chi-square test for Laplace distribution using extreme and median RSS. Advances and Applications in Statistics, 53(5):611–623, November 2018. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12166.htm.

Amer:2019:BCM

[AS19] Yasser M. Amer and Rania M. Shalabi. Bias-corrected maximum likelihood estimation of the parameters of the twoparameter Lindley distribution. Advances and Applications in Statistics, 58(2):137–158, October 2019. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12877.
htm.

Al-Saleh:2009:MMB

[ASAF09] Mohammad Ahmad Ali Al-Saleh and Suhaila Humoud Al-Farhoud. Multivariate model for bankruptcy prediction. Advances and Applications in Statistics, 11(1):29–45, February 2009. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/3970.htm.

Abdelaal:2011:DSS

[ASFS11] Medhat Mohamed Ahmed Abdelaal, Hala Abou Sena, Muhamed Wael Farouq, and Abdel Badeeh Mohamed Salem. Decision support system for the classification of breast cancer diagnosis using ROC. Advances and Applications in Statistics, 21(2):105–124, April 2011. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/5819.htm.

Ashour:2017:LEK

[ASM17] S. K. Ashour, D. R. Said, and M. A. Mostafa. Logexponentiated Kumaraswamy power series family of distributions. Advances and Applications in Statistics, 51(2):85–129, August 2017. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/11079.htm.

Ayyappan:2010:MMR

[ASS10] G. Ayyappan, Gopal Sekar, and A. Muthu Ganapathi Subramanian. M/M/1 retrial queueing system with variable service rates under Erlang k-type services. Advances and Applications in Statistics, 18(2):189–208, October 2010. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/5406.htm.

Al-Sobhi:2020:RCM

[ASSS20] Mashail Al-Sobhi, Saman Hanif Shahbaz, and Muhammad Qaiser Shahbaz. Recursive computation for moments of order statistics for transmuted exponential distribution. Advances and Applications in Statistics, 63(2):175–189, August 2020. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/13293.htm.

Al-Sobhi:2018:GHF

[ASSSAZ18] Mashail Al-Sobhi, Saman H. Shahbaz, Muhammad Qaiser Shahbaz, and Bander Al-Zahrani. The Gauss hypergeometric family of distributions. Advances and Applications in Statistics, 52(5):307–326, May 2018. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/11687.htm.

Alkenani:2013:SMO

[AY13] Ali Alkenani and Keming Yu. Sparse MAVE with oracle penalties. Advances and Applications in Statistics, 34(2):85–105, June 2013. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/7662.htm.

Abdollahnezhad:2015:TES

[AYM15] K. Abdollahnezhad, F. Yaghmaei, and S. Mahmoudi. Testing equality of the scale parameters of Pareto distributions; two computational approaches. Advances and Applications in Statistics, 45(1):45–57, April 2015. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/9141.htm.

Abo-Zaid:2019:MME

[AZ19] Ghada Abo-Zaid. Multilevel modelling to examine the effect of mothers’ smoke prior to pregnancy on childhood and adulthood asthma. Advances and Applications in Statistics, 57(2): 105–119, August 2019. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/12766.htm.

Ahmad:2019:MZI

[AZA^{+}19] Wan Muhamad Amir W. Ahmad, Nursyabiha Zafakali, Nor Azlida Aleng, Mohd Shafiq Ibrahim, Ruhaya Hasan, and Kasypi Mokhtar. Modified zero inflated Poisson regression analysis and its application to public health data. Advances and Applications in Statistics, 55(1):1–18, March 2019. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12440.htm.

Azedine:2018:PBW

[Aze18] Grine Azedine. A pseudo-bivariate Weibull-G family. Advances and Applications in Statistics, 53(5):473–486, November 2018. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/12159.htm.

Aziz:2016:SSM

[Azi16] Essam Fawzy Aziz. A suggested statistical model for whether a small industrial business is tripped or not in Great Cairo using probit model. Advances and Applications in Statistics, 48(2):141–156, February 2016. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/9782.htm.

Aziz:2018:MEE

[Azi18] Essam Fawzy Aziz. Modelling EGX30 of Egyptian stock market using spectral analysis and harmonic regression. Advances and Applications in Statistics, 53(4):325–343, October 2018. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/12108.htm.

Aziz:2019:SSM

[Azi19] Essam Fawzy Aziz. A suggested statistical model for whether the liver is cirrhosed or not. Advances and Applications in Statistics, 56(1):43–51, May 2019. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/12565.htm.

Aziz:2020:UHM

[Azi20] Essam Fawzy Aziz. Using Holt’s method and Holt–Winter’s method to improve the performance quality of e-service FIFA2018. Advances and Applications in Statistics, 60(2):169– 177, February 2020. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/13044.htm.

Ahmed:2019:EIS

[AZIE19] Amani A. Ahmed, Alyaa R. Zahran, Moshira A. Ismail, and Abd El Naser Saad. The effect of introducing a scale parameter on the estimation of offset normal distribution. Advances and Applications in Statistics, 59(2):173–197, December 2019. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/12956.htm.

Achcar:2011:TMS

[AZRS11] Jorge A. Achcar, Henrique C. Zozolotto, Eliane R. Rodrigues, and Paulo H. Nascimento Saldiva. Two multivariate stochastic volatility models applied to air pollution data from São Paulo, Brazil. Advances and Applications in Statistics, 20(1):1–23, January 2011. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/5746.htm.

BuHamra:2004:NCP

[BA04] Sana S. BuHamra and Emad-Eldin A. A. Aly. Nonparametric change point tests for LTRC data. Advances and Applications in Statistics, 4(1):27–44, April 2004. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/419.htm.

Bdair:2014:ACP

[BA14] Omar M. Bdair and Raed R. Abu Awwad. Aging and characterization properties of k th records under mean past lifetime model. Advances and Applications in Statistics, 42(1):47–64, September 2014. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/8699.htm.

Bakoban:2015:PEG

[BA15] R. A. Bakoban and M. I. Abubaker. Parameters estimation of the generalized inverted Rayleigh distribution based on progressive Type II censoring. Advances and Applications in Statistics, 47(1):19–50, October 2015. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/9471.
htm.

Bashir:2019:RBG

[BA19] Aideed Bashir and Abeera Amir. Relationship between government expenditure on education and GDP per capita in Pakistan: an ARDL approach to cointegration. Advances and Applications in Statistics, 55(1):77–103, March 2019. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12445.htm.

Badr:2020:SIG

[BA20] Majdah M. Badr and Arwa A. Alhijji. Statistical inference of generalized logistic distribution of Type I and its properties. Advances and Applications in Statistics, 61(2):193–216, April 2020. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/13145.htm.

BuHamra:2002:NCP

[BAAA02] Sana S. BuHamra, Emad-Eldin A. A. Aly, and Shafeeqah A. Al-Awadhi. Nonparametric change point tests for randomly censored data. Advances and Applications in Statistics, 2(2): 199–216, August 2002. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/119.htm.

Baklizi:2003:ASB

[Bak03] Ayman Baklizi. Acceptance sampling based on truncated life tests in the Pareto distribution of the second kind. Advances and Applications in Statistics, 3(1):33–48, April 2003. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/157.htm.

Barghout:2009:EWS

[Bar09] May Barghout. An exponentiated Weibull software reliability model. Advances and Applications in Statistics, 13(1):111– 130, October 2009. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/4207.htm.

Barghout:2010:NPN

[Bar10] May Barghout. A non-parametric NHPP software reliability model. Advances and Applications in Statistics, 14(1):31–47, January 2010. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/4518.htm.

Barghout:2012:NNP

[Bar12] May Barghout. A new non-parametric non-homogeneous Poisson process software reliability model. Advances and Applications in Statistics, 29(2):81–99, August 2012. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
7127.htm.

Bualuang:2017:EEA

[BAS17] Direk Bualuang, Yupaporn Areepong, and Saowanit Sukparungsee. Explicit expressions of average run length of double moving average control chart for ZIB models when the underlying distribution is ratio of two Poisson means. Advances and Applications in Statistics, 51(3):187–204, September 2017. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/11134.htm.

Bualuang:2019:ACC

[BAS19] Direk Bualuang, Yupaporn Areepong, and Saowanit Sukparungsee. ARL of CUSUM control schemes for ZIB model when the underlying distribution is the ratio of two Poisson means. Advances and Applications in Statistics, 56(2):95–112, June 2019. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/12637.htm.

Bahaddad:2020:MCW

[BAZ20] Ashwaq Bahaddad, Bothinah Altaf, and Alyaa R. Zahran. Multidimensional child well-being in Jordan: a dominance approach. Advances and Applications in Statistics, 61(2):119– 147, April 2020. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/13140.htm.

BenMabrouk:2006:SSM

[BB06] Anouar Ben Mabrouk and Nidhal Ben Abdallah. Study of some multinomial cascades. Advances and Applications in Statistics, 6(3):295–303, December 2006. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/2215.
htm.

Benaid:2018:STV

[BB18] B. Benaid and H. Bouzahir. A statistical test of volatility persistence in GARCH models and application to stock exchange. Advances and Applications in Statistics, 52(6):363–374, June 2018. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/11718.htm.

Bouka:2015:NPL

[BDNN15] Stéphane Bouka, Sophie Dabo-Niang, and Guy Martial Nkiet. Non-parametric level set estimation for spatial data. Advances and Applications in Statistics, 46(2):119–158, August 2015. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/9361.htm.

Beganu:2011:OEP

[Beg11] Gabriela Beganu. Optimal estimable parametric functions in a multivariate mixed linear model. Advances and Applications in Statistics, 21(1):77–83, March 2011. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/5795.htm.

Banikhademi:2016:SIM

[BF16] Hajar Banikhademi and Hasan Salehi Fathabadi. A solution to the inverse minimum cut problem in dynamic network flows. Advances and Applications in Statistics, 49(4):267–285, October 2016. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/10225.htm.

Bonelli:2021:ECR

[BFG21] Federico Bonelli, Silvia Figini, and Alessandra Grossi. Ensemble churn risk analytics. Advances and Applications in Statistics, 66(1):61–76, January 2021. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/13617.
htm.

Bentarzi:2008:MPA

[BH08] Mohamed Bentarzi and Fay?? Hamdi. Mixture periodic autoregression with periodic ARCH errors. Advances and Applications in Statistics, 8(2):219–246, April 2008. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
3086.htm.

Bishwal:2011:FES

[Bis11] Jaya Bishwal. Financial extremes: a short review. Advances and Applications in Statistics, 25(1):1–14, November 2011. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/6427.htm.

Benton:2002:PPB

[BK02] Denise Benton and K. Krishnamoorthy. Performance of the parametric bootstrap method in small sample interval estimates. Advances and Applications in Statistics, 2(3):269–285, December 2002. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/129.htm.

Badr:2012:DME

[BK12] Atef Badr and Aristides G. Karlaftis. Duration model estimation for pavement rehabilitation and service life. Advances and Applications in Statistics, 31(1):1–19, November 2012. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/7184.htm.

Budzynski:2013:GPS

[BKK13] Robert J. Budzyński, Witold Kondracki, and Andrzej Królak. On the geometry of the parameter space reduction with application to gravitational wave data analysis. Advances and Applications in Statistics, 34(1):49–63, May 2013. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
7683.htm.

Bar-Lev:2001:ANP

[BL01] Shaul K. Bar-Lev. Applying the Neyman–Pearson lemma to distributions with terminal points depending upon a parameter. Advances and Applications in Statistics, 1(3):221–225, December 2001. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/91.htm.

Bar-Lev:2010:SDO

[BLB10] Shaul K. Bar-Lev and Benzion Boukai. A simultaneous determination of the optimal sample size and sampling allocations from a stratified finite population. Advances and Applications in Statistics, 15(1):49–59, March 2010. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
4859.htm.

Bar-Lev:2015:REP

[BLB15] Shaul K. Bar-Lev and Benzion Boukai. Recycled estimation of population pharmacokinetics models. Advances and Applications in Statistics, 47(3):247–263, December 2015. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/9561.htm.

Bereta:2011:PWD

[BLF11] Estela M. P. Bereta, Francisco Louzada, and Maria A. P. Franco. The Poisson–Weibull distribution. Advances and Applications in Statistics, 22(2):107–118, June 2011. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/6042.htm.

Bar-Lev:2001:LDS

[BLLR01] Shaul K. Bar-Lev, Naftali A. Langberg, and Benjamin Reiser. Limiting distributions for statistics based on different samples. Advances and Applications in Statistics, 1(1):27–48, April 2001. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/33.htm.

Bar-Lev:2006:LBI

[BLV06] Shaul K. Bar-Lev and Frank A. Van Der Duyn Schouten. Likelihood-based inference for series system reliability. Advances and Applications in Statistics, 6(3):277–294, December 2006. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/2214.htm.

Balasubramanyam:2021:TBC

[BM21] P. Balasubramanyam and K. Sreenivasa Murthy. Trend based correction to the predicted results for time series forecasting. Advances and Applications in Statistics, 66(1):103–114, January 2021. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/13620.htm.

Berti:2005:DTA

[BMS05] G. A. Berti, M. Monti, and L. Salmaso. DoE techniques applied to the modelling of scale in hot forging and relevant effects on temperature distribution and forging load in the tooling system. Advances and Applications in Statistics, 5(3):257– 269, December 2005. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/657.htm.

Busaman:2019:TAG

[BMS^{+}19] Anurak Busaman, Khamron Mekchay, Suchada Siripant, Somporn Chuai-Aree, and Rhysa McNeil. Testing an adaptive grid-based first order finite volume scheme for dam-break simulation. Advances and Applications in Statistics, 59(1):55–73, November 2019. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/12920.htm.

Borumandnia:2017:BDS

[BMZ^{+}17] Nasrin Borumandnia, Hamid Alavi Majd, Farid Zayeri, Ahmad Reza Baghestani, Mahmood Reza Gohari, Seyyed Mohammad Tabatabaei, and Fariborz Faeghi. Bayesian 3dimensional spatial variable selection modeling of voxelspecific HRFS for localization in FMRI time series data. Advances and Applications in Statistics, 51(6):397–426, December 2017. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/11386.htm.

Borroni:2012:EPS

[Bor12] Claudio Giovanni Borroni. Extended proof of some results concerning the finite null distribution of a rank correlation measure. Advances and Applications in Statistics, 31(2):125– 152, December 2012. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/7330.htm.

Barthel:2012:CCT

[BR12] Brian P. Barthel and Mezbahur Rahman. Comparing contingency table tests in logistic regression using bootstrap sampling. Advances and Applications in Statistics, 30(1):45–75, September 2012. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/7135.htm.

Basit:2017:CEM

[BRIA17] Abdul Basit, Anam Riaz, Zafar Iqbal, and Munir Ahmad. On comparison of entropy measures for weighted and truncated weighted exponential distributions. Advances and Applications in Statistics, 50(6):477–495, June 2017. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/10942.
htm.

Bai:2011:CCA

[BRR11] Yan Bai, Gareth O. Roberts, and Jeffrey S. Rosenthal. On the containment condition for adaptive Markov chain Monte Carlo algorithms. Advances and Applications in Statistics, 21 (1):1–54, March 2011. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/5793.htm.

Bodwin:2013:SAJ

[BRY13] Kelly Bodwin, Jeffrey S. Rosenthal, and Albert H. Yoon. A statistical approach to judicial authorship: a case study of Judge Easterbrook. Advances and Applications in Statistics, 37(2):123–148, December 2013. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/7969.htm.

Bhullar:2004:UDE

[BS04] B. S. Bhullar and Pritpal Singh. On the use of design of experiments in the cluster sampling. Advances and Applications in Statistics, 4(2):195–207, August 2004. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
431.htm.

Bimali:2021:WLB

[BS21] Milan Bimali and Khimraj Shrestha. Weighted likelihoodbased approach to hierarchical clustering. Advances and Applications in Statistics, 66(2):209–226, February 2021. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13671.htm.

Berlanga-Silvente:2017:SSF

[BSGOFGPF17] Vanesa Berlanga-Silvente, Joan Guàrdia-Olmos, Pilar Figuera-Gazo, and Ernest Pons-Fanals. Salary scholarships as a factor associated with improved academic performance. Advances and Applications in Statistics, 50(4):329–348, April 2017. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/10643.htm.

Buabeng:2019:PMP

[BTANKB19] A. Buabeng, S. Twumasi-Ankrah, K. A. Nyantakyi, and B. Kumi-Boateng. Parsimonious model for predicting cocoa production in Ghana. Advances and Applications in Statistics, 55(2):193–220, April 2019. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/12499.htm.

Bahcecitapar:2016:SPS

[BU16] Melike Bahçecitapar and Mehmet Uysal. Statistical power/ sample size analysis for linear mixed effects models with MA(1) and ARMA(1, 1) autocorrelated errors in longitudinal studies. Advances and Applications in Statistics, 49(3):211– 229, September 2016. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/10080.htm.

BuHamra:2004:APT

[BuH04] Sana S. BuHamra. Application of preliminary test and shrinkage estimation in change point analysis. Advances and Applications in Statistics, 4(2):139–153, August 2004. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
427.htm.

Bui:2020:FDT

[Bui20a] Toan Ngoc Bui. Financial development and trade openness: a time series analysis. Advances and Applications in Statistics, 62(1):67–77, May 2020. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/13162.htm.

Bui:2020:HDF

[Bui20b] Toan Ngoc Bui. How does financial depth influence real estate market? Evidence from the global financial crisis. Advances and Applications in Statistics, 61(1):65–75, March 2020. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13096.htm.

Betti:2009:FMI

[BV09] Gianni Betti and Vijay Verma. Fuzzy measures of the incidence of relative poverty and deprivation: a longitudinal and comparative perspective. Advances and Applications in Statistics, 12(2):235–273, August 2009. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/4138.htm.

Beeraka:2021:MSR

[BV21] Sulochana Beeraka and B. Re. Victorbabu. Measure of slope rotatability for second order response surface designs under intra-class correlated structure of errors using balanced incomplete block designs. Advances and Applications in Statistics, 67(2):179–205, April 2021. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/13758.htm.

Cabuk:2012:AGM

[ÇA12] H. Altan Çabuk and Fikri Akdeniz. An application of generalized maximum entropy and some biased estimation methods for the Almon distributed lag model: bootstrap efficiency. Advances and Applications in Statistics, 31(2):103–124, December 2012. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/7329.htm.

Chiangpradit:2016:CPP

[CaN16] Monchaya Chiangpradit and Sa aat Niwitpong. Coverage probability of a prediction interval for a seasonal AR(1) process. Advances and Applications in Statistics, 48(4):241–255, April 2016. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/9851.htm.

Coronel-Brizio:2014:RCQ

[CBHM14] H. F. Coronel-Brizio and A. R. Hernandez-Montoya. A result concerning quadratic forms in functions of order statistics. Advances and Applications in Statistics, 41(1):1–9, July 2014. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/8625.htm.

Coronel-Brizio:2011:CTB

[CBHMRA11] H. F. Coronel-Brizio, A. R. Hernández-Montoya, and M. E. Rodríguez-Achach. A correlation test for bivariate normality. Advances and Applications in Statistics, 24(2):157–162, October 2011. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/6282.htm.

Coronel-Brizio:2012:RAE

[CBHMRA12] H. F. Coronel-Brizio, A. R. Hernández-Montoya, and M. E. Rodríguez-Achach. A result on the asymptotic eigenvectors of covariance matrices of order statistics. Advances and Applications in Statistics, 30(2):151–155, October 2012. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/7153.htm.

Coronel-Brizio:2017:TFP

[CBHMRA17] H. F. Coronel-Brizio, A. R. Hernández-Montoya, and M. E. Rodríguez-Achach. Tests of fit for the power-law distribution and some applications to the Mexican stock market index relative changes. Advances and Applications in Statistics, 50(2): 123–136, February 2017. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/10582.htm.

Coronel-Brizio:2010:ETE

[CBSBHM10] H. F. Coronel-Brizio, A. I. Shawky, R. A. Bakoban, and A. R. Hernández-Montoya. EDF tests for the exponentiated gamma distribution under type II censoring. Advances and Applications in Statistics, 17(1):95–104, July 2010. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
5168.htm.

Craiu:2008:CPF

[CC08] Mariana Craiu and Radu V. Craiu. Choice of parametric families of copulas. Advances and Applications in Statistics, 10(1): 25–40, October 2008. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/3498.htm.

Carvalho:2011:MLM

[CCM11] F. Carvalho, R. Covas, and J. T. Mexia. Manipulating lightmodels built through covering. Advances and Applications in Statistics, 22(1):13–24, May 2011. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/5912.htm.

Chutiman:2016:REA

[CCS16] Nipaporn Chutiman, Monchaya Chiangpradit, and Sujitta Suraphee. Ratio estimator in adaptive cluster sampling based on ranked set. Advances and Applications in Statistics, 49(2): 105–116, August 2016. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/10075.htm.

Chaderjian:2009:NAD

[CEG09] Bruce Chaderjian, Morteza Ebneshahrashoob, and Tangan Gao. A numerical algorithm for distributions of soonest waiting time problems in Markov dependent trials. Advances and Applications in Statistics, 13(1):95–109, October 2009. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/4206.htm.

Consul:2005:DPD

[CF05] P. C. Consul and Felix Famoye. Dev probability distribution and some of its applications. Advances and Applications in Statistics, 5(3):271–287, December 2005. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
658.htm.

Chu:2014:TBD

[CF14] David P. Chu and Ali Reza Fotouhi. A trivariate beta distribution and related distances. Advances and Applications in Statistics, 43(1):21–36, November 2014. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/8843.
htm.

Carrera-Fernandez:2010:EAC

[CFFVGO10] María Jesús Carrera-Fernández, Guillem Feixas-Viaplana, and Joan Guàrdia-Olmos. Effect of attenuation correction for the reduction of bias in psychological measurement. Advances and Applications in Statistics, 16(1):49–64, May 2010. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/5017.htm.

Chen:2007:BAS

[CG07] Jie Chen and A. K. Gupta. A Bayesian approach to the statistical analysis of a smooth-abrupt change point model. Advances and Applications in Statistics, 7(1):115–126, April 2007. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/2430.htm.

Charif:2003:CIV

[Cha03] H. A. Charif. Confidence intervals for the variance ratio in mixed linear models. Advances and Applications in Statistics, 3(3):229–241, December 2003. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/180.htm.

Chatterjee:2012:DCP

[Cha12] Arunendu Chatterjee. Detection of change points: a survey of methodologies. Advances and Applications in Statistics, 27 (2):131–165, April 2012. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/6747.htm.

Chatterjee:2018:CDW

[Cha18] Arunendu Chatterjee. Comparison of different wavelet techniques for finding change points. Advances and Applications in Statistics, 53(4):363–378, October 2018. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12110.
htm.

Chaipitak:2020:TSA

[Cha20] Saowapa Chaipitak. Time series ARIMA model for prediction of Thailand’s monthly average cassava starch domestic price. Advances and Applications in Statistics, 63(2):191–205, August 2020. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/13294.htm.

Chen:2002:EBP

[Che02] Jann-Pygn Chen. Evaluating the bootstrap p control chart. Advances and Applications in Statistics, 2(1):19–30, April 2002. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/60.htm.

Chen:2009:NAD

[Che09] Zhenmin Chen. A note on analyzing data with undetected observations. Advances and Applications in Statistics, 12(2): 137–144, August 2009. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/4131.htm.

Chiu:2014:IIV

[Chi14] Wan-Yi Chiu. The impact of investor’s view on hedging effectiveness. Advances and Applications in Statistics, 38(2): 113–128, February 2014. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/8171.htm.

Cho:2007:MCF

[Cho07] Jang Sik Cho. Multiple comparisons for Freund’s bivariate exponential populations based on fractional Bayes factor. Advances and Applications in Statistics, 7(3):441–449, December 2007. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/2723.htm.

Cho:2008:BTM

[Cho08] Jang Sik Cho. Bayesian testing in Marshall and Olkin’s bivariate exponential model with bivariate random censored data.

Advances and Applications in Statistics, 10(2):285–293, December 2008. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/3731.htm.

Christofides:2009:RRR

[Chr09] Tasos C. Christofides. Randomized response without a randomization device. Advances and Applications in Statistics, 11(1):15–28, February 2009. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/3969.htm.

Cho:2012:BIT

[CJ12] Jangsik Cho and Yong S. Joe. Bayesian independent testing in a bivariate Weibull model. Advances and Applications in Statistics, 28(2):123–131, June 2012. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
7081.htm.

Cuarteros:2020:SAH

[CJ20] Kennet G. Cuarteros and Noel G. Cuarteros Jr. Survival analysis on hemodialysis patients using Cox regression. Advances and Applications in Statistics, 65(2):165–187, December 2020. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/13572.htm.

Cho:2008:BTI

[CKC08] Jang Sik Cho, Chang Wan Kang, and Seung Bae Choi. Bayesian testing for independence in Marshall and Olkin’s bivariate exponential model. Advances and Applications in Statistics, 8(1):101–107, February 2008. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/2841.
htm.

Choi:2001:ONN

[CKS01] Seung Hoe Choi, Hae Kyung Kim, and Byung Heum Song. Optimal nonlinear L_{p}-norm estimation. Advances and Applications in Statistics, 1(1):119–131, December 2001. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/39.htm.

Chinipardaz:2006:TSD

[CL06] Rahim Chinipardaz and Seyed Mahmoud Latifi. Time series discriminant analysis of AR(p) plus noise processes: a time domain approach. Advances and Applications in Statistics, 6 (2):129–144, August 2006. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/1847.htm.

Colin:2013:ETI

[CM13] Bernard Colin and Ernest Monga. An empirical test of independence between random vectors. Advances and Applications in Statistics, 35(2):75–110, August 2013. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
7844.htm.

Chandrasekara:2016:PEM

[CMT16] N. V. Chandrasekara, M. A. Mammadov, and C. D. Tilakaratne. Parameter estimation of multivariate scaled t distribution: an application to all share price index. Advances and Applications in Statistics, 49(4):287–303, October 2016. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/10226.htm.

Congdon:2006:MHA

[Con06] Peter Congdon. Modelling hospital activity: accounting for small area and primary care practice variation. Advances and Applications in Statistics, 6(3):335–351, December 2006. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/2218.htm.

Cossari:2009:NET

[Cos09] Anthony Cossari. A note on the efficiency of Taguchi signalto-noise ratio. Advances and Applications in Statistics, 12(2): 199–206, August 2009. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/4135.htm.

Cozzucoli:2008:GRD

[Coz08] Paolo C. Cozzucoli. On the genuine replicate design for statistical calibration. Advances and Applications in Statistics, 10 (1):111–130, October 2008. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/3503.htm.

Chiou:2001:NME

[CT01] Kuo-Ching Chiou and Lee-Ing Tong. A novel means of estimating quantiles for 2-parameter Weibull distribution under the right random censoring model. Advances and Applications in Statistics, 1(1):1–26, April 2001. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/32.htm.

Calsavara:2011:FGM

[CTF11] Vinicius F. Calsavara, Vera L. D. Tomazella, and José C. Fogo. Family generalized modified Weibull for analyzing long-term survival data. Advances and Applications in Statistics, 23(1): 59–76, July 2011. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/6081.htm.

Chandrasekara:2018:ETM

[CTM18] N. V. Chandrasekara, C. D. Tilakaratne, and M. A. Mammadov. An ensemble technique for multi class imbalanced problem using probabilistic neural networks. Advances and Applications in Statistics, 53(6):647–658, December 2018. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12230.htm.

Chen:2008:SNC

[CW08] Yuh-Ing Chen and Kuang-Lu Wang. Simultaneous nonparametric confidence regions for comparing successive treatments. Advances and Applications in Statistics, 9(1):127–139, June 2008. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/3289.htm.

Chen:2010:BUW

[CW10] Zhao-Guo Chen and Ka Ho Wu. Benchmarking using working error-models. Advances and Applications in Statistics, 19(2): 113–148, December 2010. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/5679.htm.

Chan:2004:RPU

[CY04] Gary K. C. Chan and Hailiang Yang. Ruin problems under feedback model with random interest. Advances and Applications in Statistics, 4(3):379–395, December 2004. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/441.htm.

Chimka:2012:EPG

[CZ12] Justin R. Chimka and Jia Zhou. Error probabilities for gamma-distributed individual measurements. Advances and Applications in Statistics, 27(2):97–108, April 2012. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/6745.htm.

DeJuan:2009:IBA

[DA09] Aránzazu De Juan and Antonio S. M. Arroyo. An inverted beta approximation to a MPI unit root test. Advances and Applications in Statistics, 13(2):131–164, December 2009. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/4445.htm.

deAlmeida:2010:PIC

[dA10a] Norton G. de Almeida. A possible implication of the coherent access hypothesis. Advances and Applications in Statistics, 19 (1):33–38, November 2010. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/5495.htm.

DosSantos:2010:BAB

[DA10b] Carlos Aparecido Dos Santos and Jorge Alberto Achcar. A Bayesian analysis for bivariate Weibull distributions derived from copula functions in the presence of covariates and censored data. Advances and Applications in Statistics, 15(1): 1–25, March 2010. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/4857.htm.

Dimitrakopoulou:2006:GFM

[DAL06] Theodora Dimitrakopoulou, Konstantinos Adamidis, and Sotirios Loukas. On the gamma frailty model. Advances and Applications in Statistics, 6(1):41–52, April 2006. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/16.htm.

Dasgupta:2009:OAM

[Das09] Nairanjana Dasgupta. D- and A-optimal allocations in a multipoint grouped logit designed experiment. Advances and Applications in Statistics, 12(2):145–162, August 2009. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/4132.htm.

Das:2020:GEC

[Das20] Arabinda Das. Gmm estimation of a copula-based stochastic frontier model. Advances and Applications in Statistics, 64(2): 165–180, October 2020. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/13437.htm.

Doan:2020:IUD

[DB20] Thu-Trang Thi Doan and Toan Ngoc Bui. Intention to use debit card services: an analysis in the light of the unified theory of acceptance and use of technology (UTAUT). Advances and Applications in Statistics, 65(1):133–142, November 2020. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/13499.htm.

Day:2007:CAO

[DBD^{+}07] David M. Day, Irene Bevc, Thierry Duchesne, Jeffrey S. Rosenthal, Lianne Rossman, and Frances Theodor. Comparison of adult offense prediction methods based on juvenile offense trajectories using cross-validation. Advances and Applications in Statistics, 7(1):1–46, April 2007. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/2425.htm.

DeTVieira:2011:UEC

[DBSH11] Marcel De T. Vieira, Ronaldo R. Bastos, Augusto C. Souza, and Henrique S. Hippert. On the use of exploratory and confirmatory longitudinal data analysis techniques. Advances and Applications in Statistics, 22(2):129–156, June 2011. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/6044.htm.

Demirel:2018:DDE

[DÇ18] A. Kübra Demirel and H. Eray Çelik. Dirichlet distribution and estimation of parameters. Advances and Applications in Statistics, 53(4):401–421, October 2018. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12113.
htm.

Deepthi:2020:RES

[DC20] K. S. Deepthi and V. M. Chacko. Reliability estimation of stress-strength model using three parameter generalized Lindley distribution. Advances and Applications in Statistics, 65 (1):69–88, November 2020. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/13496.htm.

Dasgupta:2013:CCA

[DCKD13] Nairanjana Dasgupta, Yibing Chen, Ananth Kalyanaraman, and Sayed Daoud. Comparison of clustering algorithms: an example with proteomic data. Advances and Applications in Statistics, 33(1):63–81, March 2013. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/7458.htm.

DAmico:2009:PAD

[DDJM09] Guglielmo D’Amico, Giuseppe Di Biase, Jacques Janssen, and Raimondo Manca. Patient’s age depending HIV/AIDS evolution analysis by means of a non-homogeneous semi-Markov model. Advances and Applications in Statistics, 11(2):199–215, April 2009. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/3979.htm.

Doron:2003:DFC

[DFBL^{+}03] Gideon Doron, Camil Fuchs, Shaul K. Bar-Lev, David Navon, and Hani Zubida. Deferred forced choice questions: theoretical aspects and empirical experimentation in political surveys. Advances and Applications in Statistics, 3(1):49–66, April 2003. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/158.htm.

Dorea:2005:SDE

[DGG05] C. C. Y. Dorea, G. L. Gilardoni, and C. R. Gonçalves. Stationary distribution estimation in hidden Markov models. Advances and Applications in Statistics, 5(2):183–195, August 2005. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/452.htm.

Deutsch:2010:MLE

[DGHP10] Roland C. Deutsch, John M. Grego, Brian Habing, and Walter W. Piegorsch. Maximum likelihood estimation with binarydata regression models: small-sample and large-sample features. Advances and Applications in Statistics, 14(2):101–116, February 2010. CODEN ???? ISSN 0972-3617. URL http:/

/www.pphmj.com/abstract/4602.htm.

Dabye:2018:DES

[DGI18] A. S. Dabye, L. Gueye, and M. A. Issaka. On detection of epileptic seizure with an approach based on power spectral density with an AR model. Advances and Applications in Statistics, 52(1):1–31, January 2018. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/11482.
htm.

Darwish:2020:FLD

[DGM20] Nesma M. Darwish, Ahmed M. Gad, and Ramadan H. Mohamed. Fitting longitudinal data with missing values in the response and covariates. Advances and Applications in Statistics, 64(2):127–142, October 2020. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/13435.htm.

Darwis:2012:RE

[DGPF12] Sutawanir Darwis, Agus Yodi Gunawan, Asep K. Permadi, and Nina Fitriyati. Reserve estimation. Advances and Applications in Statistics, 28(1):45–54, May 2012. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
6857.htm.

DiNardo:2008:MAP

[DGS08] E. Di Nardo, G. Guarino, and D. Senato. A Maple algorithm for polykays and multivariate polykays. Advances and Applications in Statistics, 8(1):19–36, February 2008. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/2836.htm.

Darwis:2010:UEK

[DGW^{+}10] Sutawanir Darwis, Agus Yodi Gunawan, Sri Wahyuningsih, Nurtiti Sunusi, Aceng Komarudin Mutaqin, and Nina Fitriyati. Use of ensemble Kalman filter for bounded reservoir. Advances and Applications in Statistics, 18(2):161–173, October 2010. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/5404.htm.

Dong:2020:ESE

[DH20] P. T. Dong and T. T. Hung. An empirical study of the effect of monetary policy on the economic growth of Vietnam. Advances and Applications in Statistics, 62(1):19–29, May 2020. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/13159.htm.

Dhakonlayodhin:2018:PFC

[Dha18] Boonkong Dhakonlayodhin. The performance of fuzzy control chart for monitoring process mean and variability. Advances and Applications in Statistics, 53(4):313–324, October 2018. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/12107.htm.

Darwis:2020:ESV

[DHSW20] Sutawanir Darwis, Nusar Hajarisman, Suliadi, and Achmad Widodo. Exploring support vector regression for bearing degradation. Advances and Applications in Statistics, 60(1): 1–10, January 2020. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/12993.htm.

Darwis:2015:KFA

[DHY^{+}15] Sutawanir Darwis, Nusar Hajarisman, Teti Sofia Yanti, Mohammad Sobri, and Irfan Muharam. Kalman filtering applied to engine fault detection. Advances and Applications in Statistics, 46(3):225–236, September 2015. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
9428.htm.

Diallo:2007:ECP

[Dia07] Sara Diallo. Estimation of the change-points of the mean residual life function. Advances and Applications in Statistics, 7(2): 157–174, August 2007. CODEN ???? ISSN 0972-3617.

Dinh:2018:WCO

[Din18] Doan Van Dinh. World crude oil prices impact on consumer price index. Advances and Applications in Statistics, 52(1): 33–54, January 2018. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/11483.htm.

Dinh:2019:AMS

[Din19] Doan Van Dinh. Analysed money supply and inflation: evidence from Vietnam economy. Advances and Applications in Statistics, 56(2):125–142, June 2019. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12639.
htm.

Dinh:2020:DCB

[Din20] Doan Van Dinh. Determinants of commercial banks’ profitability: evidence from Vietnam. Advances and Applications in Statistics, 63(1):39–58, July 2020. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/13233.
htm.

Dalpatadu:2019:ISR

[DKS19] Rohan J. Dalpatadu, Kasun S. Kodikara, and Ashok K. Singh. Impact of seasonality and rainfall on tea production in Sri Lanka. Advances and Applications in Statistics, 59(2):103– 123, December 2019. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/12953.htm.

Dorea:2011:ABE

[DLM11] Chang C. Y. Dorea, Jaques S. Lopes, and Ary V. Medino. AIC, BIC and EDC criteria for higher order Markov chain with denumerable state space. Advances and Applications in Statistics, 25(1):15–29, November 2011. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/6428.htm.

Ding:2012:SSF

[DLP12] Hua Ding, Xuewen Lu, and Lingxue Pan. A statistical study on the factors influencing CSR disclosure of listed companies in China. Advances and Applications in Statistics, 27(1):65–79, March 2012. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/6743.htm.

Diniz:2010:HBG

[DMLN10] Carlos Alberto Ribeiro Diniz, Lia Hanna Martins Morita, and Francisco Louzada-Neto. Heteroscedastic von Bertalanffy growth model and an application to a Kubbard female chicken corporeal weight growth data. Advances and Applications in Statistics, 15(1):71–81, March 2010. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/4861.htm.

Dancelli:2013:PCS

[DMV13] Livia Dancelli, Marica Manisera, and Marika Vezzoli. On Pinto da Costa and Soares’ and other weighted rank correlation measures deriving from the Spearman’s rho. Advances and Applications in Statistics, 36(2):83–104, October 2013. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/7885.htm.

Dorea:2010:CNH

[DN10] C. C. Y. Dorea and D. S. B. Martins Neto. Convergence of nonhomogeneous versions of the MCEM and StEM algorithms. Advances and Applications in Statistics, 15(2):101–113, April 2010. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/4901.htm.

Dayanian:2017:SCI

[DN17] R. Dayanian and L. Navaei. On special case of r-identification problems of multiple hypotheses testing for Markov chains. Advances and Applications in Statistics, 50(1):73–83, January 2017. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/10506.htm.

Dong:2020:ALM

[DNH20] P. T. Dong, N. T. Hong Nham, and V. D. Hoa. The application of logistic model in identifying factors affecting capital access of small and medium enterprises in Vietnam. Advances and Applications in Statistics, 61(1):19–32, March 2020. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13094.htm.

Dzupire:2018:LPB

[DNO18] Nelson Christopher Dzupire, Philip Ngare, and Leo Odongo. Lévy process based Ornstein–Uhlenbeck temperature model with time varying speed of mean reversion. Advances and Applications in Statistics, 53(3):199–224, September 2018. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12012.htm.

Doan:2020:CHC

[Doa20a] Thu-Trang Thi Doan. Cash holdings of consumer goods companies: applying GMM. Advances and Applications in Statistics, 62(2):173–184, June 2020. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/13213.htm.

Doan:2020:PCB

[Doa20b] Thu-Trang Thi Doan. Profitability of commercial banks: applying GMM. Advances and Applications in Statistics, 61(1): 77–87, March 2020. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/13097.htm.

Dogo:2014:ALM

[Dog14] Samson Henry Dogo. An adaptation of LAN method for monitoring cumulative correlations. Advances and Applications in Statistics, 43(1):1–20, November 2014. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
8842.htm.

Doku:2006:CCI

[Dôk06] Isamu Dôku. A certain class of immigration superprocesses and its limit theorem. Advances and Applications in Statistics, 6(2):145–205, August 2006. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/1848.htm.

Dorea:2005:MAD

[Dor05] Chang C. Y. Dorea. Measuring asymptotic dependence of extremes and tests based on tail indexes. Advances and Applications in Statistics, 5(2):209–227, August 2005. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
1487.htm.

Das:2008:ORF

[DP08] Rabindra Nath Das and Sung H. Park. On D-optimal robust first order designs with tri-diagonal errors. Advances and Applications in Statistics, 9(1):37–51, June 2008. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/3284.htm.

DCunha:2015:ABI

[DRM15] Juliet Gratia D’Cunha, K. Aruna Rao, and T. Mallikarjunappa. Application of Bayesian inference for the analysis of stock prices. Advances and Applications in Statistics, 46(1): 57–78, July 2015. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/9332.htm.

Daudpoto:2019:IER

[DRW19] Muhammad Rafique Daudpoto, Abdul Raziq, and Muhammad Ismail Wassan. Investigating the effect of RBC on different parameters through hydraulic retention time. Advances and Applications in Statistics, 57(2):135–144, August 2019. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/12768.htm.

Dryver:2011:FLS

[Dry11] Arthur L. Dryver. Focusing on the lower scoring data in order to improve credit scoring model selection. Advances and Applications in Statistics, 20(1):25–41, January 2011. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/5748.htm.

Dalpatadu:2008:NMC

[DS08] Rohan J. Dalpatadu and Ashok K. Singh. A numerical method for computing the probability distribution of a finite sum of independent nonnegative random variables. Advances and Applications in Statistics, 9(1):145–152, June 2008. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
3291.htm.

Das:2021:CCB

[DS21] Diponkar Das and Nazneen Sultana. Climate change in Bangladesh: a quantile trend of temperature. Advances and Applications in Statistics, 66(2):145–164, February 2021. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13667.htm.

Darwis:2009:SDA

[DSG^{+}09] S. Darwis, N. Sunusi, A. Y. Gunawan, I. W. Mangku, and S. Wahyuningsih. Single decrement approach for estimating earthquake hazard rate. Advances and Applications in Statistics, 11(2):229–237, April 2009. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/3981.htm.

Dimitri:2007:MAF

[DT07] A. Dimitri and M. Talamo. A meta-analysis framework based on hierarchical mixture model. Advances and Applications in Statistics, 7(3):341–356, December 2007. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
2717.htm.

Dureh:2019:CLR

[DT19] Nurin Dureh and Phattrawan Tongkumchum. A comparison of logistic regression and machine learning algorithms applied to zero counts data in contingency tables. Advances and Applications in Statistics, 55(1):67–76, March 2019. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
12444.htm.

Dasgupta:2015:STA

[DZS15] Nairanjana Dasgupta, Ruojin Zhang, and Tory Schmidt. Significance testing of alternate bearing in ‘Granny Smith’ apple. Advances and Applications in Statistics, 45(1):29–44, April 2015. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/9140.htm.

Elatraby:2013:PSM

[EAHH13] Amr I. A. Elatraby, Dina H. Abdel-Hady, and Abeer R. Azazy Hassan. A proposed statistical model to study the factors affecting the recurrence of an individual’s stroke by myocardial infarction occurrence with application in Ain Shams University hospitals. Advances and Applications in Statistics, 32(1): 57–81, January 2013. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/7390.htm.

Eidous:2010:FOK

[EAM10] Omar M. Eidous and Abedel-Qader Al-Masri. Fourth-order kernel method using line transect sampling. Advances and Applications in Statistics, 19(1):65–79, November 2010. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/5500.htm.

ElReheemSalem:2018:SVM

[EB18] Hanaa Abd El Reheem Salem and Mona N. Abdel Bary. Selection of variables in multi-group discriminant analysis using mathematical programming and classical approaches. Advances and Applications in Statistics, 53(6):625–645, December 2018. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/12229.htm.

Elatraby:2014:SSM

[EE14] Amr Ibrahim Abdelrahman Elatraby and Ahmed Fathy Abdelaal Elwaqdy. Suggested statistical model for describing the fluctuations in the conditional variation with application on the general index of the Egyptian capital market. Advances and Applications in Statistics, 38(1):1–36, January 2014. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/8126.htm.

Ebaid:2017:NMC

[EE17] Rasha Ebaid and Walid Elbadawy. A new modification of the Cuadras copula. Advances and Applications in Statistics, 50 (5):423–434, May 2017. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/10828.htm.

Essam:2019:EPK

[EE19] A. Amin Essam and A. El-Dessouky Eman. Estimation and prediction for Kumaraswamy exponential distribution based on upper record values. Advances and Applications in Statistics, 57(1):21–39, July 2019. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/12712.htm.

Elbatal:2020:EMO

[EE20a] I. Elbatal and M. Elgarhy. Extended Marshall–Olkin lengthbiased exponential distribution: properties and application. Advances and Applications in Statistics, 64(1):113–125, September 2020. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/13338.htm.

Elbatal:2020:NGE

[EE20b] I. Elbatal and M. Elgarhy. A new generalization of Erlangtruncated exponential distribution: properties and applications. Advances and Applications in Statistics, 64(1):63–74, September 2020. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/13334.htm.

Ebaid:2017:MPA

[EFAEH17] Rasha Ebaid, Rasha Farghali, and Samah Abo-El-Hadid. A mathematical programming approach for liu-type estimator. Advances and Applications in Statistics, 50(4):293–314, April 2017. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/10641.htm.

Elatraby:2012:TEC

[EH12] Amr I. A. Elatraby and Badrya M. M. Halaby. Technical efficiency and credit risk assessment for commercial banks under Basel II accord. Advances and Applications in Statistics, 29 (1):33–61, July 2012. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/7064.htm.

Essi:2007:RCB

[EI07] I. D. Essi and J. O. Iyaniwura. On robustness and choosing between two nonlinearities. Advances and Applications in Statistics, 7(3):451–462, December 2007. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/2724.htm.

Eapen:2004:MRE

[EJ04] Merry Eapen and K. K. Jose. Modified ratio estimators of the finite population mean based on multi-auxiliary information. Advances and Applications in Statistics, 4(1):109–115, April 2004. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/424.htm.

Evangelaras:2004:SCE

[EK04] H. Evangelaras and C. Koukouvinos. Statistical criteria for evaluating fractional factorial designs. Advances and Applications in Statistics, 4(2):181–193, August 2004. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
430.htm.

Evangelaras:2003:AGB

[EKK03] H. Evangelaras, I. Kotsireas, and C. Koukouvinos. Application of Gröbner bases to the analysis of certain two or three level factorial designs. Advances and Applications in Statistics, 3 (1):1–13, April 2003. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/148.htm.

Evangelaras:2004:RFF

[EKK04] H. Evangelaras, E. Kolaiti, and C. Koukouvinos. Regular fractional factorial designs: resolution, defining relations and blocking via coding theory. Advances and Applications in Statistics, 4(1):45–58, April 2004. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/420.htm.

ElBahi:2011:ANN

[El 11] Mostefa El Bahi. Asymptotic normality of the nonparametric conditional quantile under censorship for dependent functional data. Advances and Applications in Statistics, 23(2):107–134, August 2011. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/6084.htm.

ElBiely:2018:SLT

[El 18] Mona Mustafa El Biely. Short and long-term prediction based on complex statistics and artificial intelligence techniques. Advances and Applications in Statistics, 52(4):267–280, April 2018. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/11668.htm.

Misery:2020:FEF

[El 20] Mostafa Sayed Mostafa Abd-El Hameed El Misery. Female education and first birth event in Egypt as simultaneous processes. Advances and Applications in Statistics, 63(1):23–37, July 2020. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/13232.htm.

Elatraby:2010:FAS

[Ela10] Amr I. A. Elatraby. Factors affecting the survival and solvency of small industrial businesses in Egypt. Advances and Applications in Statistics, 17(2):105–126, August 2010. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/5232.htm.

Elsaied:2021:UOD

[Els21] Hanan Elsaied. Using observation driven models to fit count daily new confirmed cases of “COVID 19” in Egypt. Advances and Applications in Statistics, 68(1):71–91, May 2021. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13789.htm.

Elabed:2010:BAS

[EM10] Asma Graja Elabed and Afif Masmoudi. Bayesian analysis of stochastic volatility models. Advances and Applications in Statistics, 19(2):81–95, December 2010. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/5677.
htm.

Eso:2020:FAS

[EMM20] Mayuening Eso, Rhysa McNeil, and Abdunfattah Masamae. Factors associated with severe violent events in the southernmost provinces of Thailand. Advances and Applications in Statistics, 63(1):59–74, July 2020. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/13234.htm.

Elrazik:2021:MCD

[EMM21] Enayat M. Abd Elrazik, Mahmoud M. Mansour, and Salah M. Mohamed. Modeling COVID-19 daily cases in Egypt. Advances and Applications in Statistics, 68(1):111–124, May 2021. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/13791.htm.

ElAdlouni:2007:OPM

[EOB07] S. El Adlouni, T. B. M. J. Ouarda, and B. Bobée. Orthogonal projection L-moment estimators for three-parameter distributions. Advances and Applications in Statistics, 7(2):193–209, August 2007. CODEN ???? ISSN 0972-3617.

ElAty:2017:GPA

[ERH17] Nihal Abd El Aty, Mahmoud Rashwan, and Ramdan Hamed. A goal programming approach for samples overlap in stratified random sampling. Advances and Applications in Statistics, 51 (1):39–57, July 2017. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/11016.htm.

Elatraby:2012:UDE

[ES12] Amr I. A. Elatraby and Wahid A. M. Shehata. Using data envelopment analysis and discriminant analysis to evaluate the technical efficiency of private sector hospitals in Egypt. Advances and Applications in Statistics, 30(2):121–133, October 2012. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/7151.htm.

El-Sagheer:2015:EPL

[ES15] Rashad M. El-Sagheer. Estimation of the parameters of life for distributions having power hazard function based on progressively type-II censored data. Advances and Applications in Statistics, 45(1):1–27, April 2015. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/9139.htm.

Elatraby:2016:FER

[ES16] Amr I. A. Elatraby and Hisham Mohamed Abdelaziz Saad. Forecasting the exchange rate of the Egyptian pound against the U.S. dollar: an empirical study. Advances and Applications in Statistics, 48(1):49–67, January 2016. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
9619.htm.

El-Sheikh:2018:PQR

[ESAAA18] Ahmed A. El-Sheikh, Alaa A. Abdel-Aziz, and Hossameldin A. Afify. On the properties of quantile regression for dynamic panel data model using two-stage approach. Advances and Applications in Statistics, 53(5):599–609, November 2018. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12165.htm.

El-Sayed:2016:EMM

[ESESA16] Sayed Mesheal El-Sayed, Ahmed Amin El-Sheikh, and Mahmoud Mohamed Abdelwahab. Estimation of AR(1) models with missing values. Advances and Applications in Statistics, 49(6):485–492, December 2016. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/10383.htm.

El-Sayed:2014:NEP

[ESESI14] Sayed Mesheal El-Sayed, Ahmed Amin El-Sheikh, and Mohamed Khalifa Ahmed Issa. New estimators of AR(2) panel data model. Advances and Applications in Statistics, 41(2): 129–136, August 2014. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/8694.htm.

El-Sayed:2015:WSE

[ESESI15] Sayed Mesheal El-Sayed, Ahmed Amin El-Sheikh, and Mohamed Khalifa Ahmed Issa. Weighted symmetric estimators of autoregressive models. Advances and Applications in Statistics, 47(2):145–152, November 2015. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/9495.htm.

El-Sayed:2017:CFB

[ESESIA17] Sayed Mesheal El-Sayed, Ahmed Amin El-Sheikh, Mohamed Khalifa Ahmed Issa, and Hadia Faried Mohamed Ahmed Azmy. A closed form of biased AR(1) model. Advances and Applications in Statistics, 50(3):191–199, March 2017. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/10627.htm.

Erickson:2006:BMB

[ESS06] Janelle S. Erickson, James D. Stamey, and John W. Seaman. Bayesian methods for bioequivalence studies. Advances and Applications in Statistics, 6(1):71–85, April 2006. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/19.htm.

Essam:2017:RVP

[Ess17] A. Amin Essam. On record values of the power Lomax distribution. Advances and Applications in Statistics, 50(6):497– 515, June 2017. CODEN ???? ISSN 0972-3617. URL http:/

/www.pphmj.com/abstract/10943.htm.

Elatraby:2015:FSU

[EW15] Amr I. A. Elatraby and Rashad R. T. Wahba. Features selection using parametric and non-parametric methods: tag SNPS selection using GA-SVM and GA-KNN. Advances and Applications in Statistics, 45(2):105–123, May 2015. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
9175.htm.

Ferreira:2020:VMH

[FAB20] Carina Ferreira, Teresa Abreu, and Mário Basto. p-value misconceptions in health professionals and patients. Advances and Applications in Statistics, 63(1):75–83, July 2020. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13235.htm.

Famoye:2008:BGP

[Fam08] Felix Famoye. Bivariate generalized Poisson regression model. Advances and Applications in Statistics, 9(1):53–69, June 2008. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/3285.htm.

Fayomi:2019:THL

[Fay19] Aisha Fayomi. Type I half logistic power Lomax distribution: statistical properties and application. Advances and Applications in Statistics, 54(1):85–98, January 2019. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12345.htm.

Ferrao:2010:DFI

[FCG10] M. E. Ferrão, P. Costa, and J. M. R. Gama. Distribution-free item response model based on marginal maximum likelihood estimation. Advances and Applications in Statistics, 18(2): 109–126, October 2010. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/5402.htm.

Furlan:2010:ELL

[FDdPF10] Camila Pedrozo Rodrigues Furlan, Carlos Alberto Ribeiro Diniz, and Maria Aparecida de Paiva Franco. Estimation of lag length in distributed lag models: a comparative study. Advances and Applications in Statistics, 17(2):127–142, August 2010. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/5233.htm.

Fitriyati:2013:EEC

[FDGP13] Nina Fitriyati, Sutawanir Darwis, Agus Yodi Gunawan, and Asep Kurnia Permadi. Exploring EnKF convergence for linear dynamical system. Advances and Applications in Statistics, 32 (2):151–162, February 2013. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/7444.htm.

Fellman:2011:DTP

[Fel11] Johan Fellman. Discontinuous transfer policies with a given Lorenz curve. Advances and Applications in Statistics, 20(2): 133–141, February 2011. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/5789.htm.

Fellman:2014:PCT

[Fel14] Johan Fellman. The properties of a class of tax policies. Advances and Applications in Statistics, 39(2):125–148, April 2014. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/8320.htm.

Fallah:2009:DMA

[FEM^{+}09] Nader Fallah, Mohammad Reza Eshraghian, Kazem Mohammad, Keramat Nourijelyani, Reza Hamidian, and Mahdi Najafi. Does magnesium affect the risk of atrial fibrillation? An approach using generalized additive logistic regression. Advances and Applications in Statistics, 12(1):11–23, June 2009. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/3998.htm.

Feng:2009:SNA

[FEW09] Changyong Feng, Carlo Ercoli, and Hongyue Wang. A simple nonparametric approach of survival data with two-stage outcomes of different scales. Advances and Applications in Statistics, 12(1):1–9, June 2009. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/3997.htm.

Fan:2001:TDT

[FFX01] Ruzong Fan, Joanna Floros, and Momiao Xiong. Transmission disequilibrium test of two unlinked disease loci; application to respiratory distress syndrome. Advances and Applications in Statistics, 1(3):277–308, December 2001. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/95.htm.

Figini:2010:SMC

[FG10] Silvia Figini and Paolo Giudici. Statistical models for customer lifetime value. Advances and Applications in Statistics, 18(1): 41–55, September 2010. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/5369.htm.

Farahmand:2011:EMA

[FG11] K. Farahmand and Jianliang Gao. Exceedance measure of amplified and de-amplified classes of algebraic polynomials. Advances and Applications in Statistics, 20(2):89–99, February 2011. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/5786.htm.

Fajardo:2014:LSM

[FH14] Val Andrei Fajardo and Mei Ling Huang. A least square method on confidence regions for high quantile of heavy tailed distributions. Advances and Applications in Statistics, 38(2): 81–111, February 2014. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/8170.htm.

Faridzah:2020:RAM

[FHS20] J. Faridzah, H. A. Hazlina, and S. Y. Sharipah Soaad. Robust alternatives to MEWMA control chart using distancebased and coordinate-wise robust estimators. Advances and Applications in Statistics, 60(1):11–33, January 2020. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12994.htm.

Farahani:2016:NSS

[FKE16] B. Farahani, M. Khodabin, and R. Ezzati. Numerical solution of stochastic nonlinear Volterra integral equations by a stochastic operational matrix based on Haar wavelets. Advances and Applications in Statistics, 48(5):317–336, May 2016. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/9911.htm.

Farag:2016:MPM

[FKH16] Aya Farag, Ragaa Kassem, and Ramadan Hamed. Mathematical programming model for selecting generalized ridge regression parameter. Advances and Applications in Statistics, 48 (6):391–409, June 2016. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/9997.htm.

Fielding:2006:RVO

[FKL06] Mark Fielding, Fima C. Klebaner, and Zinoviy Landsman. Random volatility and option prices with the generalized Student-t distribution. Advances and Applications in Statistics, 6(1):111–120, April 2006. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/21.htm.

Farahmand:2009:SRQ

[FL09] K. Farahmand and T. Li. Single retrial queues with loss of customers. Advances and Applications in Statistics, 11(2):185– 197, April 2009. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/3978.htm.

Firoozi:2011:LMT

[FLMD11] Fathali Firoozi, Donald Lien, Saeid Mahdavi, and Lynda De La Vina. A linear model test of sufficiency for nonorthogonality of matrices. Advances and Applications in Statistics, 22(2):119–127, June 2011. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/6043.
htm.

Farahmand:2010:CNR

[FM10] K. Farahmand and B. Mcguinness. Covariance of the number of real zeros of a random algebraic polynomial. Advances and Applications in Statistics, 16(1):17–24, May 2010. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/5014.htm.

Fayomi:2019:SIE

[FMA19] Aisha Fayomi, Dalal Mal, and Samia Adham. Statistical inference of the exponentiated Kumaraswamy–exponential distribution based on complete and Type II censored samples. Advances and Applications in Statistics, 54(2):273–288, February 2019. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/12401.htm.

Fawzy:2020:SSM

[FMI20] Essam Fawzy, Khaled Mohamed, and Fatma Ibrahim. Suggested statistical models to study the affecting factors on the performance of shares and forecast the daily closing price (applied study on the shares of Google company). Advances and Applications in Statistics, 65(2):247–262, December 2020. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13576.htm.

Fawzy:2019:MFW

[FMM19] Essam Fawzy, Badrya Mahmoud, and Aya Mongy. Modeling and forecasting of wheat production in Egypt. Advances and Applications in Statistics, 59(1):89–101, November 2019.

CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/12922.htm.

Feng:2010:NCC

[FMTW10] Changyong Feng, Yan Ma, Xin M. Tu, and Hongyue Wang. A note on the concordance correlation coefficient. Advances and Applications in Statistics, 15(2):195–205, April 2010. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/4907.htm.

Fernandes:2017:MMS

[FR17] C. Fernandes and P. Ramos. A method to minimize the sum of the variances of the estimators of the variance components in stair nested designs. Advances and Applications in Statistics, 51(4):277–282, October 2017. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/11235.htm.

Famoye:2003:IGP

[FS03] Felix Famoye and Karan P. Singh. On inflated generalized Poisson regression models. Advances and Applications in Statistics, 3(2):145–158, August 2003. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
167.htm.

Frydman:2008:DQE

[FS08] Halina Frydman and Gary Simon. Discrete quantile estimation. Advances and Applications in Statistics, 9(2):177–203, August 2008. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/3446.htm.

Furukatsu:2010:CUN

[FSH10] Taichi Furukatsu, Daisuke Shimamoto, and Hiroto Hyakutake. On comparison of univariate normal mean and elements of multivariate normal mean. Advances and Applications in Statistics, 16(2):147–161, June 2010. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
5159.htm.

Figini:2019:UIA

[FU19] Silvia Figini and Pierpaolo Uberti. Uncertainty interval to assess performances of credit risk models. Advances and Applications in Statistics, 57(2):121–133, August 2019. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
12767.htm.

Fujita:2013:ENP

[Fuj13] Masato Fujita. Estimation of navigation performance and offset by the EM algorithm and the variational Bayesian methods. Advances and Applications in Statistics, 35(1):1–27, July 2013. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/7782.htm.

Feng:2012:UDC

[FWT12] Changyong Feng, Hongyue Wang, and Xin M. Tu. A unified definition of conditional expectation and its applications in survival analysis. Advances and Applications in Statistics, 27 (2):81–95, April 2012. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/6744.htm.

Gil-Alana:2009:TSM

[GA09a] Luis A. Gil-Alana. Testing of I(d) statistical models with nonparametric disturbances. Advances and Applications in Statistics, 13(2):205–223, December 2009. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/4449.htm.

Gunasekera:2009:UTW

[GA09b] Sumith Gunasekera and Malwane M. A. Ananda. Unbalanced two-way ANOVA additive model under heteroscedasticity using generalized p-values. Advances and Applications in Statistics, 11(1):1–13, February 2009. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/3968.htm.

Ghitany:2004:MPL

[GAAAA04] M. E. Ghitany, F. A. Al-Awadhi, and S. A. Al-Awadhi. Modeling the presence of long-term survivors using generalized Burr XII mixture model. Advances and Applications in Statistics, 4(1):15–26, April 2004. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/319.htm.

Ghitany:2003:GGD

[GAM03] M. E. Ghitany and D. K. Al-Mutairi. A generalized geometric distribution and its application. Advances and Applications in Statistics, 3(1):67–76, April 2003. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/160.htm.

Gupt:2021:OSA

[GAP21] Bhuwaneshwar Kumar Gupt, Md. Irphan Ahamed, and Manoshi Phukon. Optimum stratification for an auxiliary variable optimum allocation under a superpopulation model. Advances and Applications in Statistics, 67(1):1–20, March 2021. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/13709.htm.

Garcia:2008:CSI

[Gar08] Jorge Garcia. Convergence of stochastic integrals and SDE’s associated to approximations of the Gaussian white noise. Advances and Applications in Statistics, 10(2):155–177, December 2008. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/3718.htm.

Gargoum:2015:OPC

[Gar15] Ali S. Gargoum. Optimizing probability computations on dynamic processes using graphical models. Advances and Applications in Statistics, 44(2):145–161, February 2015. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/9022.htm.

Gargoum:2017:SNP

[Gar17] Ali. S. Gargoum. Some new properties of Hellinger distance for validating approximations in Bayesian analysis. Advances and Applications in Statistics, 51(4):261–276, October 2017. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/11234.htm.

Gasmi:2012:SII

[Gas12] Soufiane Gasmi. Statistical inference of an imperfect maintenance model. Advances and Applications in Statistics, 30(2): 93–120, October 2012. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/7150.htm.

Goulionis:2007:MPU

[GB07] John E. Goulionis and V. K. Benos. A maintenance policy under Markovian deterioration. Advances and Applications in Statistics, 7(3):357–388, December 2007. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
2718.htm.

Grancher:2012:SIC

[GBHP^{+}12] D. Grancher, A. Bar-Hen, R. Paris, F. Lavigne, and D. Brunstein. Spatial interpolation of circular data: application to Tsunami of December 2004. Advances and Applications in Statistics, 30(1):19–29, September 2012. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/7133.
htm.

Gamage:2020:MJE

[GCN20] Ramadha D. Piyadi Gamage, Ying-Ju Chen, and Wei Ning. Modified jackknife empirical likelihood. Advances and Applications in Statistics, 60(2):201–215, February 2020. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13047.htm.

Ghattas:2010:MUD

[GD10] Badih Ghattas and Claude Deniau. Multivariate unsupervised discretization preserving mutual information. Advances and Applications in Statistics, 19(1):49–64, November 2010. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/5499.htm.

Geneti:2014:CUF

[GD14] Kasahun Takele Geneti and Teshome Kebede Deressa. Correlates of under five child mortality in Ethiopia: a logistic regression analysis. Advances and Applications in Statistics, 39 (1):61–79, March 2014. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/8239.htm.

Galvani:2019:MLT

[GF19] Marta Galvani and Silvia Figini. Machine learning techniques to understand the dynamics of terrorist events. Advances and Applications in Statistics, 56(1):53–76, May 2019. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12566.htm.

Gomez:2007:RVF

[GG07] Nicolas Gomez and Victor M. Guerrero. Restricted VAR forecasts of economic time series with contemporaneous constraints. Advances and Applications in Statistics, 7(3):313– 340, December 2007. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/2716.htm.

Gupta:2008:DMS

[GK08] A. K. Gupta and D. G. Kabe. On the distribution of the magnitude-squared coherence estimator. Advances and Applications in Statistics, 10(2):195–200, December 2008. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/3720.htm.

Guemmegne:2014:MDU

[GKKTS14] J. T. Guemmegne, J-J Kengwoung-Keumo, M. A. Tabatabai, and K. P. Singh. Modeling the dynamics of the U.S. healthcare expenditure using a hyperbolastic function. Advances and Applications in Statistics, 42(2):95–117, October 2014. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/8724.htm.

Gildeh:2018:OUR

[GLARM18] Fatemeh Keshavarz Gildeh, Farhad Hosseinzadeh Lotfi, Alireza Amirteimoori, and Mohsen Rostamy-Malkhalifeh. Optimal utilization of resources in organizations using data envelopment analysis. Advances and Applications in Statistics, 52(2):73–95, February 2018. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/11504.htm.

Gandomi:2001:LBE

[GNB01] A. Gandomi, S. Noorbaloochi, and A. Bozorgnia. Linex Bayes estimation of Weibull distribution. Advances and Applications in Statistics, 1(1):107–118, April 2001. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/38.htm.

Gurcan:2001:EPL

[GÖ01] M. Gürcan and Y. Öner. On the existence problem in logistic regression models by alternative form. Advances and Applications in Statistics, 1(2):165–174, August 2001. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/43.htm.

Gory:2015:ANI

[GO15] Jeffrey Gory and Omer Ozturk. Analysis of the NHANES III data set using ranked set and judgment post-stratified samples. Advances and Applications in Statistics, 47(1):65–89, October 2015. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/9473.htm.

Guardia-Olmos:2008:BIB

[GOdlFSLF08] Joan Guàrdia-Olmos, Emilia I. de la Fuente-Solana, and Luis Manuel Lozano-Fernández. Bayesian inference for binomial populations. Bayesian estimation for child depression prevalence. Advances and Applications in Statistics, 9(1):13– 35, June 2008. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/3283.htm.

Galan:2007:ACD

[GOG07] Ferran Galán, Francesc Oliva, and Joan Guàrdia. Assessment and comparison of discriminant analysis methods and nonparametric error estimators with mixed data. Advances and Applications in Statistics, 7(2):255–280, August 2007. CO-DEN ???? ISSN 0972-3617.

Guardia-Olmos:2009:BSE

[GOPLD09] Joan Guàrdia-Olmos, Sergi Valera Pertegas, Daniel Carro Lemos, and Emilia Inmaculada De La Fuente Solana. Bayesian structural equation models: analysis of two alternative models about safety perception. Advances and Applications in Statistics, 11(2):157–172, April 2009. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/3976.htm.

Guardia-Olmos:2012:SMA

[GOSLVA12] Joan Guàrdia-Olmos, Esther Subirà-Lobera, and José María Vilas-Aventín. A structural model to assess the factors that influence successful brand migration. Advances and Applications in Statistics, 31(1):21–42, November 2012. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
7185.htm.

Gonzalez:2010:CCG

[GP10] Eduardo Gutiérrez González and Olga Vladimirovna Panteleeva. Control charts for generalized log-gamma processes. Advances and Applications in Statistics, 18(1):89–107, September 2010. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/5372.htm.

Gerolimetto:2012:PDH

[GPP12] M. Gerolimetto, C. Pizzi, and I. Procidano. A procedure to detect hidden cointegration with the sieve bootstrap. Advances and Applications in Statistics, 30(2):77–92, October 2012. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/7149.htm.

Gaines:2013:TNS

[GR13] G. Gaines and F. Ruymgaart. A theoretical note on the simulation of Brownian motion. Advances and Applications in Statistics, 33(2):99–106, April 2013. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/7558.htm.

Guria:2015:DLL

[GR15] Sibnarayan Guria and Sugata Sen Roy. Diagnostics in loglinear models. Advances and Applications in Statistics, 44(2): 163–176, February 2015. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/9023.htm.

Grillenzoni:2007:SOD

[Gri07] Carlo Grillenzoni. Statistical optimization of digital image enhancement. Advances and Applications in Statistics, 7(2): 229–244, August 2007. CODEN ???? ISSN 0972-3617.

Grillenzoni:2011:AMF

[Gri11] Carlo Grillenzoni. Adaptive methods for financial decisions. Advances and Applications in Statistics, 21(2):125–139, April 2011. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/5820.htm.

Gor:2003:OLL

[GS03] Ravi Gor and Nita H. Shah. An order level lot-size model with time dependent deterioration and permissible delay in payments. Advances and Applications in Statistics, 3(2):159– 172, August 2003. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/170.htm.

Goulionis:2008:POM

[GS08] John E. Goulionis and Dimitrios I. Stengos. The partially observable Markov decision processes framework in medical decision making. Advances and Applications in Statistics, 9 (2):205–232, August 2008. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/3447.htm.

Guo:2010:EAS

[GS10] Bao-Cai Guo and Li-Rong Sun. The effect of the allocation of sampling resources to the efficiency of \overline{x} control chart. Advances and Applications in Statistics, 14(1):65–76, January 2010. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/4521.htm.

Garg:2014:CFA

[GS14] Davinder Kumar Garg and Gurinder Pal Singh. Construction of four associate class PBIB designs using pairing in triplets system. Advances and Applications in Statistics, 41(1):11–20, July 2014. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/8626.htm.

Gunasekera:2012:GIC

[Gun12] Sumith Gunasekera. Generalized inferences for the common shape parameter of several Pareto populations. Advances and Applications in Statistics, 26(2):137–152, February 2012. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/6685.htm.

Guo:2011:VSS

[Guo11] Baocai Guo. Variable sample size \overline{X} charts. Advances and Applications in Statistics, 20(2):143–151, February 2011. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/5790.htm.

Gencturk:2018:NMD

[GYH18] Yasemin Gençtürk, Ayten Yiğiter, and Canan Hamurkaroğlu. A new model for data with structural change at threshold: composite exponential-lognormal model. Advances and Applications in Statistics, 53(4):441–456, October 2018. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12115.htm.

Ghanbari:2018:PMB

[GZS18] Saeed Ghanbari, Najaf Zare, and Zahra Shayan. A practical method based on functional data analysis and single exponential smoothing to combine survival curves in meta-analysis: a simulation study. Advances and Applications in Statistics, 53 (2):179–197, August 2018. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/11848.htm.

Hanagal:2008:BEP

[HA08] David D. Hanagal and Kambiz Ahmadi Angali. Bayesian estimation of parameters in some bivariate exponential models. Advances and Applications in Statistics, 10(2):179–193, December 2008. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/3719.htm.

Hady:2019:UED

[Had19] Dina H. Abdel Hady. Use of exponential distribution for hybridization of distributions. Advances and Applications in Statistics, 58(1):57–75, September 2019. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12821.
htm.

Hamed:2019:OSS

[HAE19] Ramadan Hamed, Hanan Aly, and Heba ElDeeb. Optimum step stress accelerated degradation test using goal programming. Advances and Applications in Statistics, 56(1):21–36, May 2019. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/12563.htm.

Haggag:2014:CDR

[Hag14] Magda M. M. Haggag. Combining of dimension reduction regression methods. Advances and Applications in Statistics, 40(2):133–156, June 2014. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/8572.htm.

Haggag:2019:NDM

[Hag19] Magda M. M. Haggag. A new difference method for variance estimation in non-parametric regression models. Advances and Applications in Statistics, 57(2):155–189, August 2019. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12770.htm.

Hajlaoui:2008:MEQ

[Haj08] Mekki Hajlaoui. On multivariate economical quality control. Advances and Applications in Statistics, 8(2):193–207, April 2008. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/3084.htm.

Hamed:2020:MWG

[Ham20a] M. S. Hamed. The mixture Weibull-generalized gamma distribution. Advances and Applications in Statistics, 62(2):139– 171, June 2020. CODEN ???? ISSN 0972-3617. URL http:/

/www.pphmj.com/abstract/13212.htm.

Hamed:2020:SAS

[Ham20b] M. S. Hamed. A single acceptance sampling plan for truncated life test having the (P-A-L) extended Weibull distribution. Advances and Applications in Statistics, 61(2):149–159, April 2020. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/13141.htm.

Han:2003:ASC

[Han03] Kwang-Hee Han. Almost sure convergence for a linear process generated by NA sequence. Advances and Applications in Statistics, 3(2):183–189, August 2003. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
172.htm.

Han:2009:RCE

[Han09] Sun-Ah Han. Recommendation of child education contents in the mobile Internet. Advances and Applications in Statistics, 13(1):41–48, October 2009. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/4203.htm.

Han:2011:DFR

[Han11] Sun Ah Han. Difference of father’s role playing, home environment stimulation, and child’s self-regulation between fatherchildren family and normal family. Advances and Applications in Statistics, 25(1):63–71, November 2011. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
6431.htm.

Hassan:2018:JMC

[Has18] M. Y. Hassan. Joint modeling of correlated time durations and their marks using a Weibull–Poisson marked point process mixture models. Advances and Applications in Statistics, 52 (4):215–234, April 2018. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/11665.htm.

Hassan:2020:DUD

[Has20] Mohamed Yusuf Hassan. An over-dispersed and underdispersed generalized Poisson distribution. Advances and Applications in Statistics, 61(2):183–192, April 2020. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13144.htm.

Harishchandra:2017:SRM

[HB17] K. Harishchandra and Balakrishna. Software reliability model when a random number of faults occurring at a time with correction process. Advances and Applications in Statistics, 50(4):315–327, April 2017. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/10642.htm.

Habib:2004:TCV

[HBA04] Mohamed G. Habib, Adel I. Bargal, and Emad-Eldin A. A. Aly. On testing for a change in variance: new tests and a Monte Carlo study. Advances and Applications in Statistics, 4 (3):303–326, December 2004. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/438.htm.

Hurley:2009:TDP

[HBJ09] W. J. Hurley, Jack Brimberg, and D. L. James. Testing the difference in population means for a Bernoulli-exponential mixture. Advances and Applications in Statistics, 12(2):191–198, August 2009. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/4134.htm.

Hakmanage:2021:ALP

[HCJ21] N. M. Hakmanage, N. V. Chandrasekara, and D. D. M. Jayasundara. Application of linear programming models for optimal resource planning and management in sustainable cultivation — a case study in a farming village of Sri Lanka. Advances and Applications in Statistics, 67(1):21–31, March 2021. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/13710.htm.

Hassan:2016:KWG

[HE16] Amal S. Hassan and M. Elgarhy. Kumaraswamy Weibullgenerated family of distributions with applications. Advances and Applications in Statistics, 48(3):205–239, March 2016. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/9845.htm.

Hafez:2018:JCC

[HEDA18] Nada Hafez, Afaf El-Dash, and Nagwa Albehery. Joint chance constrained programming with dependent parameters. Advances and Applications in Statistics, 53(1):1–12, July 2018. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/11777.htm.

Huang:2002:CAE

[HH02] M. L. Huang and Y. H. Hung. Capability analysis for an entire product with symmetric tolerances. Advances and Applications in Statistics, 2(3):217–231, December 2002. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/123.htm.

Hung:2020:PMA

[HK20] Nguyen Hoang Hung and Tran Tuan Kien. Profitability of moving average convergence divergence and moving average on Vietnamese stock: a comparison. Advances and Applications in Statistics, 60(2):147–168, February 2020. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
13043.htm.

Hervet:2009:ERU

[HKPGN09] Eric Hervet, Mustapha Kardouchi, Thu Pham-Gia, and Francis Niyomwungere. Eye recognition using a statistical approach. Advances and Applications in Statistics, 12(2):207– 221, August 2009. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/4136.htm.

Hsu:2002:ABI

[HLBF02] Chiehwen Ed Hsu, Dejian Lai, Charles Begley, and Luisa Franzini. The association between income and mortality in Texas and Taiwan — a comparative study of spatial statistics. Advances and Applications in Statistics, 2(2):149–176, August 2002. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/110.htm.

Han:2006:ISI

[HLZ06] Fulong Han, Yiqiung Lu, and Riquan Zhang. Inferences on single index models. Advances and Applications in Statistics, 6(2):249–259, August 2006. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/1855.htm.

Harishchandra:2010:MLB

[HM10] K. Harishchandra and K. M. Manjunatha. Maximum likelihood and Bayesian estimation in a software reliability model with exponential fault correction time. Advances and Applications in Statistics, 18(2):175–188, October 2010. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
5405.htm.

Helvaci:2014:RAF

[HM14] Dilcu B. Helvaci and Saeed Maghsoodloo. Renewal and availability functions. Advances and Applications in Statistics, 43 (1):65–89, November 2014. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/8846.htm.

Hasan:2016:NAS

[HNG16] Abeer Hasan, Wei Ning, and Arjun Gupta. A new approach for skew t distribution with applications to environmental data. Advances and Applications in Statistics, 49(2):117–136, August 2016. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/10076.htm.

Hoa:2020:EFA

[Hoa20] Vu Dinh Hoa. Estimation of factors affecting competitiveness of private business in the Song Hong Delta, Vietnam. Advances and Applications in Statistics, 62(2):107–125, June 2020. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/13210.htm.

Huang:2004:NCS

[HP04] Wei-Kwang Huang and Wan-Kai Pang. The noncentral chisquare distribution revisited. Advances and Applications in Statistics, 4(1):59–64, April 2004. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/421.htm.

Hobert:2007:NCD

[HR07] James P. Hobert and Jeffrey S. Rosenthal. Norm comparisons for data augmentation. Advances and Applications in Statistics, 7(2):291–302, August 2007. CODEN ???? ISSN 0972-3617.

Hung:2009:OPE

[HSC09] Chang-Hung Hung, Jung-Chi Su, and Wen-Tzung Chen. Obesity prevalence and an evaluation model for weight control performance. Advances and Applications in Statistics, 11(2): 217–228, April 2009. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/3980.htm.

Hsieh:2008:EDE

[Hsi08] John J. Hsieh. Exact distributions of the empirical generalized survival and distribution functions. Advances and Applications in Statistics, 10(1):85–109, October 2008. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
3502.htm.

Hsieh:2010:ANS

[Hsi10a] John J. Hsieh. Asymptotic normal statistics for Poisson and compound Poisson distributions. Advances and Applications in Statistics, 17(1):61–77, July 2010. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/5166.htm.

Hsieh:2010:DPR

[Hsi10b] John J. Hsieh. On the duality of the Poisson rate. Advances and Applications in Statistics, 16(2):163–190, June 2010. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/5160.htm.

Hsieh:2009:GIA

[HSL09] Hsin-Neng Hsieh, Hsiu-Yuan Su, and Xuewen Lu. A generalized inference on assessment of similarity between dissolution profiles. Advances and Applications in Statistics, 12(2):163– 190, August 2009. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/4133.htm.

Hozzar:2018:ESM

[HTD18] Behnam Hozzar, Ghasem Tohidi, and Behroz Daneshian. On efficient solution of multiobjective linear programming. Advances and Applications in Statistics, 53(5):487–499, November 2018. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/12160.htm.

Huang:2006:NCS

[Hua06] Wei Kwang Huang. On the noncentral chi-square distribution — a computer-aided approach. Advances and Applications in Statistics, 6(1):17–32, April 2006. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/14.htm.

Hung:2005:PFI

[Hun05] Chang-Hung Hung. Physical fitness indices. Advances and Applications in Statistics, 5(3):289–299, December 2005. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/659.htm.

Hurlimann:2003:GBL

[Hür03] Werner Hürlimann. A generalized Benford law and its application. Advances and Applications in Statistics, 3(3):217–228, December 2003. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/177.htm.

Hurlimann:2005:PMD

[Hür05] Werner Hürlimann. Properties and measures of dependence for the Archimax copula. Advances and Applications in Statistics, 5(2):125–143, August 2005. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/448.htm.

Hurlimann:2012:CMS

[Hur12] Werner Hurlimann. On the CVaR measure of skew-symmetric families. Advances and Applications in Statistics, 26(1):25–39, January 2012. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/6618.htm.

Hurlimann:2015:ACI

[Hür15] Werner Hürlimann. Angle correlation inequalities of De Finetti type. Advances and Applications in Statistics, 47(3):175–187, December 2015. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/9557.htm.

Huang:2013:EMC

[HYZ13] Mei Ling Huang, Wai Kong Yuen, and Miao Zhang. Efficient methods on confidence intervals for prediction intervals. Advances and Applications in Statistics, 33(1):1–21, March 2013. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/7454.htm.

Huang:2010:ETP

[HZ10] Mei Ling Huang and Ke Zhao. On estimation of the truncated Pareto distribution. Advances and Applications in Statistics, 16(1):83–102, May 2010. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/5019.htm.

Han:2006:UCR

[HZFZ06] Fulong Han, Riquan Zhang, Jingyan Feng, and Zhiqiang Zhang. Uniform coverage rate of nonparametric kernel estimate under mixing dependent data. Advances and Applications in Statistics, 6(3):353–359, December 2006. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/2219.htm.

Ismail:2015:IMA

[IAED15] Mohamed A. Ismail, Essam A. Ahmed, and Howaida M. Ezz El-Din. Identification of moving average models: a Bayesian approach. Advances and Applications in Statistics, 46(2):79– 96, August 2015. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/9358.htm.

Ismail:2019:RRA

[IAS19] Mohamed A. Ismail, Hend A. Auda, and Mahmoud M. Sadek. Robust recursive analysis of seasonal moving average models. Advances and Applications in Statistics, 54(1):43–67, January 2019. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/12343.htm.

Islam:2014:AAR

[ICBS14] M. A. Islam, R. I. Chowdhury, S. Bae, and K. P. Singh. Assessing the association in repeated measures of depression. Advances and Applications in Statistics, 42(2):83–93, October 2014. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/8723.htm.

Ismail:2005:BIP

[IE05] Mohamed A. Ismail and Tarek I. Eldomiaty. Bayesian identification of the predictors for capital structure in Egypt. Advances and Applications in Statistics, 5(1):1–20, April 2005. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/442.htm.

Irigoien:2006:CCA

[IFVA06] I. Irigoien, E. Fernandez, S. Vives, and C. Arenas. Clapper: a clustering algorithm based on path-distances for experimental replicates. Advances and Applications in Statistics, 6(3):377– 398, December 2006. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/2221.htm.

Ijaz:2020:NRT

[IKA^{+}20] Muhammad Ijaz, Atta Ullah Khan, Syed Muhammad Asim, Yousaf Hayat, Syed Habib Shah, Muhammad Farooq, Kashif Ali, Neelam, and Seema Zubair. A new ratio type estimator to estimate the population mean using auxiliary information. Advances and Applications in Statistics, 63(1):97–108, July 2020. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/13237.htm.

Ingrassia:2009:CSE

[IM09] Salvatore Ingrassia and Isabella Morlini. Computational studies with equivalent degrees of freedoms in neural networks. Advances and Applications in Statistics, 13(1):49–81, October 2009. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/4204.htm.

Inaba:2015:BMC

[IM15] Yosuke Inaba and Etsuo Miyaoka. Bootstrapping method for constructing confidential intervals for estimated parameters of the hidden Markov model. Advances and Applications in Statistics, 46(2):97–105, August 2015. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
9359.htm.

Issa:2020:EAP

[IM20] Mohamed Khalifa Ahmed Issa and Mahmoud Mohamed Abdelwahab Mahmoud. Estimation of AR(1) panel data models with missing observations. Advances and Applications in Statistics, 63(1):109–117, July 2020. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/13238.htm.

Intarasit:2010:GBM

[IS10] A. Intarasit and P. Sattayatham. A geometric Brownian motion model with compound Poisson process and fractional stochastic volatility. Advances and Applications in Statistics, 16(1):25–47, May 2010. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/5015.htm.

Intarapak:2021:FIV

[IS21] Sukanya Intarapak and Thidaporn Supapakorn. Forecasting the import value of Thailand. Advances and Applications in Statistics, 68(1):57–70, May 2021. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/13788.htm.

Iki:2018:MDS

[IT18] Kiyotaka Iki and Sadao Tomizawa. Measure of departure from symmetry using cumulative probability for square contingency tables with ordered categories. Advances and Applications in Statistics, 52(6):401–412, June 2018. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/11721.
htm.

Ijaz:2020:BEP

[IuZHZ20] Muhammad Ijaz, Qamar uz Zaman, Yousaf Hayat, and Zafar Zaheer. Bayesian estimation of the parameter of exponential distribution under Jeffery prior. Advances and Applications in Statistics, 60(2):189–199, February 2020. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/13046.
htm.

Ip:2003:IBT

[IWC03] Wai-Cheung Ip, Heung Wong, and Qinghu Chen. Identification of ballistic texture using wavelet decomposition. Advances and Applications in Statistics, 3(3):191–198, December 2003. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/174.htm.

Jemain:2006:DPR

[JAO06] Abdul Aziz Jemain and Amer Ibrahim Al-Omari. Double percentile ranked set samples for estimating the population mean. Advances and Applications in Statistics, 6(3):261–276, December 2006. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/2213.htm.

Jemain:2007:MPR

[JAO07] Abdul Aziz Jemain and Amer Ibrahim Al-Omari. Multistage percentile ranked set samples. Advances and Applications in Statistics, 7(1):127–139, April 2007. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/2431.htm.

Jaria-Chacon:2011:SPS

[JCACTIGO11] Natalia Jaría-Chacón, Pilar Aparicio-Chueca, Xavier M. Triadó-Ivern, and Joan Guàrdia-Olmos. A study of public spending on higher education institutions in Europe in the framework of the Bologna process using statistical fitting with spatial analysis. Advances and Applications in Statistics, 25 (2):83–102, December 2011. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/6474.htm.

Jaria-Chacon:2015:HEE

[JCGO15] Natalia Jaría-Chacón and Joan Guàrdia-Olmos. A higher education expenditure prediction model: evidence from European countries. Advances and Applications in Statistics, 47(2):153– 174, November 2015. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/9496.htm.

Joshi:2016:NMA

[JD16] Venkatesh B. Joshi and Bhargavi Deshpande. A new modification to the adaptive Nadaraya–Watson kernel regression estimator. Advances and Applications in Statistics, 49(4):245–256, October 2016. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/10223.htm.

Jimenez-Gamero:2010:UKL

[JGAF10] M. D. Jiménez-Gamero and M. V. Alba-Fernández. On the use of the Kullback–Leibler distance for the two-sample problem with mixed data. Advances and Applications in Statistics, 17 (2):161–173, August 2010. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/5235.htm.

Jammalamadaka:2003:NFC

[JK03] S. Rao Jammalamadaka and Tomasz J. Kozubowski. A new family of circular models: the wrapped Laplace distributions. Advances and Applications in Statistics, 3(1):77–103, April 2003. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/162.htm.

Jayawardena:2021:EPS

[JKS21a] J. G. L. S. Jayawardena, U. Anura Kumara, and M. A. K. Sriyalatha. Effectiveness of power sector reforms on social aspect of corporate sustainability of electricity generation in Sri Lanka. Advances and Applications in Statistics, 68(1):43– 56, May 2021. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/13787.htm.

Jayawardena:2021:MRP

[JKS21b] J. G. L. S. Jayawardena, U. Anura Kumara, and M. A. K. Sriyalatha. Measurement of relationship of power sector reforms and social aspect of corporate sustainability of electricity generation of Sri Lanka — exploratory factor analysis. Advances and Applications in Statistics, 66(1):39–60, January 2021. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/13616.htm.

Jitthavech:2015:EUW

[JL15] Jirawan Jitthavech and Vichit Lorchirachoonkul. Estimators using weights based on stratum statistics in stratified random sampling. Advances and Applications in Statistics, 44(2):87– 116, February 2015. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/9019.htm.

J:2019:ERM

[JM19] Hakiim J. and Nor Idayu Mahat. The effects of resampling methods on linear discriminant analysis for data set with two imbalanced groups: an empirical evidence. Advances and Applications in Statistics, 59(1):17–42, November 2019. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12918.htm.

Jakperik:2018:SPM

[JOO18] Dioggban Jakperik, Romanus Odhiambo Otieno, and George Otieno Orwa. A semi-parametric multiplicative bias reduction density with a parametric start. Advances and Applications in Statistics, 53(6):715–729, December 2018. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12233.
htm.

Josephy:2011:PES

[JPW11] N. Josephy, M. Predescu, and S. Woolford. Parameter estimation in a system of stochastic difference equations. Advances and Applications in Statistics, 24(1):49–66, September 2011. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/6189.htm.

Jayakumar:2003:DSL

[JS03] K. Jayakumar and P. C. Sreenivas. On discrete stable laws. Advances and Applications in Statistics, 3(3):255–266, December 2003. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/309.htm.

Jahan:2018:PPM

[JSCI18] Farzana Jahan, Borhan Siddika, Rafiqul I. Chowdhury, and M. Ataharul Islam. A Poisson–Poisson model to analyze congestion data. Advances and Applications in Statistics, 52(1): 55–71, January 2018. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/11484.htm.

Jabeen:2019:RMG

[JSS19] Riffat Jabeen, Saman Hanif Shahbaz, and Muhammad Qaiser Shahbaz. Relations for moments of generalized order statistics for power Lindley distribution. Advances and Applications in Statistics, 56(2):113–124, June 2019. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12638.
htm.

Ji:2014:NSN

[JWLZ14] Ying Ji, Tienan Wang, Yijun Li, and Yong Zhou. A new smoothing nonmonotone trust region method for solving nonlinear complementarity problems. Advances and Applications in Statistics, 40(1):31–60, May 2014. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/8460.
htm.

Jabeen:2020:PEP

[JZ20] Riffat Jabeen and Azam Zaka. Percentiles estimation for the parameters of power function distribution. Advances and Applications in Statistics, 62(2):127–137, June 2020. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13211.htm.

Jaouadi:2017:PSS

[JZTS17] Said Jaouadi, Ilhem Zorgui, Slaheddine Trabelsi, and Rafik Ben Samida. The predictors of students’ satisfaction in Jazan University: evidence from partial least squares. Advances and Applications in Statistics, 50(3):229–242, March 2017. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/10629.htm.

Kim:2006:SVM

[KA06] Tae-Soo Kim and Jung-Ho Ahn. Support vector machine with kernel methods and simulations. Advances and Applications in Statistics, 6(2):207–216, August 2006. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
1849.htm.

Kanwal:2019:WFD

[KA19] Shamsa Kanwal and Munir Ahmad. Weighted Fréchet distribution (area-biased): properties and application. Advances and Applications in Statistics, 59(1):1–15, November 2019. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/12917.htm.

Khan:2020:GAS

[KA20] Khushnoor Khan and Ali Algarni. A group acceptance sampling plan using mean lifetime as a quality parameter for inverse Weibull distribution. Advances and Applications in Statistics, 64(2):237–249, October 2020. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/13440.
htm.

Katatbeh:2008:CRP

[KAA08] Q. D. Katatbeh, M. T. Alodat, and K. M. Aludaat. Upcrossings rate by the product of two Gaussian processes. Advances and Applications in Statistics, 10(2):295–301, December 2008. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/3732.htm.

Khan:2021:SSS

[KAAA21] Khushbakht Khan, Muhammad Riaz Ahmad, Mohammed Albassam, and Muhammad Aslam. A statistical study on safety factors to reduce the road accidents. Advances and Applications in Statistics, 68(1):93–109, May 2021. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
13790.htm.

Khamnei:2012:RSP

[KAF12] Hossein Jabbari Khamnei, Ali Abolhasani, and Parvin Fathipour. Reliability for six parameter generalized Burrtype XII distribution with transformation method. Advances and Applications in Statistics, 29(2):101–110, August 2012. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/7128.htm.

Khan:2018:RTE

[KAK18] Hina Khan, Muhammad Aslam, and Masood Amjad Khan. Ratio type estimators in unequal probability sampling: a numerical study. Advances and Applications in Statistics, 53(4): 457–471, October 2018. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/12116.htm.

Kamel:2011:UPC

[Kam11] Maie M. Kamel. Use of Pareto charts and principle component analysis to determine the main reasons for train accidents in Egypt. Advances and Applications in Statistics, 23(1):29–41, July 2011. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/6078.htm.

Kamel:2019:IME

[Kam19] Maie M. Kamel. Intervention model for exchange rate in Egypt. Advances and Applications in Statistics, 54(1):31–42, January 2019. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/12342.htm.

Kang:2001:AER

[Kan01] Kee-Hoon Kang. On adaptive estimation of regression coefficients with symmetric errors. Advances and Applications in Statistics, 1(2):207–220, August 2001. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/46.htm.

Kang:2016:ABR

[Kan16] Kee-Hoon Kang. An approach for bias reduction in density estimation. Advances and Applications in Statistics, 48(2): 109–121, February 2016. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/9780.htm.

Karuppusamy:2020:IPL

[Kar20] Sadasivan Karuppusamy. Improved Poisson–Lindley distribution. Advances and Applications in Statistics, 65(1):57–68, November 2020. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/13495.htm.

Kawasaki:2010:CEC

[Kaw10] Youhei Kawasaki. Comparison of exact confidence intervals for the difference between two independent binomial proportions. Advances and Applications in Statistics, 15(2):157–170, April 2010. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/4904.htm.

Kawasaki:2014:NCI

[Kaw14] Yohei Kawasaki. A new confidence interval for the difference between two independent binomial proportions. Advances and Applications in Statistics, 38(1):71–80, January 2014. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/8129.htm.

Kharroubi:2011:NFA

[KB11] Samer A. Kharroubi and Alan Brennan. A novel formulation for approximation Bayesian computation based on signed roots of log-density ratios. Advances and Applications in Statistics, 22(1):57–76, May 2011. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/5914.htm.

Khoshnevisan:2005:OPL

[KBS05] M. Khoshnevisan, Sukanto Bhattacharya, and Florentin Smarandache. Optimal plant layout design for process-focused systems. Advances and Applications in Statistics, 5(2):197– 208, August 2005. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/454.htm.

Kallah-Dagadu:2018:FSM

[KDNH18] G. Kallah-Dagadu, B. K. Nkansah, and N. K. Howard. Flexible statistical modelling of the occurrences of transcription factor binding sites along a DNA sequence. Advances and Applications in Statistics, 53(6):659–691, December 2018. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12231.htm.

Kalantan:2019:PEB

[KE19] Zakiah Ibrahim Kalantan and Mervat Khalifah Abd Elaal. Parameters estimations on bivariate Weibull G distributions: a new family. Advances and Applications in Statistics, 58(2): 77–122, October 2019. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/12875.htm.

Kelarev:2005:SAC

[Kel05] A. V. Kelarev. A statistical algorithm for computing information rates of codes with computer algebra systems. Advances and Applications in Statistics, 5(1):87–90, April 2005. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/445.htm.

Kelarev:2006:CSP

[Kel06] A. V. Kelarev. Computing statistics for polynomial codes: an algorithm based on the Mann–Whitney U-test. Advances and Applications in Statistics, 6(1):53–56, April 2006. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/17.htm.

Kurakami:2013:IDP

[KFT13] Hiroyuki Kurakami, Akihiro Fujimura, and Sadao Tomizawa. Incomplete diagonals-parameter symmetry model for ordered square contingency tables. Advances and Applications in Statistics, 33(2):107–113, April 2013. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/7559.
htm.

Kumari:2021:AMR

[KG21] Anupma Kumari and Amisha Gupta. Application of multiple regression model to determine the impact of foreign direct investment and export on economic growth of food processing industry in India. Advances and Applications in Statistics, 68 (1):1–21, May 2021. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/13785.htm.

Khalaf:2013:HRR

[Kha13] Ghadban Khalaf. The helpful ridge regression estimator. Advances and Applications in Statistics, 34(1):39–48, May 2013. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/7682.htm.

Khodabakhshi:2008:UFD

[Kho08] M. Khodabakhshi. Using feasible directions to prove the existence of strong efficient unit in data envelopment analysis. Advances and Applications in Statistics, 10(2):201–206, December 2008. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/3721.htm.

Khamis:2010:SRB

[KHY10] Faisal G. Khamis, Muna F. Hanoon, and Saad G. Yaseen. On a study of the relationship between education and mortality using fully and partially latent models. Advances and Applications in Statistics, 14(2):157–172, February 2010. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/4605.htm.

Kipchirchir:2011:UEP

[Kip11] I. C. Kipchirchir. On ultimate extinction probabilities and mean behaviour of spatial patterns. Advances and Applications in Statistics, 25(1):31–45, November 2011. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
6429.htm.

Kipchirchir:2013:CAD

[Kip13] I. C. Kipchirchir. Comparative analysis of dispersion models. Advances and Applications in Statistics, 37(1):13–35, November 2013. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/7934.htm.

Kipchirchir:2017:CAE

[Kip17] I. C. Kipchirchir. Comparative analysis of the effect of dispersion and competition on population sustainability. Advances and Applications in Statistics, 51(1):17–38, July 2017. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/11015.htm.

Kipchirchir:2020:EED

[Kip20] I. C. Kipchirchir. Estimating an exponentially decaying function of rate parameter of a Poisson process. Advances and Applications in Statistics, 65(1):1–17, November 2020. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13492.htm.

Kabir:2009:EUN

[KK09] Md. Enamul Kabir and Shahjahan Khan. Estimation of univariate normal mean using p-value. Advances and Applications in Statistics, 13(2):241–256, December 2009. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
4451.htm.

Karanina:2018:STD

[KKBB18] S. Yu. Karanina, N. A. Kocheeva, M. Yu. Belikova, and N. V. Baranovskiy. Spatial and temporal distribution of lightning discharges over Tomsk region according to WWLLN data for 2010–2015. Advances and Applications in Statistics, 52(5): 339–362, May 2018. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/11689.htm.

Kokonendji:2009:AKR

[KKD09] Célestin C. Kokonendji, Tristan Senga Kiessé, and Clarice G. B. Demétrio. Appropriate kernel regression on a count explanatory variable and applications. Advances and Applications in Statistics, 12(1):99–125, June 2009. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
4003.htm.

Kossa-Kongbowali:2020:PDS

[KKMMK20] José-Salathiel Kossa-Kongbowali, Boniface Mfourga, Roger Armand Makany, and Gabriel Kissita. Proposal of dual simultaneous co-inertia analysis: Dscia. Advances and Applications in Statistics, 61(1):95–118, March 2020. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/13099.
htm.

Kubrusly:2015:CRI

[KL15] Jessica Kubrusly and Hélio Lopes. Constructive regression on implicit regions. Advances and Applications in Statistics, 45 (3):201–223, June 2015. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/9260.htm.

Kim:2004:CRN

[KLK04] Tae Soo Kim, Joung Nam Lee, and Jong Hyouk Kim. On the combinations of random number generators and empirical tests. Advances and Applications in Statistics, 4(2):155–166, August 2004. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/428.htm.

Koukouvinos:2002:OSD

[KM02] C. Koukouvinos and P. Mantas. E(s^{2})-optimal supersaturated designs constructed from conference matrices. Advances and Applications in Statistics, 2(3):287–300, December 2002. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/132.htm.

Kumar:2008:MSG

[KM08] K. Aneesh Kumar and M. Manoharan. Modelling seed germination — an application of phase type distribution. Advances and Applications in Statistics, 8(1):1–18, February 2008. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/2835.htm.

Koske:2012:CPO

[KMT12] J. K. Koske, J. M. Mutiso, and I. K. Tum. Construction of a practical optimum second order rotatable design in three dimensions. Advances and Applications in Statistics, 30(1): 31–43, September 2012. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/7134.htm.

Kasturiratna:2011:CVN

[KNG11] Dhanuja Kasturiratna, Truc T. Nguyen, and Arjun K. Gupta. Characterization of k-variate normal distribution with covariance structure, σ^{2}Σ_{0} and EDF goodness-of-fit tests. Advances and Applications in Statistics, 20(1):67–88, January 2011. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/5751.htm.

Kissita:2013:CCA

[KNMM13] Gabriel Kissita, Léonard Niere, Roger Armand Makany, and Dominique Mizere. Canonical correlation analysis between two vertical multi-block tables and applications to qualitative variables. Advances and Applications in Statistics, 34(1):11–38, May 2013. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/7681.htm.

Konate:2013:LPC

[Kon13] Lamine Konaté. On LAN for parametrized continuous signals in non-homogeneous diffusion. Advances and Applications in Statistics, 37(1):1–12, November 2013. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
7933.htm.

Kouider:2002:LSE

[Kou02] Elies Kouider. Least-squares estimation of Pareto parameters by judgment ordered ranked-set samples. Advances and Applications in Statistics, 2(2):101–117, August 2002. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/99.htm.

Kiani:2009:ESS

[KPP09] Mehdi Kiani, John Panaretos, and Stelios Psarakis. Extensions to the strategy of the steepest ascent for response surface methodology. Advances and Applications in Statistics, 13 (1):1–17, October 2009. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/4201.htm.

Kourlaba:2008:DAC

[KPS08] G. Kourlaba, D. Panagiotakos, and V. Stavrinos. The diagnostic accuracy of composite indices is associated with the number of partitions of their components: a simulation study. Advances and Applications in Statistics, 8(1):89–99, February 2008. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/2840.htm.

Kim:2001:CLT

[KR01] Tae-Sung Kim and Dae-Hee Ryu. A central limit theorem for stationary linear processes generated by associated process. Advances and Applications in Statistics, 1(1):89–98, April 2001. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/36.htm.

Khosravani:2013:DBD

[KR13] Azar Khosravani and Constantin Rasinariu. n-digit Benford distributed random variables. Advances and Applications in Statistics, 36(2):119–130, October 2013. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/7887.
htm.

Kumar:2009:MGG

[KRS09] Kuldeep Kumar, Gulasekaran Rajaguru, and Samir Shrivastava. Modelling the glitter in gold. Advances and Applications in Statistics, 13(2):225–239, December 2009. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
4450.htm.

Kumar:2003:SES

[KS03] Kuldeep Kumar and Mukesh Srivastava. On the specification of error structure in the linear regression model. Advances and Applications in Statistics, 3(2):105–133, August 2003. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/165.htm.

Kaushal:2018:POU

[KS18] Kshma Kaushal and Sukhdev Singh. Portfolio optimization using wind driven optimization technique. Advances and Applications in Statistics, 53(3):225–241, September 2018. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12013.htm.

Kurosawa:2016:NUM

[KSM16] Takuma Kurosawa, Asanao Shimokawa, and Etsuo Miyaoka. A note on using the maximum partial likelihood estimator of transition model for binary time series. Advances and Applications in Statistics, 48(1):69–90, January 2016. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
9620.htm.

Kokla:2002:MMG

[KT02] Anna Kokla and George Tzavelas. A mathematical model for the Gensini index. Advances and Applications in Statistics, 2 (1):39–50, April 2002. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/68.htm.

Kreuzmair:2008:SRL

[KT08] Ingrid Kreuzmair and Helge Toutenburg. Sensitivity of the restricted least squares estimator under misspecified restrictions and multicollinearity. Advances and Applications in Statistics, 8(2):153–175, April 2008. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/3082.htm.

Kubaychuk:2019:FCA

[Kub19] Oksana Kubaychuk. Fast correction algorithms for weighted empirical distribution functions. Advances and Applications in Statistics, 54(2):315–326, February 2019. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12404.
htm.

Kvaalseth:2011:LDA

[Kvå11] Tarald O. Kvålseth. The lambda distribution and its applications to categorical summary measures. Advances and Applications in Statistics, 24(2):83–106, October 2011. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/6278.htm.

Krzysko:2002:PMD

[KW02] Miroslaw Krzysko and Waldemar Wolynski. Pairwise multiple discriminant analysis. Advances and Applications in Statistics, 2(3):249–268, December 2002. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/126.htm.

Kawasaki:2015:NIT

[KY15] Yohei Kawasaki and Hiroshi Yamada. Non-inferiority test statistics for two independent binomial proportions: comparison of eight methods. Advances and Applications in Statistics, 46(3):183–196, September 2015. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/9426.htm.

Laaksonen:2012:NVU

[Laa12] Seppo Laaksonen. A new visualization for uncertainty intervals with survey examples. Advances and Applications in Statistics, 28(2):133–140, June 2012. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/7082.
htm.

Laaksonen:2014:ISD

[Laa14] Seppo Laaksonen. Introduction to sampling design data file. Advances and Applications in Statistics, 40(2):169–179, June 2014. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/8574.htm.

Lawal:2012:MDS

[Law12] Bayo H. Lawal. Measure of departure from symmetry model for square contingency tables. Advances and Applications in Statistics, 28(2):109–121, June 2012. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/7080.
htm.

Lavalle:2016:ABG

[LB16] Andrea Lavalle and Sergio Bramardi. An algorithm based on generalized Procrustes analysis to find the consensus of several configurations of individuals connected by common checks. Advances and Applications in Statistics, 49(1):31–48, July 2016. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/10025.htm.

Leu:2001:MCS

[LC01] Ching-Ho Leu and K. S. Chen. On a modification of the circular systematic sampling. Advances and Applications in Statistics, 1(3):227–246, December 2001. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/92.htm.

Luengo:2009:NST

[LC09] Amelia V. García Luengo and Inmaculada Oña Casado. A note on sampling over two occasions for the estimation of current population product. Advances and Applications in Statistics, 11(1):87–99, February 2009. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/3973.htm.

Liao:2002:CEP

[LCL02] S. J. Liao, K. S. Chen, and R. K. Li. Capability evaluation for processes of the larger-the-better type for non-normal populations. Advances and Applications in Statistics, 2(2):189–198, August 2002. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/115.htm.

lmaz:2019:HSQ

[lÇl19] Seda Başar Yılmaz, H. Eray Çelik, and Veysel Yılmaz. Healthcare service quality — customer satisfaction: PLS path model. Advances and Applications in Statistics, 54(2):289–300, February 2019. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/12402.htm.

Lin:2005:BRS

[LCP05] Jyh-Jiuan Lin, Ching-Hui Chang, and Nabendu Pal. A brief review of shrinkage estimation of a multivariate normal mean with extension to multiple linear regression. Advances and Applications in Statistics, 5(3):371–399, December 2005. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/668.htm.

Lee:2001:CPG

[Lee01] Seonha Lee. Classification of the person groups using the individual behavior model in the information society. Advances and Applications in Statistics, 1(2):195–205, August 2001. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/45.htm.

Levantesi:2013:SCR

[Lev13] Susanna Levantesi. Solvency capital requirements for longevity risk under different stochastic mortality models. Advances and Applications in Statistics, 33(2):137–160, April 2013. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/7561.htm.

Liang:2001:USN

[LHC01] S. K. Liang, Y. H. Hung, and K. S. Chen. Using the signal-to-noise ratio to evaluate the teller performance in the banking system. Advances and Applications in Statistics, 1(2):133–148, August 2001. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/41.htm.

Li:2011:EPW

[Li11] Jiexiang Li. Estimation of the parameters of Weibull distribution when samples are doubly censored. Advances and Applications in Statistics, 25(1):73–81, November 2011. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/6432.htm.

Li:2012:EPR

[Li12] Jiexiang Li. Estimation of parameters of Rayleigh distributions. Advances and Applications in Statistics, 27(1):1–8, March 2012. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/6738.htm.

Li:2015:IEL

[Li15] Jiexiang Li. Interval estimation for a lifetime distribution with increasing failure rate when samples are doubly censored. Advances and Applications in Statistics, 44(3):177–189, March 2015. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/9121.htm.

Lin:2002:PII

[Lin02] G. H. Lin. Process incapability index for contaminated normal processes. Advances and Applications in Statistics, 2(2):119– 130, August 2002. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/102.htm.

Lin:2009:RDA

[Lin09] Yue-Xian Lin. A rocket discarding algorithm for the valuation of Asian reset options. Advances and Applications in Statistics, 13(1):19–39, October 2009. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/4202.htm.

Liu:2010:BAN

[Liu10] Wei Liu. A Bayesian approach to nonlinear mixed-effects models with measurement errors and missingness in covariates. Advances and Applications in Statistics, 18(1):73–87, September 2010. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/5371.htm.

Loizos:2006:NNN

[LK06] Andreas Loizos and Matthew G. Karlaftis. Neural networks and non-parametric statistical models: a comparative analysis in pavement condition assessment. Advances and Applications in Statistics, 6(1):87–110, April 2006. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/20.htm.

Lim:2004:TEN

[LKP04] Jae-Hak Lim, Dae-Kyoung Kim, and Dong Ho Park. Testing the equality of NBU-ness at a specified age. Advances and Applications in Statistics, 4(1):97–108, April 2004. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/423.htm.

Lee:2009:CIS

[LL09] Seung-Hwan Lee and Eun-Joo Lee. Confidence interval of survival function for extreme value distribution with censored data. Advances and Applications in Statistics, 11(2):239–253, April 2009. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/3982.htm.

Lassina:2017:SBM

[LLDD17] Diabaté Lassina, Kané Ladji, Diawara Daouda, and Diarra Diby. A study of the bonus–malus system in the case of the car insurance sector in Mali. Advances and Applications in Statistics, 51(4):295–307, October 2017. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/11237.
htm.

Latif:2011:NES

[LM11] Sumaia A. Latif and Pedro A. Morettin. Nonparametric estimation of Sibuya’s measure of local dependence for time series. Advances and Applications in Statistics, 23(1):1–27, July 2011. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/6076.htm.

Louzada-Neto:2010:MTS

[LNC10] Francisco Louzada-Neto and Juliana Cobre. A multiple time scale survival model. Advances and Applications in Statistics, 14(1):1–16, January 2010. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/4516.htm.

Lopes:2014:PDC

[Lop14] Kim Samejima M. Lopes. Partial directed coherence applications on EEG data. Advances and Applications in Statistics, 42(1):65–82, September 2014. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/8700.htm.

Li:2011:PEM

[LP11] M. Li and M. Predescu. Parameter estimation in a model of the labor market. Advances and Applications in Statistics, 21 (2):141–158, April 2011. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/5821.htm.

Li:2013:NUM

[LP13] M. Li and M. Predescu. A note on an unemployment model dynamics and estimation of parameters. Advances and Applications in Statistics, 37(2):171–186, December 2013. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/7971.htm.

Lee:2010:SEC

[LPKPS10] Sang Eun Lee, C. J. Park, Yonghee Kim-Park, and Key-Il Shin. Stepwise estimation of cell probabilities in 2 by 2 independent contingency tables. Advances and Applications in Statistics, 16(2):123–139, June 2010. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/5157.
htm.

Liu:2008:NCS

[LPLH08] Wei Liu, Wan-Kai Pang, Ping-Kei Leung, and Shui-Hung Hou. A note on the comparison of several linear regression models. Advances and Applications in Statistics, 9(2):293–301, August 2008. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/3452.htm.

Leu:2007:OSS

[LS07] Lii-Yuh Leu and Kuan-Fu Shen. Optimal step-stress accelerated life test plans with progressive type-II censoring. Advances and Applications in Statistics, 7(1):81–95, April 2007. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/2428.htm.

Luo:2016:EDA

[LS16] Jia Luo and Junfeng Shang. Exploratory data analysis on unemployment rates in USA. Advances and Applications in Statistics, 48(4):303–316, April 2016. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/9854.
htm.

Laipaporn:2020:TVC

[LT20] Jetsada Laipaporn and Phattrawan Tongkumchum. The timevarying correlation estimator using the natural cubic spline volatility. Advances and Applications in Statistics, 62(1):79– 95, May 2020. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/13163.htm.

Lim:2020:CDS

[LTTD20] Apiradee Lim, Muhamad Rifki Taufik, Phattrawan Tongkumchum, and Nurin Dureh. Comparison of different supervised machine learning algorithms for the prediction of tuberculosis mortality. Advances and Applications in Statistics, 62(2): 185–201, June 2020. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/13214.htm.

Lombardo:2009:ANL

[LTvR09] R. Lombardo, G. Tessitore, and J. L. A. van Rijckevorsel. Adaptive non-linear principal component and surface analysis. Advances and Applications in Statistics, 12(1):85–98, June 2009. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/4002.htm.

Luguterah:2016:OGE

[Lug16] Albert Luguterah. Odd generalized exponential Rayleigh distribution. Advances and Applications in Statistics, 48(1):33– 48, January 2016. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/9618.htm.

Li:2018:LRN

[LV18] Na Li and Jan Vrbik. Linear regression with non-normal errors. Advances and Applications in Statistics, 53(2):153–164, August 2018. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/11846.htm.

Lovaglio:2011:EER

[LVM11] P. G. Lovaglio, S. Verzillo, and M. Mezzanzanica. Estimation of educational returns using university and labour market administrative archives. Advances and Applications in Statistics, 23(2):77–105, August 2011. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/6083.htm.

Li:2019:LRM

[LVX19] Na Li, Jan Vrbik, and Xiaojian Xu. Linear regression with mis-specified error distribution. Advances and Applications in Statistics, 54(2):237–253, February 2019. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12399.
htm.

Lau:2004:MSC

[LW04] Salina S. T. Lau and H. Wong. Modelling seasonal conditional heteroscedasticity. Advances and Applications in Statistics, 4 (3):327–355, December 2004. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/439.htm.

Li:2017:CUD

[LWZ^{+}17] Xiaowu Li, Lin Wang, Mingsheng Zhang, Linke Hou, and Juan Liang. On convolution of unimodal distribution and multimodal distribution. Advances and Applications in Statistics, 50(5):385–395, May 2017. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/10825.htm.

Li:2012:BFA

[LX12] Xiangke Li and Dejun Xie. A behavior finance approach for understanding prepayment propensity of mortgage loans. Advances and Applications in Statistics, 29(2):125–134, August 2012. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/7130.htm.

Lee:2003:NNB

[LYK03] Teag-Hee Lee, Bo-Hyun Yun, and Young-Chul Kim. Neural network-based image retrieval using nonlinear combining of heterogeneous features. Advances and Applications in Statistics, 3(3):243–254, December 2003. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/181.htm.

Li:2015:CEO

[LYZ15] Jiaorui Li, Yuan Yao, and Li Zheng. Comprehensive evaluation of operational risks of Chinese public pension fund. Advances and Applications in Statistics, 44(2):117–132, February 2015. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/9020.htm.

Ma:2005:CML

[Ma05] Shuangge Ma. Constraint maximum likelihood estimation of relative risk. Advances and Applications in Statistics, 5(3): 313–323, December 2005. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/662.htm.

Mowafy:2017:SMP

[MA17] Mamdouh Abdel Alim Saad Mowafy and Hanaa Hussein Ali Abou Alela. A statistical model to predict the food gap of the wheat crop in Egypt. Advances and Applications in Statistics, 51(3):205–230, September 2017. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/11135.
htm.

Mesfioui:2019:NOP

[MA19] Mhamed Mesfioui and Ali Algarni. New ordering properties of order statistics from dependent random variables. Advances and Applications in Statistics, 54(1):69–84, January 2019. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12344.htm.

Mohammed:2020:ULR

[MA20] Mohammed Omar Musa Mohammed and Ahmed Saied Rahama Abdallah. Using logistic regression to investigate the risk factors of diabetes in Hotat Bani Tamim Governorate. Advances and Applications in Statistics, 64(2):181–201, October 2020. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/13438.htm.

Mady:2008:BAE

[Mad08] Afaf M. Mady. A Bayesian approach for an exponential parameter with delayed observations. Advances and Applications in Statistics, 10(2):277–284, December 2008. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
3730.htm.

Maghsoodloo:2019:SEN

[Mag19] Saeed Maghsoodloo. On sigma-estimators from a normal universe. Advances and Applications in Statistics, 54(2):199–236, February 2019. CODEN ???? ISSN 0972-3617. URL http:/

/www.pphmj.com/abstract/12398.htm.

Mahmoud:2010:SLM

[Mah10] Marie M. Mahmoud. Statistical logistic model for fitting the relationship between students’ grades and their gender. Advances and Applications in Statistics, 15(2):137–155, April 2010. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/4903.htm.

Maki:2014:CLT

[Mak14] Daiki Maki. A comparison of linearity tests based on wild bootstrap. Advances and Applications in Statistics, 40(2):93– 107, June 2014. CODEN ???? ISSN 0972-3617. URL http:/

/www.pphmj.com/abstract/8570.htm.

Marsh:2013:SAE

[Mar13] Patrick Marsh. Saddlepoint approximations for the estimator on a lagged dependent variable. Advances and Applications in Statistics, 36(2):105–118, October 2013. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
7886.htm.

Marsh:2014:ACD

[Mar14] Patrick Marsh. Approximation for the conditional distribution of the MLE with application to autoregression. Advances and Applications in Statistics, 42(1):15–32, September 2014. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/8697.htm.

Mazmanoglu:2004:VAA

[Maz04] Adnan Mazmanoğlu. Variance analyses of advantages and disadvantages of the service that active financial associations provide their customers solution by the model of hierarchy classification with two factors and their applications. Advances and Applications in Statistics, 4(2):209–222, August 2004. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/432.htm.

Maqsood:2019:ODL

[MBSJ19] Arfa Maqsood, S. M. Aqil Burney, Suboohi Safdar, and Tahseen Ahmed Jilani. Outlier detection in linear time series regression models. Advances and Applications in Statistics, 55 (2):253–268, April 2019. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/12502.htm.

Mutaqin:2010:EMV

[MDTD10] Aceng Komarudin Mutaqin, Sutawanir Darwis, Dumaria Rulina Tampubolon, and Rianto Ahmadi Djojosugito. Estimating mean and variance of sum of correlated random variables using empirical moments. Advances and Applications in Statistics, 16(1):65–81, May 2010. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/5018.htm.

Mahmoud:2014:ENE

[MED14] Osama H. Mahmoud and Eman A. El-Dessouky. Estimating the number excess of loss claims using the exponentiatedexponential distribution. Advances and Applications in Statistics, 38(1):51–69, January 2014. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/8128.htm.

Mekki:2010:GQC

[Mek10] Hajlaoui Mekki. A graphical quality control procedure using data depth. Advances and Applications in Statistics, 19(2): 97–111, December 2010. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/5678.htm.

Mousa:2016:GDG

[MEL16] A. M. Mousa, S. M. El Sayed, and S. H. Abdel Latif. A general delete-group jackknife variance estimator for stratified sampling survey. Advances and Applications in Statistics, 49 (5):327–341, November 2016. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/10271.htm.

Me-ead:2019:CRS

[MeMW19] Cherdchai Me-ead, Rhysa McNeil, and Wandee Wanishsakpong. Clustering regions in South Asia based on factor analysis of temperature anomalies. Advances and Applications in Statistics, 58(1):1–12, September 2019. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12816.
htm.

Mousa:2016:GRB

[MESAF16] Amany M. Mousa, Ahmed A. El-Sheikh, and Mahmoud A. Abdel-Fattah. A gamma regression for bounded continuous variables. Advances and Applications in Statistics, 49(4):305– 326, October 2016. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/10227.htm.

Medina:2020:MER

[MFPO20] Marcos Salazar Medina, José Luis Martinez Flores, Diana Sánchez Partida, and Patricia Cano Olivo. Motor equipment reliability analysis. Advances and Applications in Statistics, 63(2):151–174, August 2020. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/13292.htm.

Mehdi:2019:CBZ

[MFR^{+}19] Tabatabaee Seyed Mehdi, Zayeri Farid, Baghestani Ahmad Reza, Ghorbani Zahra, and Akbarzadeh Baghban Alireza. A comparison between zero and doubly inflated Poisson models based on a simulation study. Advances and Applications in Statistics, 56(1):77–93, May 2019. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/12567.htm.

Marchev:2009:ESS

[MH09] Dobrin Marchev and James P. Hobert. Exact sampling from the Student’s t model. Advances and Applications in Statistics, 13(2):165–179, December 2009. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/4446.htm.

Motta:2012:ISV

[MH12] Anderson C. O. Motta and Luiz K. Hotta. Influence in stochastic volatility models. Advances and Applications in Statistics, 27(1):27–45, March 2012. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/6740.htm.

Mbah:2019:TDS

[MHW^{+}19] Alfred K. Mbah, Wei Wei He, Ronee Wilson, O’Neil Lynch, Karen Bruder, and Hamisu M. Salihu. Transitions in depressive symptoms among pregnant women exposed to environmental tobacco smoke: an application of latent transition analysis. Advances and Applications in Statistics, 55(1):19–34, March 2019. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/12441.htm.

Monia:2011:AFE

[MJ11] Landolsi Monia and Ben Rejeb Jaleleddine. ARIMA forecasting of energy consumption in Tunisia. Advances and Applications in Statistics, 25(2):153–164, December 2011. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/6480.htm.

Mutai:2012:CFD

[MKM12] Charles K. Mutai, Joseph K. Koske, and John M. Mutiso. Construction of a four dimensional third order rotatable design through balanced incomplete block designs. Advances and Applications in Statistics, 27(1):47–53, March 2012. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/6741.htm.

Morrell:2015:EPC

[MKM^{+}15] Christopher H. Morrell, Michael W. Kattan, F. Roy MacKintosh, Stephen Van Den Eeden, and Thomas B. Neville. Estimating prostate cancer relative survival and cancer-specific death. Advances and Applications in Statistics, 46(3):197–223, September 2015. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/9427.htm.

Mutiso:2016:CFL

[MKN16] J. M. Mutiso, G. K. Kerich, and H. M. Ng’eno. Construction of five level second order rotatable designs using supplementary difference sets. Advances and Applications in Statistics, 49 (1):21–30, July 2016. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/10024.htm.

Mecha:2021:GDM

[MKO21] Peter N. Mecha, Isaac C. Kipchirchir, and Joseph A. M. Ottieno. Generators for discrete mixtures based on order statistics with application to exponential distribution. Advances and Applications in Statistics, 67(2):161–178, April 2021. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13757.htm.

Mun:2002:TDP

[MM02] Yong Mun and Jisun Mun. A time-dependent Poisson process approach to chain smoking behavior. Advances and Applications in Statistics, 2(2):177–187, August 2002. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
113.htm.

Mansour:2021:ISE

[MME21] Mahmoud M. Mansour, Salah M. Mohamed, and Enayat M. Abd Elrazik. An improved strategy to estimate the rare sensitive parameter under Poisson distribution by using twostage unrelated question RRT model. Advances and Applications in Statistics, 67(1):85–116, March 2021. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
13714.htm.

Mak:2015:PAE

[MN15] T. K. Mak and F. Nebebe. A parametric approach for estimating conditional probability distributions. Advances and Applications in Statistics, 44(2):133–143, February 2015. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/9021.htm.

Mak:2018:ENS

[MN18] T. K. Mak and F. Nebebe. Estimating nonlinear structural relationships. Advances and Applications in Statistics, 52(5): 327–338, May 2018. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/11688.htm.

Mak:2020:SLE

[MN20] Tak K. Mak and Fassil Nebebe. Simulated likelihood estimation using iterative importance sampling. Advances and Applications in Statistics, 60(1):45–62, January 2020. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12996.htm.

Maiti:2017:SML

[MNC17] Suchismita Maiti, Debashis Nandi, and Baisakhi Chakraborty. Statistical modelling of log-compressed Nakagami distributed ultrasound echo envelope. Advances and Applications in Statistics, 50(5):349–366, May 2017. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/10823.
htm.

Moala:2010:BAW

[Moa10] Fernando A. Moala. Bayesian analysis for the Weibull parameters by using noninformative prior distributions. Advances and Applications in Statistics, 14(2):117–143, February 2010. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/4603.htm.

Mohamed:2013:CBM

[Moh13] Rania Ahmed Hamed Mohamed. A comparison between the models ARIMA/GARCH and artificial neural networks in modeling financial returns to the Arab Republic of Egypt market securities. Advances and Applications in Statistics, 36(2): 131–150, October 2013. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/7888.htm.

Mohamed:2014:CCL

[Moh14] Rania Ahmed Hamed Mohamed. Classification of the companies listed on the Egyptian stock using SOM neural network, K-means and hierarchical clustering methods. Advances and Applications in Statistics, 38(2):149–159, February 2014. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/8174.htm.

Mostafa:2015:BAT

[Mos15] Ayman A. Mostafa. Bayesian analysis technique for generalized Cox’s proportional hazards model using bugs: applications in medical data. Advances and Applications in Statistics, 46(1):1–36, July 2015. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/9330.htm.

Menendez:2009:PTS

[MPZ09] M. L. Menéndez, L. Pardo, and K. Zografos. Preliminary test and Stein-type estimation of location parameter for elliptically contoured distributions. Advances and Applications in Statistics, 11(1):101–136, February 2009. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/3974.htm.

Mettle:2021:QMS

[MQAA21] Felix Okoe Mettle, Enoch NiiBoi Quaye, Dennis Arku, and Emmanuel Kojo Aidoo. A quantitative model for a sustainable pension plan in a developing economy. Advances and Applications in Statistics, 67(1):47–64, March 2021. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
13712.htm.

Mettle:2019:ASP

[MQLA19] Felix O. Mettle, Enoch N. B. Quaye, Ravenhill A. Laryea, and Louis Asiedu. Analysis of share prices as Markov chains with countably infinite states. Advances and Applications in Statistics, 54(1):1–20, January 2019. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12340.
htm.

Morlanes:2009:EEA

[MRS09] José Igor Morlanes, Antti Rasila, and Tommi Sottinen. Empirical evidence on arbitrage by changing the stock exchange. Advances and Applications in Statistics, 12(2):223–233, August 2009. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/4137.htm.

Md:2019:ACD

[MRS19] Kamruzzaman Md, Md. Arafat Rahman, and Muhammad Shuaib. Analysis of count data fitting GEE model: an application to adolescent behavioral monitoring data. Advances and Applications in Statistics, 54(1):137–149, January 2019. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/12347.htm.

Mata:2010:FHM

[MRSR10] M. C. Juan Antonio González Mata, M. C. Ramón Luévanos Rojas, M. C. Cesar Uranga Sifuentes, and Arnulfo Luévanos Rojas. Forecasting of human migration flow in the Comarca Lagunera through the application of Markov processes. Advances and Applications in Statistics, 18(1):1–26, September 2010. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/5367.htm.

Mandal:2006:BMM

[MS06] S. Mandal and M. Samanta. O-BLUE and minimum MSE linear estimator of the location parameter of an exponential distribution with known coefficient of variation. Advances and Applications in Statistics, 6(2):121–128, August 2006. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/1846.htm.

Miyamoto:2005:NDS

[MSIT05] Nobuko Miyamoto, Satoshi Shinohara, Akira Inoue, and Sadao Tomizawa. A note on different selection of best-fitting model by likelihood ratio test and Akaike information criterion for the analysis of contingency tables. Advances and Applications in Statistics, 5(3):301–312, December 2005. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/660.htm.

Morrell:2013:TPS

[MSP^{+}13] Christopher H. Morrell, Veena Shetty, Terry Phillips, Thiruma V. Arumugam, Mark P. Mattson, and Ruiqian Wan. Testing the product of slopes in related regressions. Advances and Applications in Statistics, 36(1):29–46, September 2013. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/7904.htm.

Muller:2002:FBT

[Mül02] Matthias J. Müller. Finite biprediction tests in treatmentresponse contingency tables. Advances and Applications in Statistics, 2(1):1–18, April 2002. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/57.htm.

Mun:2011:ICF

[Mun11] Yong Mun. Issues of controlling family-wise type-I error rate (FWER) when simultaneously evaluating non-inferiority and superiority using sequence testing methods. Advances and Applications in Statistics, 23(1):53–57, July 2011. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
6080.htm.

Muzzioli:2012:PCP

[Muz12] S. Muzzioli. Put-call parity and options’ forecasting power. Advances and Applications in Statistics, 30(1):1–17, September 2012. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/7132.htm.

Maghsoodloo:2017:SSV

[MWB17] Saeed Maghsoodloo, Randy West, and Dilcu Barnes. A statistical study on the variability of aggregate characteristics of asphalt mixtures. Advances and Applications in Statistics, 51 (6):427–444, December 2017. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/11387.htm.

Mandal:2015:ODP

[MY15] S. Mandal and Y. Yang. D_{s}-optimal designs in polynomial regression models. Advances and Applications in Statistics, 45 (3):167–179, June 2015. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/9258.htm.

Nagar:2005:PHF

[NA05] Daya K. Nagar and José Angel Alvarez. Properties of the hypergeometric function type I distribution. Advances and Applications in Statistics, 5(3):341–351, December 2005. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/666.htm.

Nagar:2004:NCC

[NBA04] Daya K. Nagar, Elizabeth Bedoya, and Elkin Lubin Arias. Non-central complex matrix-variate beta distribution. Advances and Applications in Statistics, 4(3):287–302, December 2004. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/437.htm.

Nandram:2015:BAR

[NBL15] Balgobin Nandram, Dhiman Bhadra, and Yiwei Liu. A Bayesian analysis of racial differences in treatment among breast cancer patients. Advances and Applications in Statistics, 45(2):75–104, May 2015. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/9174.htm.

Nagar:2003:TMC

[NC03] Daya K. Nagar and María Eugenia Castañeda. Testing multisample compound symmetry. Advances and Applications in Statistics, 3(3):285–292, December 2003. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
313.htm.

Navaei:2017:SPP

[ND17] L. Navaei and R. Dayanian. On some properties of a pseudoprobability space. Advances and Applications in Statistics, 50 (2):85–95, February 2017. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/10579.htm.

Ning:2008:DUI

[Nin08] Wei Ning. Detecting an unconditionally identifiable pattern in two-way ANOVA. Advances and Applications in Statistics, 9(2):247–260, August 2008. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/3449.htm.

Nishiyama:2009:MCA

[Nis09] Takahiro Nishiyama. On the multiple comparisons among the components of the mean vector in repeated measure design.

Advances and Applications in Statistics, 11(2):255–272, April 2009. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/3983.htm.

Niwitpong:2017:GCI

[Niw17] Suparat Niwitpong. Generalized confidence intervals for function of variances of lognormal distributions. Advances and Applications in Statistics, 51(2):151–163, August 2017. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/11081.htm.

Nishino:2013:BWE

[NK13] Haruhisa Nishino and Kazuhiko Kakamu. Bayesian Whittle estimation of ARFIMA model. Advances and Applications in Statistics, 37(2):149–170, December 2013. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
7970.htm.

Niwitpong:2011:CIR

[NKaN11] Suparat Niwitpong, Sanoe Koonprasert, and Sa aat Niwitpong. Confidence intervals for the ratio of normal means with a known coefficient of variation. Advances and Applications in Statistics, 25(1):47–61, November 2011. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
6430.htm.

Nkodia:2009:APE

[Nko09] Merlin Guy Nkodia. Amortization possibilities in the explosive evolution of poverty in Congo–Brazzaville: an econometrical approach. Advances and Applications in Statistics, 11(1):65– 85, February 2009. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/3972.htm.

Nkodia:2010:ILS

[Nko10] Merlin Guy Nkodia. Impact of the labour supply on an economic system: an econometrical approach on Congolese data. Advances and Applications in Statistics, 18(2):127–160, October 2010. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/5403.htm.

Nidsunkid:2021:PCC

[NL21] Sudarat Nidsunkid and Mena Lao. The performance of control chart for individual measurements when the process data are uniformly distributed. Advances and Applications in Statistics, 67(1):65–83, March 2021. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/13713.htm.

Nasiru:2016:GET

[NLI16] Suleman Nasiru, Albert Luguterah, and Mohammed Muniru Iddrisu. Generalized Erlang-truncated exponential distribution. Advances and Applications in Statistics, 48(4):273–301, April 2016. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/9853.htm.

Nag:2004:TSL

[NM04] Ashok K. Nag and Amit Mitra. Time series lag selection using genetic algorithm. Advances and Applications in Statistics, 4 (2):233–251, August 2004. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/434.htm.

Nasiru:2018:EGH

[NMN18] Suleman Nasiru, Peter N. Mwita, and Oscar Ngesa. Exponentiated generalized half logistic Burr X distribution. Advances and Applications in Statistics, 52(3):145–169, March 2018. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/11562.htm.

Nduati:2020:AKI

[NMW20] Dorcas W. Nduati, Levi N. Mbugua, and Moses W. Wakoli. Application of Kumaraswamy inverse Weibull distribution in modeling survival data. Advances and Applications in Statistics, 64(1):1–22, September 2020. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/13331.htm.

Navarro:2005:NDN

[NR05] Jorge Navarro and José M. Ruiz. A note on the discrete normal distribution. Advances and Applications in Statistics, 5(2): 229–245, August 2005. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/558.htm.

Noskov:2020:ACD

[NV20] S. I. Noskov and A. S. Vergasov. Application criterion of displacement when choosing the method of evaluating the parameters of regression equations. Advances and Applications in Statistics, 61(1):89–94, March 2020. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/13098.
htm.

Niwitpong:2016:CIDa

[NW16] Suparat Niwitpong and Arunee Wongkhao. Confidence intervals for the difference between inverse of normal means. Advances and Applications in Statistics, 48(5):337–347, May 2016. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/9912.htm.

Nwabueze:2009:PER

[Nwa09] Joy Chioma Nwabueze. Pattern of exchange rate movement of seven underdeveloped countries. Advances and Applications in Statistics, 13(2):193–204, December 2009. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
4448.htm.

Oukid:2009:BBP

[OA09] Nadia Oukid and Amar Aissani. Bounds on busy period for queues with breakdowns. Advances and Applications in Statistics, 11(2):137–156, April 2009. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/3975.htm.

Okagbue:2019:CFE

[OAA19] Hilary I. Okagbue, Muminu O. Adamu, and Timothy A. Anake. Closed form expression of the quantile function of Maxwell–Boltzmann distribution. Advances and Applications in Statistics, 54(2):179–197, February 2019. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
12397.htm.

Oida:2011:EVB

[Oid11] Kazumasa Oida. Enhanced variance-based suspended stream detection. Advances and Applications in Statistics, 24(1):7– 33, September 2011. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/6187.htm.

Okango:2016:OGC

[OKM16] Ayubu Anapapa Okango, Joseph Kipsigei Koske, and John Muindi Mutiso. Optimization of germination conditions of Melia volkensii by response surface methodology. Advances and Applications in Statistics, 49(3):195–210, September 2016. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/10079.htm.

Omara:2018:WLW

[Oma18] Tarek M. Omara. Weight LAD and weight LAD ridge estimator for seemingly unrelated regression models. Advances and Applications in Statistics, 52(6):375–389, June 2018. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/11719.htm.

Omara:2019:MLT

[Oma19] Tarek M. Omara. Modification of Liu-type estimator for two SUR model. Advances and Applications in Statistics, 55(1): 47–66, March 2019. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/12443.htm.

Ogola:2019:GIW

[OMO19] Anthony C. Ogola, Levi N. Mbugua, and George O. Orwa. The gamma inverse Weibull Poisson distribution with application to survival data. Advances and Applications in Statistics, 59 (2):149–172, December 2019. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/12955.htm.

Odiase:2007:EPP

[OO07] J. I. Odiase and S. M. Ogbonmwan. Exact permutation paradigm in multisample experiments. Advances and Applications in Statistics, 7(3):417–434, December 2007. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/2721.htm.

Omoregie:2020:FNP

[OO20] David E. Omoregie and Patrick Osatohanmwen. Forecasting Nigeria’s per capita disposable income using the SARIMA time series framework. Advances and Applications in Statistics, 64(1):23–32, September 2020. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/13332.htm.

Olatayo:2011:RBM

[OOA11] T. O. Olatayo, A. I. Oredein, and A. A. Akomolafe. Residual bootstrap method: an approach to model validation. Advances and Applications in Statistics, 25(2):109–113, December 2011. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/6476.htm.

Omidi:2017:RSA

[ORMPL17] Mostafa Omidi, Mohsen Rostamy-Malkhalifeh, Ali Payan, and Farhad Hosseinzadeh Lotfi. Radial sensitivity analysis of left, right, and overall returns to scale using the virtual anti-ideal unit. Advances and Applications in Statistics, 50(2):137–171, February 2017. CODEN ???? ISSN 0972-3617. URL http:/

/www.pphmj.com/abstract/10583.htm.

Oluyede:2007:WCV

[OT07] Broderick O. Oluyede and Mekki Terbeche. On weighted covariance and vitality orderings with application to the exponential family. Advances and Applications in Statistics, 7(2): 175–192, August 2007. CODEN ???? ISSN 0972-3617.

Ozturk:2001:RER

[Özt01] Ömer Öztürk. Robust estimation of regression and scale parameters in linear models. Advances and Applications in Statistics, 1(1):59–87, April 2001. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/35.htm.

Pintar:2013:PBM

[PACW13] Adam L. Pintar, Christine M. Anderson-Cook, and Huaiqing Wu. Prediction-based model selection for Bayesian multiple regression models. Advances and Applications in Statistics, 32 (2):83–117, February 2013. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/7441.htm.

Pappas:2011:TPL

[PAL11] Vasileios Pappas, Konstantinos Adamidis, and Sotirios Loukas. A three-parameter lifetime distribution. Advances and Applications in Statistics, 20(2):159–167, February 2011. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/5792.htm.

Pararai:2010:UGP

[Par10] Mavis Pararai. Underreporting in generalized Poisson regression model with an application to demand for medical care data. Advances and Applications in Statistics, 16(1):1–15, May 2010. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/5013.htm.

Peschell:2004:TEA

[PBS^{+}04] K. J. Peschell, S. Bae, R. Shanmugam, A. A. Bartolucci, and K. P. Singh. Therapeutic equivalence assessment (TEA) for exponential data. Advances and Applications in Statistics, 4 (1):1–14, April 2004. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/317.htm.

Pailden:2013:ELR

[PC13] Junvie Pailden and Hanfeng Chen. Empirical likelihood ratio test for difference of the means of zero-inflated populations. Advances and Applications in Statistics, 35(1):29–55, July 2013. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/7783.htm.

Paneru:2014:RAU

[PC14] Khyam Paneru and Hanfeng Chen. Regression analysis under complex probability sampling designs in presence of many zero-value responses. Advances and Applications in Statistics, 40(1):1–29, May 2014. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/8459.htm.

Panityakul:2018:AMA

[PC18] Thammarat Panityakul and Ronnason Chinram. Applied meta-analysis for teacher research in southern border provinces of Thailand. Advances and Applications in Statistics, 52(6):391–400, June 2018. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/11720.htm.

Percontini:2013:NBF

[PCB13] Ana Percontini, Gauss M. Cordeiro, and Marcelo Bourguignon. The G-negative binomial family: general properties and applications. Advances and Applications in Statistics, 35 (2):127–160, August 2013. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/7846.htm.

Pero-Cebollero:2018:SRS

[PCGOMF18] Maribel Peró-Cebollero, Joan Guàrdia-Olmos, and Núria Mancho-Fora. A systematic review of simulation procedures for FMRI connectivity studies. Advances and Applications in Statistics, 53(1):29–69, July 2018. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/11779.htm.

Pero-Cebollero:2011:CCI

[PCSMGO11] Maribel Peró-Cebollero, Juan Delgado Sánchez-Mateos, and Joan Guàrdia-Olmos. Comparison of confidence intervals based on the median as a robust alternative to the classic hypotheses test. Advances and Applications in Statistics, 24(1): 67–82, September 2011. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/6190.htm.

Panagiotakos:2007:CTD

[PD07] Demosthenes B. Panagiotakos and George Dedoussis. Classification trees in determining health outcomes, using genetic information. Advances and Applications in Statistics, 7(2): 303–311, August 2007. CODEN ???? ISSN 0972-3617.

Panagiotakos:2009:ESR

[PD09] Demosthenes B. Panagiotakos and Urania Dafni. Evaluation of stopping rules for stochastic curtailment. Advances and Applications in Statistics, 12(1):25–36, June 2009. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
3999.htm.

Paraguas:2006:SLA

[PDK06] Ferdinand J. Paraguas, Madan Mohan Dey, and Anton Abdulbasah Kamil. Spatial linear approximate to almost ideal demand system. Advances and Applications in Statistics, 6 (3):361–375, December 2006. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/2220.htm.

Pararai:2006:GPR

[PFL06] Mavis Pararai, Felix Famoye, and Carl Lee. Generalized Poisson regression model for underreported counts. Advances and Applications in Statistics, 6(3):305–322, December 2006. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/2216.htm.

Pacella:2010:PCE

[PG10] Massimo Pacella and Antonio Grieco. Production cost estimation by linear regression analysis: a case study of model selection in manufacturing. Advances and Applications in Statistics, 16(2):103–121, June 2010. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/5156.htm.

Pang:2007:BAE

[PHTY07] Wan Kai Pang, Shui Hung Hou, Marvin D. Troutt, and Bosco Wing Tong Yu. A Bayesian approach to the estimation of Pareto distribution with an application in insurance. Advances and Applications in Statistics, 7(3):389–401, December 2007. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/2719.htm.

Pirouz:2016:CAB

[PK16] Behzad Pirouz and Esmaile Khorram. A computational approach based on the e-constraint method in multi-objective optimization problems. Advances and Applications in Statistics, 49(6):453–483, December 2016. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/10382.
htm.

Pandit:2017:TAP

[PK17] Parameshwar V. Pandit and Savitha Kumari. On testing against positive quadrant dependence based on sub-sample order statistics. Advances and Applications in Statistics, 50(2): 107–121, February 2017. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/10581.htm.

Pourahmad:2016:OCS

[PKKHM^{+}16] Saeedeh Pourahmad, Mahmood Kharrati-Kopaei, Zahra Hosseini-Moghadam, Sadat Nouri, and Mojgan Akbarzadeh Jahromi. An optimal combination of several binary markers based on Neyman–Pearson lemma: with an application to immunohistochemical markers for the diagnosis of classical Hodgkin lymphoma. Advances and Applications in Statistics, 48(5):369–389, May 2016. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/9915.htm.

Piri:2018:SPW

[PLRaB18] Maliheh Piri, Farhad Hosseinzadeh Lotfi, Mohsen Rostamyalkhalifeh, and Mohammad Hasan Behzadi. The selection of projects within constrained resources using stochastic data envelopment analysis. Advances and Applications in Statistics, 52(3):171–201, March 2018. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/11563.htm.

Prado:2015:DSM

[PLRV15] Silvia Maria Prado, Francisco Louzada, José Gilberto Rinaldi, and Marcio Luis Lanfredi Viola. A distribution for service model. Advances and Applications in Statistics, 44(1):1–19, January 2015. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/9014.htm.

Punathumparambath:2021:NAE

[PM21] Bindu Punathumparambath and Kannan Vadakkadath Meethal. New approach for estimating differentially expressed genes in microarray. Advances and Applications in Statistics, 66(2): 191–208, February 2021. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/13670.htm.

Porto:2015:TIM

[PMS15] Rogério F. Porto, Oscar E. Molina, and Gladys E. Salcedo. A tv-ivar model for multivariate irregular time series. Advances and Applications in Statistics, 45(3):181–200, June 2015. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/9259.htm.

Polymenis:2019:NAD

[Pol19] Athanase Polymenis. A note on the asymptotic distribution of Student’s t for AR(1) processes. Advances and Applications in Statistics, 56(1):37–42, May 2019. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/12564.htm.

Panagiotakos:2002:AAS

[PPT02] D. B. Panagiotakos, C. Pitsavos, and G. Tzavelas. An attempt to analyze series of epidemiological data using multiple correspondence analysis: an application to cardio2000 database. Advances and Applications in Statistics, 2(2):139–147, August 2002. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/107.htm.

Palmeros-Rojas:2014:ECP

[PRVATS14] Oscar Palmeros-Rojas, José A. Villaseñor-Alva, and Francisco S. Tajonar-Sanabria. Estimation of a change-point in the Weibull regression model. Advances and Applications in Statistics, 43(2):107–118, December 2014. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
8860.htm.

Phantu:2018:DCM

[PSA18] Suganya Phantu, Saowanit Sukparungsee, and Yupaporn Areepong. Dma chart monitoring of the first integer-valued autoregressive processes of Poisson counts. Advances and Applications in Statistics, 52(2):97–119, February 2018. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/11505.htm.

Pouresmaeil:2017:EMU

[PSKF17] Hassan Pouresmaeil, Elyas Shivanian, Esmaile Khorram, and Hassan Salehi Fathabadi. An extended method using TOP-SIS and VIKOR for multiple attribute decision making with multiple decision makers and single valued neutrosophic numbers. Advances and Applications in Statistics, 50(4):261–292, April 2017. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/10640.htm.

Pavlenko:2004:OWH

[PvR04] Tatjana Pavlenko and Dietrich von Rosen. On the optimal weighting of high-dimensional Bayesian networks. Advances and Applications in Statistics, 4(3):357–377, December 2004. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/440.htm.

Qumsiyeh:2017:BAT

[QDG17] Maher Qumsiyeh, Robert Deis, and Dalton Gannon. Bootstrapping an autoregressive time series model using SAS. Advances and Applications in Statistics, 50(6):435–467, June 2017. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/10940.htm.

Qyrana:2020:FSM

[QF20] Mishel Qyrana and Silvia Figini. Feature selection and multicollinearity in supervised models. Advances and Applications in Statistics, 65(2):151–164, December 2020. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
13571.htm.

Qureshi:2018:GEP

[QH18] Naureen Riaz Muhammad Nouman Qureshi and Muhammad Hanif. Generalized estimator for population variance in systematic random sampling. Advances and Applications in Statistics, 53(6):771–788, December 2018. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12235.
htm.

Qureshi:2018:JIE

[QKH18] Muhammad Nouman Qureshi, Sadia Khalil, and Muhammad Hanif. Joint influence of exponential ratio and exponential product estimator for the estimation clustered population mean in adaptive cluster sampling. Advances and Applications in Statistics, 53(1):13–28, July 2018. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/11778.htm.

Qin:2004:ELR

[QTH04] Gengsheng Qin, Min Tsao, and Yu-Sheng Hsu. Empirical likelihood ratio confidence interval for the Winsorized mean. Advances and Applications in Statistics, 4(2):167–180, August 2004. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/429.htm.

Rashid:2021:SWP

[RA21] Muhammad Zahid Rashid and Ahmad Saeed Akhter. Survival weighted power function distribution with applications to medical, oceanology and metrology data. Advances and Applications in Statistics, 67(2):133–160, April 2021. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13756.htm.

Rattan:2015:SRO

[RAM15] Preety Rattan, Sangeeta Arora, and Kalpana K. Mahajan. Some results on ordering of Lorenz and related curves for power function distribution. Advances and Applications in Statistics, 44(1):21–41, January 2015. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/9015.
htm.

Ramadan:2016:EMG

[Ram16a] Saleem Z. Ramadan. An efficient method for global optimization of black-box functions using one-dimensional interpolation and reduced design space. Advances and Applications in Statistics, 48(1):1–31, January 2016. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/9617.htm.

Ramadan:2016:NPS

[Ram16b] Saleem Z. Ramadan. A note on planning the simple stepstress accelerated life test under Bayesian method and Type-I right censoring. Advances and Applications in Statistics, 48 (3):185–204, March 2016. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/9844.htm.

Ramadan:2016:STD

[Ram16c] Saleem Z. Ramadan. Sampling target distributions based on support alteration and discretization. Advances and Applications in Statistics, 48(6):445–471, June 2016. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
10000.htm.

Rasch:2011:DSE

[Ras11] Dieter Rasch. Determining the size of experiments for ANOVA models. Advances and Applications in Statistics, 21(2):85–104, April 2011. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/5818.htm.

Rashwan:2013:LBV

[Ras13] Nasr Ibrahim Rashwan. A length-biased version of the generalized gamma distribution. Advances and Applications in Statistics, 32(2):119–137, February 2013. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/7442.htm.

Rashwan:2015:EPM

[Ras15] Nasr Ibrahim Rashwan. Estimation of the parameters of mixed exponentiated Weibull and exponentiated exponential from censored type I samples. Advances and Applications in Statistics, 47(1):1–18, October 2015. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/9469.htm.

Reschenhofer:2003:ACS

[RB03] Erhard Reschenhofer and Lakdere Benkherouf. Automatic continuous smoothing. Advances and Applications in Statistics, 3(1):15–32, April 2003. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/156.htm.

Reyes-Cervantes:2019:CCI

[RCAAMC19] Hortensia Reyes-Cervantes, Félix Almendra-Arao, and Marcos Morales-Cortés. A comparison of confidence intervals for a proportion and criteria for their application. Advances and Applications in Statistics, 58(1):35–43, September 2019. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12819.htm.

Rodrigues:2011:WNB

[RCDO11] Cristiane Rodrigues, Gauss M. Cordeiro, Clarice G. B. Demétrio, and Edwin M. M. Ortega. The Weibull negative binomial distribution. Advances and Applications in Statistics, 22(1):25–55, May 2011. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/5913.htm.

Riad:2018:NGF

[RE18] Mahmoud Riad and Ola Farouk Abd El Mabood. A new goodness of fit test for the beta distribution based on the empirical Laplace transform. Advances and Applications in Statistics, 53(2):165–177, August 2018. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/11847.htm.

Reschenhofer:2004:SSB

[Res04] Erhard Reschenhofer. On subset selection and beyond. Advances and Applications in Statistics, 4(3):265–286, December 2004. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/436.htm.

Robinson:2015:ACE

[RG15] R. A. Robinson and R. S. Gill. Asymptotics and confidence estimation in segmented regression models. Advances and Applications in Statistics, 45(2):125–151, May 2015. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
9176.htm.

Rondero-Guerrero:2020:GUD

[RGlHnSC20] Carlos Rondero-Guerrero, Isidro GonzÃlez HernÃ ndez, and Carlos Soto-Campos. On a generalized uniform distribution. Advances and Applications in Statistics, 60(1):93–103, January 2020. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/12999.htm.

Rashad:2020:CEM

[RIH20] Muhammad Rashad, Zafar Iqbal, and Muhammad Hanif. Characterizations and entropy measures of the Libby–Novick generalized beta distribution. Advances and Applications in Statistics, 63(2):235–259, August 2020. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/13297.
htm.

Raziq:2018:DAF

[RIT18] Abdul Raziq, Raja Muhammad Ilyas, and Ghulam Hyder Talpur. Differences among factors of job satisfaction in classification of the academic institute (school, college and university) teachers of Balochistan, Pakistan. Advances and Applications in Statistics, 53(1):71–85, July 2018. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/11780.
htm.

Ramli:2020:MVF

[RJ20] Suhailywati Ramli and Saiful Hafizah Jaaman. Medianvariance fuzzy with trapezoidal fuzzy numbers for portfolio selection model. Advances and Applications in Statistics, 60 (1):35–44, January 2020. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/12995.htm.

Rasheed:2014:MQS

[RM14] K. V. Abdul Rasheed and M. Manoharan. Markovian queueing system with self regulatory servers. Advances and Applications in Statistics, 41(2):109–128, August 2014. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
8693.htm.

Rahman:2018:MLP

[RM18] Mezbahur Rahman and Sabiha Mahzabeen. Maximum likelihood parameter estimation for the mixtures of normal distributions. Advances and Applications in Statistics, 53(5):501– 518, November 2018. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/12161.htm.

Ramos:2013:ELP

[RMdSC13] Manoel Wallace A. Ramos, Pedro Rafael D. Marinho, Ronaldo V. da Silva, and Gauss M. Cordeiro. The exponentiated Lomax Poisson distribution with an application to lifetime data. Advances and Applications in Statistics, 34(2): 107–135, June 2013. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/7663.htm.

Rojas:2011:VSP

[RMS^{+}11] Arnulfo Luévanos Rojas, Facundo Cortés Martinez, Cesar Uranga Sifuentes, Ramon Luévanos Rojas, and Ma Enriqueta Luévanos Soto. Vibrations in systems of pipes with different movements in the supports, considering the deformations by sharp. Advances and Applications in Statistics, 20(2): 109–132, February 2011. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/5788.htm.

Ruiz-Olivares:2020:FMP

[RORGTR20] Alejandro Ruiz-Olivares, Martha Elva Ramírez-Guzmán, and Sandy Yaredd Trujano-Ramos. Forecasting Mexican peso–US dollar exchange rate after the 2020 pandemic with GARCH and XGBoost time series models. Advances and Applications in Statistics, 65(1):89–106, November 2020. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
13497.htm.

Rosenthal:2020:MBP

[Ros20] Jeffrey S. Rosenthal. Maximum binomial probabilities and game theory voter models. Advances and Applications in Statistics, 64(1):75–85, September 2020. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/13335.
htm.

Rahman:2015:NUP

[RP15] Mezbahur Rahman and Mohammad Shaha Alam Patwary. A note on unequal probability sampling in bootstrap sampling. Advances and Applications in Statistics, 44(3):191–200, March 2015. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/9122.htm.

Roberts:2003:DTK

[RR03] Gareth O. Roberts and Jeffrey S. Rosenthal. Downweighting tightly knit communities in World Wide Web rankings. Advances and Applications in Statistics, 3(3):199–216, December 2003. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/176.htm.

Rashwan:2014:BME

[RS14] Nasr Rashwan and Hanaa Salem. A Bayesian method to estimate the parameters of regression model. Advances and Applications in Statistics, 43(1):37–52, November 2014. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/8844.htm.

Rojas:2011:STM

[RSM^{+}11] Arnulfo Luévanos Rojas, Alberto Diosdado Salazar, Facundo Cortés Martinez, Ramón Luévanos Rojas, and Armando Cesar Uranga Sifuentes. Study of two models for a continuous beam reinforced concrete for three different sections. Advances and Applications in Statistics, 25(2):131–151, December 2011. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/6479.htm.

Rusiman:2018:PMM

[RSR^{+}18] Mohd Saifullah Rusiman, Suliadi Sufahani, Nur Afifi Izzati Mohd Robi, Abdul Wahab Abdullah, and Nur Amira Azmi. Predictive modelling of marine fish landings in Malaysia. Advances and Applications in Statistics, 53(2):123– 135, August 2018. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/11844.htm.

Rahman:2014:FRB

[RSW14] Mezbahur Rahman, Faysal Ibne Satter, and Han Wu. Fixed regressors in bootstrap regression. Advances and Applications in Statistics, 39(2):95–108, April 2014. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/8318.htm.

Rekkas:2018:TMA

[RW18] M. Rekkas and O. Wong. Transformations, means, and accurate confidence intervals. Advances and Applications in Statistics, 52(2):121–144, February 2018. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/11506.htm.

Rady:2012:MCS

[RYAD12] El-Houssainy Abd Elbar Rady, Ahmed H. Youssef, and Tareq Abdul-Aziz Al-Doub. A Monte Carlo simulation study for (GLS) and (GMM) estimation. Advances and Applications in Statistics, 26(1):1–23, January 2012. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/6617.htm.

Riggs:2008:SIB

[RYS08] Kent Riggs, Dean M. Young, and James D. Stamey. Statistical inference for a bivariate Poisson regression model. Advances and Applications in Statistics, 10(1):55–73, October 2008. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/3500.htm.

Ryu:2004:CCW

[Ryu04] Dae-Hee Ryu. Complete convergence for weighted sums of AANA random variables. Advances and Applications in Statistics, 4(2):117–128, August 2004. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/425.htm.

Rady:2014:EPT

[RZ14] El-Housseiny A. Rady and Dalia Ziedan. Estimation of population total using nonparametric regression models. Advances and Applications in Statistics, 39(1):37–59, March 2014. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/8238.htm.

Ren:2008:DCC

[RZC08] Jiafu Ren, Zongfang Zhou, and Lin Chen. On dynamic combined credit evaluation and error estimation to the ETF’s service agents. Advances and Applications in Statistics, 10(1): 41–53, October 2008. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/3499.htm.

Saidi:2010:BMG

[SA10] G. Saidi and A. Aissani. Bounds on moment generating functions of HNBUE class and its variants. Advances and Applications in Statistics, 15(2):181–194, April 2010. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
4906.htm.

SaadMowafy:2016:SSP

[SA16] Mamdouh Abdel Alim Saad Mowafy and Hanaa Hussein Ali Abou Alaela. A statistical study to predict to development index of information and communication technology (ICT) in the Arab countries. Advances and Applications in Statistics, 49(3):167–194, September 2016. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/10078.htm.

Sadri:2010:SMA

[Sad10] Pejmon Sadri. A simple method for assessing bivariate normality. Advances and Applications in Statistics, 19(1):39–48, November 2010. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/5497.htm.

Saenchan:2017:CIE

[Sae17] Lampang Saenchan. Comparisons of interval estimators for the linear function of variances under non-normality. Advances and Applications in Statistics, 51(5):337–350, November 2017. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/11275.htm.

Shahbaz:2017:BEP

[SAGKS17] Saman Hanif Shahbaz, Saeed Al-Ghamdi, Khushnoor Khan, and Muhammad Qaiser Shahbaz. Bayesian estimation for the parameters of Poisson–Lomax distribution using conditional and unconditional gamma priors. Advances and Applications in Statistics, 51(1):1–15, July 2017. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/11014.htm.

Salmaso:2005:PTS

[Sal05] Luigi Salmaso. Permutation tests in screening two-level factorial experiments. Advances and Applications in Statistics, 5 (1):91–110, April 2005. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/446.htm.

Salmensuu:2018:PPF

[Sal18] Olli Salmensuu. Potato price forecasts in Delhi retail from Bayesian regime switching. Advances and Applications in Statistics, 53(6):731–770, December 2018. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12234.
htm.

Salem:2019:ABE

[Sal19a] Hanaa Abd El-Reheem Salem. Approximate Bayesian estimation for odd Burr III Lomax distribution. Advances and Applications in Statistics, 56(2):173–185, June 2019. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12642.htm.

Salem:2019:EEP

[Sal19b] Hanaa Abd El-Reheem Salem. Exponentiated exponential Pareto distribution: properties and estimation. Advances and Applications in Statistics, 57(1):89–104, July 2019. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12715.htm.

Safdar:2019:PCI

[SAM19] Suboohi Safdar, Ejaz Ahmed, and Arfa Maqsood. Process capability indices for fatigue life model. Advances and Applications in Statistics, 55(2):237–251, April 2019. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
12501.htm.

Stoica:2006:FDR

[SB06] George Stoica and Michael T. Bradley. False discovery rate due to additional outcomes. Advances and Applications in Statistics, 6(1):33–39, April 2006. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/15.htm.

Suzuki:2009:BSP

[SB09] Sumihiro Suzuki and Michael I. Baron. ϵ-Bayes sequential plan. Advances and Applications in Statistics, 11(2):173–183, April 2009. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/3977.htm.

Stoica:2011:PAD

[SB11] George Stoica and Michael T. Bradley. A priori action-decision models: technical results. Advances and Applications in Statistics, 24(1):1–6, September 2011. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/6186.htm.

Sulochana:2020:SSO

[SB20] B. Sulochana and B. Re. Victor Babu. A study on second order slope rotatable designs under tri-diagonal correlated structure of errors using a pair of balanced incomplete block designs. Advances and Applications in Statistics, 65(2):189–208, December 2020. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/13573.htm.

Singh:2003:MEL

[SBZ03] Karan P. Singh, Alfred A. Bartolucci, and Shimin Zheng. Mixed-effects linear models based on elliptically contoured distribution assumptions. Advances and Applications in Statistics, 3(2):135–143, August 2003. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/166.htm.

Saumyamala:2019:HSR

[SC19a] M. G. A. Saumyamala and N. V. Chandrasekara. Hourly solar radiation forecasting using artificial neural network model for Colombo, Sri Lanka. Advances and Applications in Statistics, 56(2):143–151, June 2019. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/12640.htm.

Sezer:2019:TND

[SC19b] Demet Sezer and Nuri Celik. Transmuted normal distribution and an application to finance. Advances and Applications in Statistics, 54(2):301–313, February 2019. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12403.
htm.

Sung:2001:EME

[SCG01] Wen-Pei Sung, Kuen-Suan Chen, and Cheer Germ Go. The effective method of evaluation for the quality control of steel reinforcing bar. Advances and Applications in Statistics, 1(2): 175–193, August 2001. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/44.htm.

Shanubhogue:2019:MCI

[SD19] Ashok Shanubhogue and Rajendra G. Desai. A method of constructing interval estimations for the parameters of a family of distributions under progressive censoring. Advances and Applications in Statistics, 56(1):1–20, May 2019. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
12562.htm.

Sakthivel:2021:TLG

[SD21] K. M. Sakthivel and K. Dhivakar. Topp–Leone generated-Dagum distribution and its properties. Advances and Applications in Statistics, 66(1):13–38, January 2021. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
13615.htm.

Sales:2012:GMP

[SdA12] J. S. Sales and N. G. de Almeida. The generalized Q Mandel’s parameter cannot reveal the nonclassicality of the squeezed coherent states. Advances and Applications in Statistics, 30 (2):135–149, October 2012. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/7152.htm.

Selim:2015:BVB

[Sel15a] Kamal Samy Selim. Bounds on variance based on differences of mass points. Advances and Applications in Statistics, 47 (3):211–224, December 2015. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/9559.htm.

Selim:2015:DMD

[Sel15b] Kamal Samy Selim. Distribution and moments of the double exponential sample range. Advances and Applications in Statistics, 44(3):201–217, March 2015. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/9123.
htm.

Sabry:2020:FJM

[SGM20] Dina M. Sabry, Ahmed M. Gad, and Ramadan H. Mohamed. Fitting joint modeling of longitudinal and time-to-event data using stochastic EM approach. Advances and Applications in Statistics, 64(1):33–62, September 2020. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/13333.
htm.

Sun:2009:ESA

[SH09] Xiaoqian Sun and Zhuoqiong He. An efficient simulation algorithm for Bayesian spatial analysis via the generalized ratioof-uniforms method. Advances and Applications in Statistics, 11(1):47–63, February 2009. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/3971.htm.

Shang:2008:NRB

[Sha08] Junfeng Shang. A note on the relationship between influential observations and model selection. Advances and Applications in Statistics, 10(2):237–246, December 2008. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/3727.htm.

Shang:2010:BBS

[Sha10] Junfeng Shang. The bootstrap-based selection criteria: an optimal choice for model selection in linear regression. Advances and Applications in Statistics, 14(2):173–189, February 2010. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/4606.htm.

Shalabi:2019:NBB

[Sha19] Rania M. Shalabi. Non Bayesian and Bayesian estimation for the bivariate generalized Lindley distribution. Advances and Applications in Statistics, 54(2):327–344, February 2019.

CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/12405.htm.

Shen:2020:OBI

[She20] Yujan Shen. Optimum Bayesian inference for 3-step step-stress tests under Type-II censoring. Advances and Applications in Statistics, 62(1):1–18, May 2020. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/13157.htm.

Sultana:2019:FIW

[SHM19] Nazneen Sultana, Shamima Hossain, and Kazi Md. Farhad Mahmud. Factors influencing women’s opinions regarding violence in Bangladesh. Advances and Applications in Statistics, 59(2):125–147, December 2019. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/12954.htm.

Shoji:2008:SNM

[Sho08] Isao Shoji. A simple numerical method of checking normality in statistical models. Advances and Applications in Statistics, 10(2):259–275, December 2008. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/3729.htm.

Simwa:2006:SMF

[Sim06] R. O. Simwa. On statistical modelling of floods and drought insurance products for developing countries. Advances and Applications in Statistics, 6(2):241–248, August 2006. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/1854.htm.

Singh:2016:EPW

[Sin16a] Trijya Singh. Estimation of parameters of a Weibull-type nonlinear regression model. Advances and Applications in Statistics, 49(6):409–416, December 2016. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/10379.
htm.

Singh:2016:MMA

[Sin16b] Trijya Singh. A method of moments approach to parameter estimation in intrinsically nonlinear regression models. Advances and Applications in Statistics, 49(1):1–20, July 2016. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/10023.htm.

Shanubhogue:2010:MVU

[SJ10] Ashok Shanubhogue and N. R. Jain. Minimum variance unbiased estimation in exponential distribution using progressively type II censored data with binomial removals. Advances and Applications in Statistics, 18(1):27–39, September 2010. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/5368.htm.

Sanubhogue:2012:BPT

[SJ12] Ashok Sanubhogue and A. K. Jiheel. Bayes pre-test estimation of mean of exponential distribution under asymmetric loss function using progressive type II censored sample. Advances and Applications in Statistics, 27(2):109–130, April 2012. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/6746.htm.

Siew:2018:EIT

[SJH18] Lam Weng Siew, Saiful Hafizah Jaaman, and Lam Weng Hoe. Enhanced index tracking model with entropy maximization. Advances and Applications in Statistics, 53(3):243–258, September 2018. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/12014.htm.

Sepanski:2008:FGB

[SK08] J. H. Sepanski and Lingji Kong. A family of generalized beta distributions for income. Advances and Applications in Statistics, 10(1):75–84, October 2008. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/3501.htm.

Shangodoyin:2008:CDT

[SKS08] D. K. Shangodoyin, P. M. Kgosi, and K. Setlhare. A comparison of distribution test statistics for detecting outliers in dataset. Advances and Applications in Statistics, 9(1):109– 125, June 2008. CODEN ???? ISSN 0972-3617. URL http:/

/www.pphmj.com/abstract/3288.htm.

Suzuki:2011:FBL

[SLCB11] Adriano K. Suzuki, Francisco Louzada, Vicente G. Cancho, and Gladys D. C. Barriga. The FGM bivariate lifetime copula model: a Bayesian approach. Advances and Applications in Statistics, 21(1):55–76, March 2011. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/5794.htm.

Shin:2008:NSA

[SLKP08] Key-Il Shin, Sang Eun Lee, and Yonghee Kim-Park. A note on small area estimations for rare event data. Advances and Applications in Statistics, 8(1):73–88, February 2008. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/2839.htm.

Sloboda:2013:TNT

[Slo13] Brian W. Sloboda. Testing the neoclassical theories of labor demand: an application to the time series of the United States. Advances and Applications in Statistics, 35(1):57–73, July 2013. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/7784.htm.

Safadi:2012:WAC

[SM12a] Thelma Safadi and Pedro A. Morettin. A wavelet analysis to compare environmental time series. Advances and Applications in Statistics, 26(2):79–96, February 2012. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
6682.htm.

Shimokawa:2012:AGA

[SM12b] Asanao Shimokawa and Etsuo Miyaoka. Application of genetic algorithm for classification of medical images. Advances and Applications in Statistics, 29(1):1–31, July 2012. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/7063.htm.

Sani:2012:DBS

[SMG12] Ali Akbar Rahimzadeh Sani, Hamid Reza Mostafaei, and Mahdiyeh Ghohyeei. Difference between sample median and population median with odd sample sizes. Advances and Applications in Statistics, 28(2):141–147, June 2012. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
7083.htm.

Sadeghi:2019:DED

[SMHN19] Elham Sadeghi, Reza Moghaddasi, Seyed Safdar Hosseini, and Amir Mohammadi Nejad. Is diversified export a driver for agricultural growth in developing countries? A dynamic model. Advances and Applications in Statistics, 58(2):123–135, October 2019. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/12876.htm.

Sreedhar:2021:SAA

[SMP21] B. R. Sreedhar, K. Muthyalappa, and K. Pushpanjali. Statistical analysis of agricultural performance in different crops by using sampling techniques. Advances and Applications in Statistics, 66(1):1–11, January 2021. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/13614.
htm.

Sedghi:2019:DFP

[SMYH19] Neda Sedghi, Reza Moghaddasi, Saeed Yazdani, and Seyed Safdar Hosseini. On the dependency of food price and production: an Iranian modern time varying copula-based analysis. Advances and Applications in Statistics, 55(2):157–174, April 2019. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/12497.htm.

Somkhuean:2014:UBG

[SNaN14] Rada Somkhuean, Suparat Niwitpong, and Sa aat Niwitpong. On upper bound of the generalized p-value for the mean of lognormal distribution when the coefficient of variation is known. Advances and Applications in Statistics, 41(2):97–107, August 2014. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/8692.htm.

Shangodoyin:2006:TSS

[SOAD06] D. K. Shangodoyin, O. O. Obe, R. Arnab, and S. S. Dlamini. Tool support for systematic test data generation using genetic algorithms. Advances and Applications in Statistics, 6(3):399– 409, December 2006. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/2222.htm.

Shittu:2008:SBU

[SOO08] O. I. Shittu, T. A. Oyeniyi, and T. O. Olatayo. The seasonality of birth in urban southwestern Nigeria. Advances and Applications in Statistics, 10(2):219–235, December 2008. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/3726.htm.

Shanchao:2017:SCR

[SP17] Yang Shanchao and Xiang Ping. Strong consistency of robust nonparametric kernel regression estimation for a-mixing processes. Advances and Applications in Statistics, 51(3):231– 244, September 2017. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/11136.htm.

Sukparungsee:2018:EFA

[SPA18] Saowanit Sukparungsee, Suganya Phantu, and Yupaporn Areepong. Explicit formula of average run length of moving average control chart for Poisson INMA(1) process. Advances and Applications in Statistics, 52(4):235–250, April 2018. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/11666.htm.

Sugito:2020:AQP

[SPM20] Sugito, Alan Prahutama, and Mustafid. Analysis of queue and performance of toll Muktiharjo Semarang, Indonesia as automatic toll service systems. Advances and Applications in Statistics, 60(2):179–188, February 2020. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/13045.
htm.

Shanubhogue:2014:TME

[SR14a] A. Shanubhogue and D. P. Raykundaliya. A test for main effects when observations are randomly right censored. Advances and Applications in Statistics, 39(1):1–23, March 2014. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/8236.htm.

Sotres-Ramos:2014:HRV

[SR14b] David Sotres-Ramos. How randomization validates the significance levels of the Fisher exact test for 2×2 tables. Advances and Applications in Statistics, 42(2):157–164, October 2014. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/8726.htm.

Sujatha:2014:NNP

[SR14c] Ch. Lakshmi Sujatha and M. Vishnu Vardhana Rao. A new non parametric method of estimation of prediction limits and comparison of its performance with semi parametric method. Advances and Applications in Statistics, 42(1):33–46, September 2014. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/8698.htm.

Sotres-Ramos:2017:ADP

[SRCT17] David Sotres-Ramos and Yoni Castillo-Tzec. The asymptotic distribution of the Pearson chi-square statistic for the 2 × 2 table. Advances and Applications in Statistics, 50(6):469–476, June 2017. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/10941.htm.

Sotres-Ramos:2013:ALJ

[SRCTAA13] David Sotres-Ramos, Yoni Castillo-Tzec, and Félix Almendra-Arao. Adjusting the Laster–Johnson–Kotler non-inferiority test for comparing two proportions. Advances and Applications in Statistics, 36(2):151–163, October 2013. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
7889.htm.

Sadri:2017:ILP

[SRMS17] Soroor Sadri, Mohsen Rostamy-Malkhalifeh, and Naghi Shoja. Inverse linear programming in cost efficiency and network. Advances and Applications in Statistics, 51(2):131–149, August 2017. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/11080.htm.

Siray:2010:ILE

[ŞS10] Gülesen Üstündağ Şiray and Sadullah Sakallioğlu. Inadmissibility of the Liu estimator to the ordinary least squares estimator. Advances and Applications in Statistics, 19(1):21–32, November 2010. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/5494.htm.

Sarma:2016:EFM

[SS16] B. Bhargavarama Sarma and B. Shobha. Estimation in functional measurement error models with compound symmetric covariance structured equation errors. Advances and Applications in Statistics, 49(2):137–166, August 2016. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
10077.htm.

Sumalinab:2018:BEM

[SS18] Bryan P. Sumalinab and Arnulfo P. Supe. Bayesian estimation of GARCH(p, q) model. Advances and Applications in Statistics, 53(4):345–361, October 2018. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12109.
htm.

Shahbaz:2019:CDG

[SS19] Saman Hanif Shahbaz and Muhammad Qaiser Shahbaz. On concomitants of dual generalized order statistics for a bivariate inverse Weibull distribution. Advances and Applications in Statistics, 55(1):35–46, March 2019. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/12442.htm.

Sides:2021:UMT

[SS21] Ryan Sides and James Stamey. Underreporting in mixed treatment comparisons meta-analysis for Poisson data. Advances and Applications in Statistics, 67(2):225–236, April 2021. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13760.htm.

Sukparungsee:2016:EEA

[SSA16] Saowanit Sukparungsee, Sophana Somran, and Yupaporn Areepong. Exact expression of average run length for twosided exponential cusum procedure. Advances and Applications in Statistics, 48(6):429–444, June 2016. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
9999.htm.

Sharghi:2020:UTB

[SSBA20] Sima Sharghi, Kevin E. Stoll, Kimberlyn K. Brooks, and Andrew Alt. Using tree-based methods to predict student retention with an emphasis on the education, wealth, and density of the hometown neighborhood. Advances and Applications in Statistics, 60(2):105–135, February 2020. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/13041.
htm.

Shen:2007:MGJ

[SSL07] Yu-Jan Shen, Kuan-Fu Shen, and Ming-Chih Lee. Mixed GARCH-jump models with generalized error distribution for assets returns. Advances and Applications in Statistics, 7(3): 403–416, December 2007. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/2720.htm.

Sanchez:2016:PSC

[SSRG16] Julio César Martínez Sánchez, David Sotres-Ramos, and Martha Elva Ramírez Guzmán. Portfolio selection with conditional covariance matrix and nonlinear programming. Advances and Applications in Statistics, 49(5):343–355, November 2016. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/10272.htm.

Sinaj:2011:APC

[ST11] Valentina Sinaj and Arjan Tushaj. Analysis problems and creating forecast models for CPI in Albania. Advances and Applications in Statistics, 24(2):107–128, October 2011. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/6279.htm.

Calis:2011:EAM

[şTE^{+}11] Nazif Çalış, Ayça Hatice Türkan, Murat Erişoğlu, Ülkü Erişoğlu, and Hamza Erol. EM algorithm for the mixtures of the Marshall–Olkin bivariate Weibull distributions. Advances and Applications in Statistics, 24(1):35–48, September 2011. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/6188.htm.

Stoica:2006:BKL

[Sto06a] George Stoica. A Baum–Katz law under regularly varying moment conditions. Advances and Applications in Statistics, 6(3):411–417, December 2006. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/2223.htm.

Stoica:2006:NPE

[Sto06b] George Stoica. Non-parametric estimation of Barlow’s model. Advances and Applications in Statistics, 6(2):235–240, August 2006. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/1851.htm.

Stoica:2008:CCS

[Sto08] George Stoica. A counterexample concerning a series of T. L. Lai. Advances and Applications in Statistics, 9(1):141–144, June 2008. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/3290.htm.

Stoica:2011:SNW

[Sto11] George Stoica. A self-normalized weak law of large numbers for exchangeable sequences. Advances and Applications in Statistics, 25(2):103–108, December 2011. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/6475.htm.

Sontimoon:2010:ERT

[STPS10] Nuengruithai Sontimoon, Montip Tiensuwan, Simona Panunzi, and Wutthi Sumetchotimaytha. Estimating recurrent time using nonlinear mixed effects model in patients with colorectal cancer. Advances and Applications in Statistics, 17 (1):41–60, July 2010. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/5165.htm.

Stummer:2002:DRR

[Stu02] Wolfgang Stummer. Decision risk reductions for stock indices. Advances and Applications in Statistics, 2(1):79–99, April 2002. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/77.htm.

Sun:2011:NCD

[Sun11] Li Rong Sun. A new chart for defecting asymmetric loss. Advances and Applications in Statistics, 20(1):43–53, January 2011. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/5749.htm.

Salicru:2004:CIC

[SV04] Miquel Salicrú and Antonio Vaamonde. Confidence intervals for capability indices. Advances and Applications in Statistics, 4(2):129–137, August 2004. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/426.htm.

Santos:2012:AFT

[SV12] Patrícia Borchardt Santos and Dione Maria Valenca. Accelerated failure time models with random effects to estimate operation time of oil wells. Advances and Applications in Statistics, 28(1):23–44, May 2012. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/6856.htm.

Salicru:2002:ABM

[SVSO02] M. Salicrú, S. Vives, J. Sanchez, and F. Oliva. Asymptotic and bootstrap methodologies on the estimation of the sonorous sensation. Advances and Applications in Statistics, 2(3):311– 322, December 2002. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/135.htm.

Sirikantisophon:2019:LUC

[SWCT19] K. Sirikantisophon, W. Wanishsakpong, P. Chuangchang, and O. Thinnukool. Land use change over time in Chiang Mai Province, Thailand. Advances and Applications in Statistics, 54(2):255–272, February 2019. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/12400.htm.

Swift:2008:BDP

[Swi08] Randall J. Swift. A birth-death process with deterministic mean population. Advances and Applications in Statistics, 9 (1):101–108, June 2008. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/3287.htm.

Satoh:2004:CMC

[SYO04] Kenichi Satoh, Hirokazu Yanagihara, and Megu Ohtaki. Clustering method by connected neighborhoods and its application. Advances and Applications in Statistics, 4(2):223–231, August 2004. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/433.htm.

Shi:2008:ASL

[SZZ08] Yang Shi, Zhiqiang Zhang, and Riquan Zhang. On almost sure limit theorem and strong law of large numbers for squares sequence from ARCH(1). Advances and Applications in Statistics, 10(1):149–154, October 2008. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/3505.htm.

Tanusit:2010:CSP

[Tan10] Manlika Tanusit. A comparative study of parameter inference methods for Poisson distribution: small sample sizes. Advances and Applications in Statistics, 14(1):49–55, January 2010. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/4519.htm.

Tsou:2005:PRT

[TC05] Tsung-Shan Tsou and Li-Chu Chien. Parametric robust tests for multiple regression parameters under generalized linear models. Advances and Applications in Statistics, 5(1):51–86, April 2005. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/444.htm.

Tarepe:2020:SAP

[TCD20] Dennis A. Tarepe, Kennet G. Cuarteros, and Palawan M. Diangca. A simulation assessment of the properties of the estimators of a partially linear model. Advances and Applications in Statistics, 65(2):209–226, December 2020. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/13574.
htm.

Teimouri:2011:BWD

[TG11] Mahdi Teimouri and Arjun K. Gupta. On a bivariate Weibull distribution. Advances and Applications in Statistics, 22(2): 77–106, June 2011. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/6041.htm.

Tomizawa:2002:PDT

[TH02] Sadao Tomizawa and Tomoyuki Hashimoto. Power-divergence type measure of departure from uniform association for crossclassifications having ordered categories. Advances and Applications in Statistics, 2(3):233–248, December 2002. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/125.htm.

Tsou:2014:RRM

[TH14] Tsung-Shan Tsou and Wei-Cheng Hsiao. Robust regression modeling via parametric likelihoods: an alternative to using the skew t distribution and several rank-based methods. Advances and Applications in Statistics, 40(2):109–132, June 2014. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/8571.htm.

Thiagarajah:2008:FEP

[Thi08] K. Thiagarajah. Fuzzy estimation of the parameters in location-scale models. Advances and Applications in Statistics, 10(2):247–258, December 2008. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/3728.htm.

Thiagarajah:2009:FVM

[Thi09] K. Thiagarajah. Forecasts of volatility models: fuzzy estimation approach. Advances and Applications in Statistics, 13(1): 83–94, October 2009. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/4205.htm.

Tinungki:2018:EBA

[Tin18] Georgina Maria Tinungki. The election of the best autoregressive integrated moving average model to forecasting rice production in Indonesia. Advances and Applications in Statistics, 52(4):251–265, April 2018. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/11667.htm.

Tanokura:2004:PCA

[TK04] Yoko Tanokura and Genshiro Kitagawa. Power contribution analysis for multivariate time series with correlated noise sources. Advances and Applications in Statistics, 4(1):65–95, April 2004. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/422.htm.

Thuong:2019:ASM

[TKD19] N. Q. Thuong, V. A. Kolobkova, and P. T. Dong. Application of the statistical method of the stable design solutions synthesis for UAV’s development. Advances and Applications in Statistics, 59(1):43–54, November 2019. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12919.
htm.

Tomizawa:2007:IAU

[TMO07] Sadao Tomizawa, Nobuko Miyamoto, and Noriaki Ohba. Improved approximate unbiased estimators of measures of asymmetry for square contingency tables. Advances and Applications in Statistics, 7(1):47–63, April 2007. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
2426.htm.

Tomizawa:2008:DIM

[TMS08] Sadao Tomizawa, Nobuko Miyamoto, and Masaya Sakurai. Decomposition of independence model and separability of its test statistic for two-way contingency tables with ordered categories. Advances and Applications in Statistics, 8(2):209–218, April 2008. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/3085.htm.

Tripodis:2014:EMV

[TN14] Yorghos Tripodis and Nagaraj K. Neerchal. Estimation of missing values in linear models. Advances and Applications in Statistics, 39(2):109–123, April 2014. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/8319.
htm.

Thangjai:2018:CIC

[TN18] Warisa Thangjai and Suparat Niwitpong. Confidence intervals for common variance of several one-parameter exponential populations. Advances and Applications in Statistics, 53 (3):285–311, September 2018. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/12016.htm.

Thangjai:2017:CIC

[TNN17a] Warisa Thangjai, Sa-Aat Niwitpong, and Suparat Niwitpong. Confidence intervals for the common mean of several oneparameter exponential populations. Advances and Applications in Statistics, 51(4):245–260, October 2017. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
11233.htm.

Thangjai:2017:LSC

[TNN17b] Warisa Thangjai, Suparat Niwitpong, and Sa-Aat Niwitpong. On large sample confidence intervals for the common inverse mean of several normal populations. Advances and Applications in Statistics, 51(1):59–84, July 2017. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/11017.
htm.

Thangjai:2018:CIV

[TNN18] Warisa Thangjai, Sa-Aat Niwitpong, and Suparat Niwitpong. Confidence intervals for variance and difference between variances of one-parameter exponential distributions. Advances and Applications in Statistics, 53(3):259–283, September 2018. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/12015.htm.

Towhidi:2010:OME

[TNR10] Mina Towhidi and Mohammad-Reza Namazi-Rad. An optimal method of estimation in rotation sampling. Advances and Applications in Statistics, 15(2):115–136, April 2010. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/4902.htm.

Terbeche:2005:TSD

[TO05] Mekki Terbeche and Broderick O. Oluyede. Two stage design for estimation of mean difference in the exponential family. Advances and Applications in Statistics, 5(3):325–339, December 2005. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/664.htm.

Tolotti:2012:LFA

[Tol12] Marco Tolotti. A linear filtering approach for incomplete accounting information models. Advances and Applications in Statistics, 29(2):135–151, August 2012. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/7131.
htm.

Toropov:2020:PAF

[Tor20] V. D. Toropov. A possible approach to filling gaps in the source data. Advances and Applications in Statistics, 61(2): 161–167, April 2020. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/13142.htm.

Toropov:2021:AMM

[Tor21] V. D. Toropov. Application of the method of multiple estimation of linear regression parameters for modeling the level of economic transparency. Advances and Applications in Statistics, 66(1):97–102, January 2021. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/13619.htm.

Tzavelas:2002:OPD

[TP02] George Tzavelas and Demosthenes B. Panagiotakos. Optimal prediction for diagnostic tests using logistic regression analysis: an algorithm and large sample properties. Advances and Applications in Statistics, 2(3):301–309, December 2002. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/134.htm;   http://www.pphmj.com/
abstract/139.htm.

Tzavelas:2005:DEA

[TPP05] George Tzavelas, Athanasios G. Paliatsos, and Nikolaos Prezerakos. Distribution of the extreme annual precipitation over urban areas. Advances and Applications in Statistics, 5(1): 111–123, April 2005. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/447.htm.

Tripsiannis:2007:MGR

[TPP07a] G. A. Tripsiannis, A. N. Philippou, and A. A. Papathanasiou. Multivariate generalized run-related distributions. Advances and Applications in Statistics, 7(1):141–156, April 2007. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/2432.htm.

Tripsiannis:2007:MEG

[TPP07b] Gregory A. Tripsiannis, Afroditi A. Papathanasiou, and Andreas N. Philippou. Multivariate extended generalized distributions of order k. Advances and Applications in Statistics, 7 (2):211–227, August 2007. CODEN ???? ISSN 0972-3617.

Taj:2018:RAU

[TRA^{+}18] S. Z. Taj, S. M. Rizwan, B. M. Alkali, D. K. Harrison, and G. Taneja. Reliability analysis of a 3-unit subsystem of a cable plant. Advances and Applications in Statistics, 52(6): 413–429, June 2018. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/11722.htm.

Turner:2012:ACC

[TRCH12] Matthew A. Turner, Jeffrey S. Rosenthal, Jian Chen, and Chunyan Hao. Adaptation to climate change: evidence from 18th and 19th century Iceland. Advances and Applications in Statistics, 28(2):77–108, June 2012. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/7079.htm.

Tiensuwan:2010:ABJ

[TS10] Montip Tiensuwan and Sukuman Sarikavanich. Applications of Box–Jenkins models to rainfall data of Chiang Mai province in Thailand. Advances and Applications in Statistics, 15(1): 27–48, March 2010. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/4858.htm.

Tsao:2002:UVL

[Tsa02] Min Tsao. Uniform validity of Lugannani–Rice expansion for the tail probability of a sample mean. Advances and Applications in Statistics, 2(1):31–37, April 2002. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/64.htm.

Tomizawa:2001:PDT

[TSU01] Sadao Tomizawa, Masahiko Sagae, and Makiko Uehara. Power-divergence type measure of departure from an entropy model. Advances and Applications in Statistics, 1(2):149–163, August 2001. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/42.htm.

Thomas:2008:MVG

[TTM08] Seemon Thomas, Alex Thannippara, and A. M. Mathai. On a matrix-variate generalized Type-2 Dirichlet density. Advances and Applications in Statistics, 8(1):37–56, February 2008. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/2837.htm.

Tahata:2008:IAU

[TTT08] Kouji Tahata, Ryota Tomisato, and Sadao Tomizawa. An improved approximate unbiased estimator of log-odds ratio for 2 × 2 contingency tables. Advances and Applications in Statistics, 9(1):1–12, June 2008. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/3282.htm.

Thandrayen:2013:LCA

[TW13] Joanne Thandrayen and Yan Wang. A latent class approach with covariates and local dependence in capture-recapture models. Advances and Applications in Statistics, 34(1):65–84, May 2013. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/7684.htm.

Tan:2016:REE

[TW16] Ninghui Tan and Jibo Wu. On a restricted estimator in an errors-in-variable model. Advances and Applications in Statistics, 49(4):257–265, October 2016. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/10224.htm.

Thaithanan:2020:CFM

[TW20] Jeeraporn Thaithanan and Chantha Wongoutong. A combined forecasting model for predicting the number of road traffic accident deaths in Thailand. Advances and Applications in Statistics, 64(2):143–163, October 2020. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/13436.
htm.

Tzavelas:2008:SNR

[Tza08] George Tzavelas. Some new results concerning the efficiency of the method of moments estimators for the location parameter case. Advances and Applications in Statistics, 10(1):1–14, October 2008. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/3496.htm.

Ugwuanyim:2021:RAC

[UABW21] Geoffrey U. Ugwuanyim, Harrison O. Amuji, Desmond C. Bartholomew, and Hyginus Wisdom. The reliability of art in the control of Human Immune Virus (HIV). Advances and Applications in Statistics, 67(2):117–132, April 2021. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13755.htm.

ulanSabir:2020:IIR

[uaSYHNQ20] Qurat ul an Sabir, Uzma Yasmeen, Muhammad Hanif, and Tri Nguyen-Quang. Identification of influence of the risk factors to algal bloom in fresh water reservoirs using multinomial logistics regression. Advances and Applications in Statistics, 63(2):219–233, August 2020. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/13296.htm.

Ueda:2013:NEP

[UH13] Kengo Ueda and Hiroto Hyakutake. Note on estimating parameters in heterogeneous random effects models for two sample problem. Advances and Applications in Statistics, 37(2): 187–193, December 2013. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/7972.htm.

Ula:2015:CEX

[Ula15] Taylan A. Ula. Corrections for estimated XBAR control charts. Advances and Applications in Statistics, 47(1):51–63, October 2015. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/9472.htm.

Um:2005:NGC

[Um05] Yonghwan Um. A new generalization of Cohen’s kappa to multivariate interval data and several observers. Advances and Applications in Statistics, 5(2):159–169, August 2005. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/450.htm.

Um:2008:NMM

[Um08] Yonghwan Um. A new multivariate measure of association for nominal independent variable. Advances and Applications in Statistics, 10(1):15–24, October 2008. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/3497.
htm.

Udayabaskaran:2013:POH

[UMV13] S. Udayabaskaran, R. Murthy, and J. Viswanath. The problem of optimal harvesting in a stochastically branching population. Advances and Applications in Statistics, 33(1):55–62, March 2013. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/7457.htm.

Van:2020:IFF

[Van20] Nguyen Hong Van. Influencing factors on the formation and development of the leather footwear industry association cluster in the northern key economic region in Vietnam. Advances and Applications in Statistics, 65(1):107–131, November 2020. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/13498.htm.

Venema:2015:GEN

[Ven15] Rieken S. Venema. Geometric ergodicity of the normalized perceptron algorithm for the classification of non-separable Gaussian input vectors. Advances and Applications in Statistics, 46(1):37–55, July 2015. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/9331.htm.

Victorbabu:2008:MSO

[Vic08] B. Re. Victorbabu. On modified second order slope rotatable designs using incomplete block designs with unequal block sizes. Advances and Applications in Statistics, 8(1):131–151, February 2008. CODEN ???? ISSN 0972-3617. URL http:/

/www.pphmj.com/abstract/2843.htm.

Visek:2014:DRI

[Víš14] Jan Ámos Víšek. Diagnostics of robust identification of model. Advances and Applications in Statistics, 43(2):119–161, December 2014. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/8861.htm.

Visek:2015:RLW

[Víš15] Jan Ámos Víšek. Representation of the least weighted squares. Advances and Applications in Statistics, 47(2):91–144, November 2015. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/9494.htm.

Verma:2014:ICV

[VPG^{+}14] U. Verma, H. P. Piepho, M. Goyal, A. Goyal, and P. Verma. Impact of climatic variables on sugarcane yield prediction in Haryana (India). Advances and Applications in Statistics, 39 (1):25–35, March 2014. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/8237.htm.

Verma:2012:DAS

[VPKG12] U. Verma, H. P. Piepho, M. H. Kalubarme, and M. Goyal. Development of agrometeorological and spectral wheat yield models using various statistical procedures in Haryana State (India). Advances and Applications in Statistics, 26(2):113– 136, February 2012. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/6684.htm.

Verma:2012:MLM

[VPO^{+}12] U. Verma, H. P. Piepho, J. O. Ogutu, M. H. Kalubarme, and M. Goyal. Multi-level mixed modelling for weather–crop-yield relationships on agro-climatic zone basis in Haryana (Northern India). Advances and Applications in Statistics, 28(1):1–22, May 2012. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/6855.htm.

Vrbik:2010:MMP

[Vrb10] Jan Vrbik. Moments of AR(2)-model parameter estimators. Advances and Applications in Statistics, 18(1):57–72, September 2010. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/5370.htm.

Vrbik:2011:EAC

[Vrb11] Jan Vrbik. Edgeworth approximation of confidence regions. Advances and Applications in Statistics, 24(2):129–138, October 2011. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/6280.htm.

Vrbik:2013:ACR

[Vrb13a] Jan Vrbik. Accurate confidence regions based on MLEs. Advances and Applications in Statistics, 32(1):33–56, January 2013. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/7389.htm. See errata and addendum

[Vrb13b].

Vrbik:2013:EAA

[Vrb13b] Jan Vrbik. Errata and addendum to “Accurate confidence regions based on MLEs” [MR3060565]. Advances and Applications in Statistics, 35(2):161–163, August 2013. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/

7847.htm. See [Vrb13a].

Vrbik:2014:AES

[Vrb14a] Jan Vrbik. Asymptotic expansion of sampling distribution of an MLE to 1/n accuracy. Advances and Applications in Statistics, 38(1):37–50, January 2014. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/8127.htm.

Vrbik:2014:IAT

[Vrb14b] Jan Vrbik. Improving accuracy of χ^{2} test of independence. Advances and Applications in Statistics, 39(2):81–94, April 2014. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/8317.htm.

Vrbik:2015:ACR

[Vrb15] Jan Vrbik. Accurate confidence regions for Markov-model parameters. Advances and Applications in Statistics, 45(1):59– 74, April 2015. CODEN ???? ISSN 0972-3617. URL http:/

/www.pphmj.com/abstract/9142.htm.

Vrbik:2017:ADS

[Vrb17a] Jan Vrbik. Approximate distribution of sample quantiles. Advances and Applications in Statistics, 50(5):397–410, May 2017. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/10826.htm.

Vrbik:2017:GFS

[Vrb17b] Jan Vrbik. General formulas for serial correlation, variance and likelihood function relating to AR(k) models. Advances and Applications in Statistics, 50(5):411–421, May 2017. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/10827.htm.

Victorbabu:2006:MSO

[VVV06] B. Re. Victorbabu, V. Vasundharadevi, and B. Viswanadham. Modified second order response surface designs, rotatable designs using pairwise balanced designs. Advances and Applications in Statistics, 6(3):323–334, December 2006. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/2217.htm.

Wang:2007:SOB

[Wan07] Xikui Wang. Stochastic order and bandit processes with survival time responses. Advances and Applications in Statistics, 7(2):245–254, August 2007. CODEN ???? ISSN 0972-3617.

Wang:2008:BFD

[Wan08] Jiantian Wang. On beta-family distributions. Advances and Applications in Statistics, 9(2):233–246, August 2008. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/3448.htm.

Wang:2010:CDT

[Wan10] Jiantian Wang. Characterization of distributions through hazard functions and reversed hazard functions. Advances and Applications in Statistics, 19(2):149–153, December 2010. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/5680.htm.

Wang:2013:NRB

[Wan13] Jiantian Wang. A note on the relationships between Spearman’s ρ and Kendall’s τ under bivariate homogeneous shock model. Advances and Applications in Statistics, 34(1):1–9, May 2013. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/7680.htm.

Wang:2014:LDA

[Wan14] Zailong Wang. Longitudinal data analysis strategies — an application of mixed model to post-transplant serum creatinine data. Advances and Applications in Statistics, 41(2):137–165, August 2014. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/8695.htm.

Wongkhao:2014:CII

[WaNN14] Arunee Wongkhao, Sa aat Niwitpong, and Suparat Niwitpong. Confidence intervals for the inverse of mean in a normal distribution with a known coefficient of variation. Advances and Applications in Statistics, 42(1):1–14, September 2014. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/8696.htm.

Wu:2014:RPM

[WCW14] Ka Ho Wu, Zhao-Guo Chen, and Wensheng Wang. A robust and practical method of unit root tests. Advances and Applications in Statistics, 41(1):21–67, July 2014. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
8627.htm.

Wahyuningsih:2010:TVA

[WDGP10] Sri Wahyuningsih, Sutawanir Darwis, Agus Yodi Gunawan, and Asep Kurnia Permadi. Time varying autoregressions for hyperbolic decline curve. Advances and Applications in Statistics, 15(1):61–70, March 2010. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/4860.htm.

Wigena:2014:ETR

[WDM14] A. H. Wigena, A. Djuraidah, and I. W. Mangku. Ensemble two return levels of generalized Pareto and modified Champernowne distributions using linear regression. Advances and Applications in Statistics, 40(2):157–167, June 2014. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/8573.htm.

Wei:2015:ELR

[Wei15] Chuanhua Wei. Empirical likelihood ratio test for serial correlation in additive models. Advances and Applications in Statistics, 45(3):153–165, June 2015. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/9257.htm.

Wang:2017:FCP

[WH17] Yong Wang and Juan Hu. Fast computation of predictive scores under cross-validation for multivariate spatial models. Advances and Applications in Statistics, 50(3):173–189, March 2017. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/10626.htm.

Wijayawardhana:2020:PAM

[WH20] H. N. A. M. Wijayawardhana and A. P. Hewaarachchi. Periodic autoregressive model for temperature data in Nuwara Eliya, Sri Lanka. Advances and Applications in Statistics, 63 (1):85–96, July 2020. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/13236.htm.

Woolford:2013:NAB

[WJ13] S. Woolford and N. Josephy. A note on the asymptotic behavior of a nonlinear system of stochastic difference equations. Advances and Applications in Statistics, 33(1):23–35, March 2013. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/7455.htm.

Woo:2018:CEM

[WK18] Song Jea Woo and Kee-Hoon Kang. On comparison of estimation methods in quantile regression. Advances and Applications in Statistics, 52(3):203–213, March 2018. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
11564.htm.

Wang:2008:SCE

[WM08] Wen-Xin Wang and Yan-Peng Ma. Stochastic comparisons of excess lifetimes of renewal processes based upon generalized stochastic orderings. Advances and Applications in Statistics, 9(2):283–292, August 2008. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/3451.htm.

Wei:2013:BAU

[WNB13] Rong Wei, Balgobin Nandram, and Dilli Bhatta. A Bayesian analysis of US mortality curves for race-sex domains by small area. Advances and Applications in Statistics, 34(2):137–165, June 2013. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/7664.htm.

Wanishsakpong:2020:CET

[WNT20] Wandee Wanishsakpong, Khairil Anwar Notodiputro, and Jeeraporn Thaithanan. On comparing the efficiency of two classification methods for precipitation occurrence in Bangkok. Advances and Applications in Statistics, 60(1):79–91, January 2020. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/12998.htm.

Wongoutong:2020:ICM

[Won20] Chantha Wongoutong. Imputation for consecutive missing values in non-stationary time series data. Advances and Applications in Statistics, 64(1):87–102, September 2020. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13336.htm.

Walter:2021:NCG

[WPJC21a] Onchere Walter, Weke Patrick, Otieno JAM, and Ogutu Carolyne. Non-central gamma frailty with application to life term assurance data. Advances and Applications in Statistics, 67 (2):237–253, April 2021. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/13761.htm.

Walter:2021:SFM

[WPJC21b] Onchere Walter, Weke Patrick, Otieno JAM, and Ogutu Carolyne. Shared frailty model with application in joint-life annuity insurance. Advances and Applications in Statistics, 68 (1):23–42, May 2021. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/13786.htm.

Wu:2016:SGM

[WRP16] Han Wu, Mezbahur Rahman, and Mohammad Shaha Alam Patwary. A study of a generalized modified Weibull model: an exponentiated Weibull case. Advances and Applications in Statistics, 49(5):357–367, November 2016. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/10273.
htm.

Wang:2015:IAR

[WS15a] Shengxi Wang and Junfeng Shang. Independence among random variables between semi-martingales. Advances and Applications in Statistics, 44(3):219–239, March 2015. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
9124.htm.

Wu:2015:AUL

[WS15b] Jibo Wu and Abdul Shakoor. An almost unbiased Liu estimator in linear regression model with measurement error. Advances and Applications in Statistics, 44(1):77–85, January 2015. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/9018.htm.

Wu:2001:NMA

[Wu01] Dane W. Wu. A new measure of attributable risk and its stratification. Advances and Applications in Statistics, 1(1): 49–58, April 2001. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/34.htm.

Wu:2002:SSK

[Wu02] Dane W. Wu. Some study of kurtosis: a new measure of peakedness for distributions. Advances and Applications in Statistics, 2(2):131–137, August 2002. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
104.htm.

Wu:2014:SMW

[Wu14] Shiang-Jen Wu. Stochastic modeling of wave equation with uncertainties in initial and boundary conditions. Advances and Applications in Statistics, 42(2):119–156, October 2014. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/8725.htm.

Wu:2016:NDB

[Wu16] Jibo Wu. A note on the difference-based almost unbiased Liu estimator in partial linear model. Advances and Applications in Statistics, 48(5):363–368, May 2016. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
9914.htm.

Wu:2017:RTE

[Wu17] Jibo Wu. Ridge-type estimation in semiparametric EV model with linear measurement error. Advances and Applications in Statistics, 51(6):389–395, December 2017. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/11385.
htm.

Wang:2014:BIP

[WY14] Xiaoyin Wang and Dongqing Yao. Bayesian inference of predictor’s relative importance in multiple regression. Advances and Applications in Statistics, 41(1):69–96, July 2014. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/8628.htm.

Wang:2010:MRA

[WYZ10] Shaojun Wang, Xiaoping Yang, and Peibiao Zhao. Management risk analysis of real estate companies. Advances and Applications in Statistics, 15(2):171–180, April 2010. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/4905.htm.

Wang:2011:RSE

[WZG11] Nan Wang, Chun-Xiao Zhang, and You Gao. Research on symmetric engines parameters difference in aero-engine condition monitoring. Advances and Applications in Statistics, 20 (1):55–65, January 2011. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/5750.htm.

Xie:2010:SPE

[XB10] Xian-Jin Xie and Aihua Bian. A SAS package for evaluating logistic and proportional odds model fit. Advances and Applications in Statistics, 14(1):57–64, January 2010. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/4520.htm.

Xu:2016:AEV

[XDL16] Yingmei Xu, Diawara Daouda, and Kané Ladji. An application of extreme value theory in automobile insurance: a new approach to the convex combination of two variables thresholds minimizing the variance of this combination. Advances and Applications in Statistics, 48(2):91–108, February 2016. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/9779.htm.

Xia:2014:PFS

[XW14] Jing Xia and Michelle Yongmei Wang. Particle filtering with sequential parameter learning for nonlinear bold fMRI signals. Advances and Applications in Statistics, 40(1):61–74, May 2014. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/8461.htm.

Xiao:2019:SPM

[XW19] Yongshun Xiao and Tony Wragg. A stochastic process model for the survival and splintering of terrorist organizations. Advances and Applications in Statistics, 55(1):105–155, March 2019. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/12447.htm.

Youssef:2016:SSS

[YAA16] Ahmed H. Youssef, Ali H. Almarshadi, and Abdulqader H. Almusabi. A simulation study for sure estimators with an application in air navigation in Saudi Arabia. Advances and Applications in Statistics, 48(6):411–428, June 2016. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/9998.htm.

Yang:2008:EJS

[Yan08] Y. S. Yang. Evaluation of Johnson’s system for transforming non-normal data. Advances and Applications in Statistics, 8 (1):57–71, February 2008. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/2838.htm.

Yang:2012:DBC

[Yan12] Xin Yang. The data-based choice of bandwidth for kernel quantile estimator of VaR. Advances and Applications in Statistics, 31(1):43–54, November 2012. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/7186.
htm.

Yoo:2013:ESB

[YB13] Changwon Yoo and Erik M. Brilz. Efficient and scalable Bayesian statistical method for identifying causal relationships from intervention studies. Advances and Applications in Statistics, 37(2):95–122, December 2013. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
7968.htm.

Yehia:2016:SCO

[Yeh16] Enas Gawdat Yehia. The shqc criterion for order selection in autoregressive models. Advances and Applications in Statistics, 48(3):169–183, March 2016. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/9843.htm.

Yelamanchili:2020:MSM

[Yel20] Rama Krishna Yelamanchili. Modeling stock market daily returns volatility using symmetric and asymmetric GARCH models with three different distributions. Advances and Applications in Statistics, 63(2):119–140, August 2020. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/13290.htm.

Yuen:2009:WBD

[YH09] Wai Kong Yuen and Mei Ling Huang. A weighted bivariate density estimation method. Advances and Applications in Statistics, 13(2):181–191, December 2009. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
4447.htm.

Yigiter:2010:TCP

[Yİ10] Ayten Yiğiter and Ceyhan İnal. Test for a change point in Poisson process with applications to the British coal mining disasters and the air traffic flow data set. Advances and Applications in Statistics, 14(2):145–155, February 2010. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/4604.htm.

Yoo:2003:BAO

[YKK03] Hee-Kyung Yoo, Hea-Jung Kim, and Dae-Hwang Kim. A Bayesian approach for ordering products of independent normal means. Advances and Applications in Statistics, 3(3):267– 283, December 2003. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/310.htm.

Yamamoto:2014:MPS

[YKT14] Kouji Yamamoto, Kanau Kawasaki, and Sadao Tomizawa. Marginal point-symmetry measure for multi-way contingency tables. Advances and Applications in Statistics, 40(1):75–91, May 2014. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/8462.htm.

Youssef:2012:DFD

[YM12] Ahmed H. Youssef and Nada Ali-Mohammed Muqaibal. Determinants of foreign direct investment inflows of Gulf Cooperation Council: panel gravity model approach. Advances and Applications in Statistics, 28(1):55–75, May 2012. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/6858.htm.

Yigit:2011:TER

[YMM11] Soner Yiǧit, Mehmet Mendeş, and Hamıt Mırtagıoğlu. Type I error rates and test power for some outlier detecting tests. Advances and Applications in Statistics, 21(2):159–167, April 2011. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/5822.htm.

Yamamoto:2010:HGA

[YMT10] Kouji Yamamoto, Nobuko Miyamoto, and Sadao Tomizawa. Harmonic, geometric and arithmetic means type uncertainty measures for two-way contingency tables with nominal categories. Advances and Applications in Statistics, 17(2):143–159, August 2010. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/5234.htm.

Yamamoto:2013:GMM

[YMT13] Kouji Yamamoto, Kazuho Masumura, and Sadao Tomizawa. Generalized measures on marginal symmetry for nominal square contingency tables. Advances and Applications in Statistics, 32(2):139–149, February 2013. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/7443.
htm.

Yosef:2015:SCN

[Yos15] Arthur Yosef. Some classical-new results on the set-indexed Brownian motion. Advances and Applications in Statistics, 44 (1):57–76, January 2015. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/9017.htm.

Youssef:2005:FSP

[You05] Ahmed H. Youssef. Finite sample properties of the random coefficients regression model. Advances and Applications in Statistics, 5(2):247–256, August 2005. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
560.htm.

Yang:2010:HOR

[YQCSG10] Wei Yang, Fares Qeadan, Daniel Cook, and Julie Smith-Gagen. Heterogeneity of odds ratios over time in crosssectional trend studies: an overweight-obesity example from a nationwide population-based study. Advances and Applications in Statistics, 17(2):175–190, August 2010. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
5236.htm.

Yosef:2017:MRI

[YS17] Arthur Yosef and Eli Shnaider. On measuring the relative importance of explanatory variables in a soft regression method. Advances and Applications in Statistics, 50(3):201–228, March 2017. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/10628.htm.

Yang:2006:ODP

[YSK06] Yang Yang, Isao Shoji, and Sumei Kanehiro. Optimal dividend policy of life insurance company with policyholders’ dynamically inconsistent time preferences. Advances and Applications in Statistics, 6(1):1–15, April 2006. CODEN ???? ISSN 0972-

3617. URL http://www.pphmj.com/abstract/9.htm.

Yang:2010:LIC

[YT10] Chi-Chuan Yang and Tsung-Shan Tsou. Likelihood inferences for correlated nominal data without knowing the intra-cluster correlations. Advances and Applications in Statistics, 17(1):

29–40, July 2010. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/5164.htm.

Yin:2013:IFD

[YT13] Jian Yin and Lixin Tian. The impact of foreign direct investment on economic growth in China. Advances and Applications in Statistics, 36(1):75–81, September 2013. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/7906.htm.

Yang:2019:GMI

[YXY19] Shanchao Yang, Weiwei Xu, and Xin Yang. G-M integrated type instantaneous volatility estimation. Advances and Applications in Statistics, 55(2):175–192, April 2019. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
12498.htm.

Ying:2020:KWV

[YY20] Guobing Ying and Shanchao Yang. Kernel weighted volatility estimation for stochastic diffusion model with jump. Advances and Applications in Statistics, 64(2):203–235, October 2020. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/13439.htm.

Zahran:2014:RTW

[Zah14] Alyaa R. Zahran. An rspca-type wealth index for telephone surveys. Advances and Applications in Statistics, 43(1):53–63, November 2014. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/8845.htm.

Zaka:2020:EGP

[ZAJ20] Azam Zaka, Ahmad Saeed Akhter, and Riffat Jabeen. The exponentiated generalized power function distribution: theory and real life applications. Advances and Applications in Statistics, 61(1):33–63, March 2020. CODEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/abstract/13095.htm.

Zakari:2020:HPR

[Zak20] Ibrahim Sidi Zakari. Harnessing the potential of real-time data analysis and statistical modeling for mapping exposure to air pollution. Advances and Applications in Statistics, 61(2):169– 181, April 2020. CODEN ???? ISSN 0972-3617. URL http:

//www.pphmj.com/abstract/13143.htm.

Zaman:2019:EEP

[Zam19] Tolga Zaman. Efficient estimators of population mean using auxiliary attribute in stratified random sampling. Advances and Applications in Statistics, 56(2):153–171, June 2019. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/12641.htm.

Zaroudi:2018:CAF

[ZBF18] Samira Zaroudi, Mohammad Hassan Behzadi, and Mohammad Reza Faridrohani. A copula approach for finding the type of dependency with mortality force function in insurance market. Advances and Applications in Statistics, 53(2):103–121, August 2018. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/11843.htm.

Zghoul:2003:ECL

[Zgh03] Ahmad A. Zghoul. Estimating the Cauchy location and scale parameters using empirical characteristic function. Advances and Applications in Statistics, 3(2):173–181, August 2003. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/171.htm.

Zhang:2020:MCT

[Zha20] Yanhui Zhang. K-means clustering to ttr based lexical diversity analysis. Advances and Applications in Statistics, 64(2): 267–276, October 2020. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/13442.htm.

Zhang:2009:APU

[ZHMK09] Jing Zhang, Chong Z. He, David R. Mehr, and Robin L. Kruse. Analyzing pressure ulcer development in 36 nursing homes using Bayesian hierarchical modeling. Advances and Applications in Statistics, 12(1):67–84, June 2009. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
4001.htm.

Zimbidis:2008:OMS

[Zim08] Alexandros A. Zimbidis. Optimal management of the safety loading and the exposure to proportional reinsurance coverage of an ordinary insurance system using a diffusion approximation. Advances and Applications in Statistics, 9(1):71–100, June 2008. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/3286.htm.

Zheng:2013:MMV

[ZKZ13] Shimin Zheng, Jeff Knisley, and Chunming Zhang. Moments of matrix variate skew elliptically contoured distributions. Advances and Applications in Statistics, 36(1):13–27, September 2013. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/7903.htm.

Zhang:2006:FCP

[ZLZ06] Riquan Zhang, Yiqiang Lu, and Yusheng Zhang. Functionalcoefficient partially linear time series models for hydrology. Advances and Applications in Statistics, 6(1):57–69, April 2006. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/18.htm.

Zoia:2009:URT

[Zoi09] Maria Grazia Zoia. A unified representation theorem on new algebraic bases, for (co)integrated processes up to the second order. Advances and Applications in Statistics, 12(1):37–66, June 2009. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/4000.htm.

Zimbidis:2008:DHM

[ZP08] Alexandros A. Zimbidis and Athanasios A. Pantelous. Dynamic hedging of the mortality risk via a continuous control strategy of the portfolio of investments of a pension fund. Advances and Applications in Statistics, 8(2):247–289, April 2008. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/3087.htm.

Zhao:2010:PFT

[ZRX10] Yan D. Zhao, Dewi Rahardja, and Xian-Jin Xie. Particle filter for target tracking and applications to bearing-angle problems. Advances and Applications in Statistics, 14(1):17–30, January 2010. CODEN ???? ISSN 0972-3617. URL http://www.

pphmj.com/abstract/4517.htm.

Zurbenko:2001:ALI

[Zur01] Igor Zurbenko. Algorithm of local information. Advances and Applications in Statistics, 1(3):247–275, December 2001. CODEN ???? ISSN 0972-3617. URL http:/

/www.pphmj.com/abstract/93.htm;   http://www.pphmj.
com/abstract/96.htm.

Zhou:2010:SSS

[ZW10] Weihua Zhou and Jin Wang. A simulation study for some graphic-based multivariate two sample tests. Advances and Applications in Statistics, 19(1):1–19, November 2010. CO-DEN ???? ISSN 0972-3617. URL http://www.pphmj.com/

abstract/5492.htm.

Zhao:2005:IMD

[ZX05] Yichuan Zhao and Zhiheng Xu. Inference for the mixed discrete and continuous Cox regression model via empirical likelihood. Advances and Applications in Statistics, 5(2):145–157, August 2005. CODEN ???? ISSN 0972-3617. URL http://

www.pphmj.com/abstract/449.htm.

Zhang:2008:SSS

[ZZF08] Zhiqiang Zhang, Riquan Zhang, and Jingyan Feng. Stationarity of squares sequences from ARCH(1). Advances and Applications in Statistics, 8(2):291–296, April 2008. CODEN ????

ISSN 0972-3617. URL http://www.pphmj.com/abstract/
3088.htm.

Zheng:2013:SRM

[ZZK13] Shimin Zheng, Chunming Zhang, and Jeff Knisley. Stochastic representations of the matrix variate skew elliptically contoured distributions. Advances and Applications in Statistics, 33(2):83–98, April 2013. CODEN ???? ISSN 0972-3617. URL

http://www.pphmj.com/abstract/7557.htm.

Zhang:2013:MCP

[ZZN13] Maojun Zhang, Xueni Zhao, and Jiangxia Nan. The model for CDS pricing based on the Gaussian Copula method. Advances and Applications in Statistics, 36(1):1–11, September 2013. CODEN ???? ISSN 0972-3617. URL http://www.pphmj.

com/abstract/7902.htm.